### v3

# Foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems and their applications for multi-strength intelligent classifiers

Abstract
Hesitant fuzzy sets are widely used in certain instances of uncertainty and hesitation. In sets, the inclusion relationship is an important and foundational definition. Thus, as a kind of set, hesitant fuzzy sets require an explicit definition of inclusion relationship. Based on the hesitant fuzzy membership degree of discrete form, several kinds of inclusion relationships for hesitant fuzzy sets are proposed in this work. Then, some foundational propositions of hesitant fuzzy sets are presented, along with propositions of families of hesitant fuzzy sets. Some foundational propositions of hesitant fuzzy information systems are proposed with respect to parameter reductions and an example and an algorithm are given to illustrate the processes of parameter reduction. Finally, a multi-strength intelligent classifier is proposed to make health state diagnoses for complex systems.
keywords:
Hesitant fuzzy sets, hesitant fuzzy information systems, parameter reductions, multi-strength intelligent classifier
1 Introduction
Set theory established by Cantor FP (hereafter referred to as classical set) is the foundation of modern mathematics. Meanwhile, modern mathematics, such as group theory EJB ; DDJ , topology SWA , graph theory GHH and so on, are the foundations of computer science. After classical set theory was perfected, other kinds of sets were developed, such as fuzzy set ZADE , interval-valued fuzzy set GOMB , intuitionistic fuzzy set AKT and hesitant fuzzy set TV and so on. Several kinds of sets had developed their foundational theoretical systems except hesitant fuzzy set. Especially, one of the most foundational definitions, the definition of the inclusion relationship between two hesitant fuzzy sets, is not sufficiently clear. This article develops the foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems, and proposes a multi-strength intelligent classifier to make health state diagnoses for complex systems.
For the classical set FP , a set describes the membership degree of an element in two cases, namely, 0% and 100%. The relationship between an element and a set is determined depending on the element in or not in the set. However, there are uncertainties in the real world. To deal with the uncertain cases, Zadeh proposed the concept of fuzzy set ZADE wherein the membership degree of an element of a set is a value in $[0,1]$. The membership degree of a fuzzy set represents the probability that an element belongs to a set TZH ; YSB and is no longer just 0% and 100%. Subsequently, Gorzalczany GOMB proposed the concept of interval-valued fuzzy set in which the membership degree of an element of a set is an interval included in $[0,1]$. Interval-valued fuzzy set can deal with cases wherein we cannot determine the exact probability of an element belonging to a set, but we can determine an interval that the probability must appear in this interval BSO ; FKM . Furthermore, Atanassov AKT proposed the concept of intuitionistic fuzzy set wherein the membership degree of an element of a set is a bibasic array. Intuitionistic fuzzy set can describe things with a dual character, one number of the bibasic array expresses the positive aspect and the other number expresses the negative aspect KRA ; PLDY .
Torra TV proposed the concept of hesitant fuzzy set wherein the membership degree of an element of a set is a multidimensional array. Hesitant fuzzy sets focus on cases of uncertainty and hesitation, which are common in the real world. For instance, about the data sampling for equipment monitoring and decision scoring by an expert team, the feedback of such cases always comes in the form of arrays, and the multiple different values can make us indecisive. Thus, hesitant fuzzy sets with a membership degree of a discrete array are more suitable in describing the indecisiveness in decisions compared with the classical, fuzzy, interval-valued fuzzy, and intuitionistic fuzzy sets. Hesitant fuzzy sets are widely used in different fields, such as decision-making HJH ; LLYX ; LNZQY ; GXLY , risk evaluation HXZS , agglomerative hierarchical clustering JLGL , investment policy DYM , modeling for complex product manufacturing tasks BLYY , linguistic perceptual AWNJE , forecasting HLJW ; PMNME , and so on.
However, there is no sufficiently clear definition of the inclusion relationship between two hesitant fuzzy sets. The definition of inclusion relationship is a significant foundation for sets, because if a set is a subset of another set and vice versa, then these two sets are equivalent. Based on the hesitant fuzzy membership degree of discrete form, we propose in this work several kinds of inclusion relationships for hesitant fuzzy sets and then present some foundational propositions of hesitant fuzzy sets and the families of hesitant fuzzy sets. It is worth noting that some foundational propositions of classical sets do not hold in hesitant fuzzy sets.
In recent years, information systems with uncertain information have become a hot topic due to the rising instances of uncertain information in the real world. Plenty of studies have been conducted on this topic, such as fuzzy information system NXSZ , fuzzy incomplete information system WWZJ , interval-valued information system DJWZ ; SDZX , and intuitionistic fuzzy information system HBLH . Based on the foundational propositions of hesitant fuzzy sets, we then study some theories with respect to parameter reductions in hesitant fuzzy information systems.
Analyzing high-dimensional data with small sample sizes is a challenge, the health status of a complex system with such data is difficult to diagnose. Based on the foundational propositions of hesitant fuzzy information systems, a multi-strength intelligent classifier in knowledge bases PALA ; PASA is proposed to make health state diagnoses for complex systems. This multi-strength intelligent classifier is a candidate tool for dimensionality reduction of high-dimensional data, classification, diagnosis, evaluation, decision-making, and so on.
The remainder of the paper is constructed as follows. Section 2 presents several kinds of inclusion relationships of hesitant fuzzy sets, while some foundational propositions of hesitant fuzzy sets are discussed, showing some rules that hold in classical sets or hesitating fuzzy sets. Section 3 presents some propositions on families of hesitant fuzzy sets. Section 4 proposes some propositions with respect to parameter reductions in hesitant fuzzy information systems and provides an algorithm for parameter reductions. Section 5 proposes a multi-strength intelligent classifier in knowledge bases to make health state diagnoses for complex systems. Section 6 presents a conclusion.
2 Foundations of hesitant fuzzy sets
In this section, we briefly recall some basic concepts about hesitant fuzzy sets and then present several kinds of inclusion relationships of hesitant fuzzy sets that are suitable for hesitant fuzzy membership degree of discrete form. We focus on whether some rules of classical sets hold in hesitant fuzzy sets. In subsequent studies of hesitant fuzzy sets, researchers must avoid the intuitions of classical sets as well as the rules held in the classical sets but not in the hesitant fuzzy sets.
If two sets $A$ and $B$ are classical sets, $A\sqcap B$ and $A\sqcup B$ represent the intersection and union of $A$ and $B$, respectively. Furthermore, $A\sqsubset B$ represents that $A$ is a subset of the classical set $B$. Let $U$ be a universal set and $E$ be a set of parameters. Let $HF(U)$ be the set of all hesitant fuzzy sets defined over $U$.
2.1 Reviews of fuzzy sets and hesitant fuzzy sets
Definition 2.1.
ZADE A fuzzy set $F$ on $U$ is a mapping $F:U\rightarrow[0,1]$.
Definition 2.2.
ZADE $F_{1}$ and $F_{2}$ are two fuzzy sets on $U$, $F_{1}$ is a subset of $F_{2}$ if $F_{1}(x)\leqslant F_{2}(x)$ for all $x\in U$, denoted as $F_{1}\subset F_{2}$.
Definition 2.3.
TV A hesitant fuzzy set on $U$ is defined as a function that when applied to $U$ returns a subset of $[0,1]$.
Definition 2.4.
TV For each $x\in U$, and a hesitant fuzzy set $H$, the lower bound and upper bound of $H(x)$ are defined
lower bound $H^{-}(x)=min\ H(x)$,
upper bound $H^{+}(x)=max\ H(x)$.
Definition 2.5.
TV Given two hesitant fuzzy sets represented by their membership functions $H_{1}$ and $H_{2}$, their union and intersection are defined
union $(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant max(H_{1}^{-}(x),H_{2}^{-}(x))\}$,
intersection $(H_{1}\cap H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\leqslant min(H_{1}^{+}(x),H_{2}^{+}(x))\}$.
Let $U=\{x,y\}$, there are three common expression forms of hesitant fuzzy set, such as $A=\{\frac{0.6,0.5,0.3}{x},\frac{0.5,0.3,0.2}{y}\}$, $A=\frac{0.6,0.5,0.3}{x}+\frac{0.5,0.3,0.2}{y}$ and $A=\{<x,(0.6,0.5,0.3)>,<y,(0.5,0.3,0.2)>\}$. $A=B$ means that $A$ and $B$ are perfectly consistent in this article, i.e., $B=\frac{0.3,0.6,0.5}{x}+\frac{0.2,0.5,0.3}{y}$.
2.2 Inclusion relationships of hesitant fuzzy sets
For two hesitant fuzzy sets $H_{1}$ and $H_{2}$, Babitha et al BJ introduced that $H_{1}$ is a hesitant fuzzy subset of $H_{2}$ if $H_{1}(x)\subset H_{2}(x)$ for all $x\in U$. However, what is the meaning of $H_{1}(x)\subset H_{2}(x)$? There is no detailed description in BJ .
Carlos et al JCVT introduced a definition of inclusion relationship for hesitant fuzzy sets, i.e., $H_{1}\subset H_{2}$Leftrightarrow $H_{1}\cup H_{2}=H_{2}$. By the description in JCVT , $H_{1}(x)\subset H_{2}(x)\Leftrightarrow\{h:h\in H_{2}(x)\}=H_{2}(x)=(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant\max(H_{1}^{-}(x),H_{2}^{-}(x))\}$, i.e., $H_{1}(x)\subset H_{2}(x)$ implies $h<H_{2}^{-}(x)$ for all $h\in H_{1}(x)$. The definition of inclusion relationship for hesitant fuzzy sets in JCVT is not perfect enough for these aspects, shown as (i) and (ii),
(i) The inclusion relationship of sets is also the ordering relationship of sets. A foundational and important proposition of inclusion relationship, $A\subset B$ and $B\subset A$ if and only if $A=B$, i.e., the equivalent sets contain each other. However, by the definition of inclusion relationship in JCVT , if $H_{1}(x)\subset H_{2}(x)$ and $H_{2}(x)\subset H_{1}(x)$, then a contradiction is produced, i.e., $H_{1}^{+}(x)<H_{2}^{-}(x)<H_{2}^{+}(x)<H_{1}^{-}(x)<H_{1}^{+}(x)$.
(ii) The definition of inclusion relationship in JCVT is too specific to have a small range of applications. For example, let $H_{1}(x)=\{0.1,0.2,0.5\}$, $H_{2}(x)=\{0.6,0.7,0.9\}$ and $H_{3}(x)=\{0.1,0.5,0.8\}$, then $H_{1}(x)\subset H_{2}(x)$ is obvious. However, this definition cannot describe the relationship of $H_{2}(x)$ and $H_{3}(x)$ and is not applicable for many cases of hesitant fuzzy sets.
In this work, we propose several kinds of inclusion relationships for hesitant fuzzy sets based on the strength of the information in hesitant fuzzy membership degrees. First, we present a detailed example to illustrate the Definition 2.7.
Example 2.6.
Let $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\}$ be a set of decision-making schemes and $H$ be an expert team that consists of three experts. $H(U)=\frac{0.9,0.2}{x_{1}}+\frac{0.6,0.6,0.5}{x_{2}}+\frac{0.7,0.5,0.5}{x_{3}}+\frac{0.8,0.6,0.5}{x_{4}}+\frac{0.9,0.3,0.1}{x_{5}}+\frac{0.9,0.8,0.7}{x_{6}}$ are the schemes’ scores that are given by experts, in which the scores for $x_{1}$ are 0.9 and 0.2 that are obtained through the evaluations made by experts. One of the three experts fails to evaluate the scheme $x_{1}$.
(1) Here, $x_{1}$ has a score of 0.9, which is greater than or equal to all the scores for $x_{2}$, it is possible that the scheme $x_{1}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{p}H(x_{1})$.
(2) Here, $0.5670.55=mean[H(x_{1})]<mean[H(x_{2})]=$ 0.567, where $mean[\cdot]$ is the mean value operator. In comparing the mean values of scores, we find that the scheme $x_{2}$ is better than the scheme $x_{1}$, denoted as $H(x_{1})\subset_{m}H(x_{2})$.
(3) On the one hand, the best score of $x_{3}$ is 0.7, which is greater than or equal to the best score of $x_{2}$. On the other hand, the worst score of $x_{3}$ is 0.5, which is greater than or equal to the worst score of $x_{2}$. To compare the respective best and worst cases of schemes $x_{2}$ and $x_{3}$, it is acceptable that the scheme $x_{3}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{a}H(x_{3})$.
(4) To compare the scores for schemes $x_{3}$ and $x_{4}$ one by one ($0.7\leqslant 0.8$; $0.5\leqslant 0.6$; and $0.5\leqslant 0.5$), it is strongly credible that the scheme $x_{4}$ is better than the scheme $x_{3}$, denoted as $H(x_{3})\subset_{s}H(x_{4})$.
(5) We can obtain $H(x_{1})\subset_{s}H(x_{5})$ after truncating the tail score of $x_{5}$, i.e., deleting the score 0.1 of $x_{5}$, this case is denoted as $H(x_{1})\subset_{st}H(x_{5})$ and is recorded briefly as $H(x_{1})\subset_{t}H(x_{5})$.
(6) The worst score of $x_{6}$ is greater than or equal to the best score of $x_{3}$. Thus, it is necessary that the scheme $x_{6}$ is better than the scheme $x_{3}$, which is denoted as $H(x_{3})\subset_{n}H(x_{6})$.
Comparing the mean value of scores is a common thought for decision-making; however, while doing so, some important information may be lost, such as the best and the worst scores.
Definition 2.7.
Let $H_{1}$ and $H_{2}$ be two hesitant fuzzy sets on $U$. Several kinds of inclusion relationships of two hesitant fuzzy sets are defined as follows,
(1) If $H^{+}_{1}(x)\leqslant H^{+}_{2}(x)$, then $H_{1}(x)\subset_{p}H_{2}(x)$. If $H_{1}(x)\subset_{p}H_{2}(x)$ for all $x\in U$, then $H_{1}\subset_{p}H_{2}$. If $H_{1}\subset_{p}H_{2}$ and $H_{2}\subset_{p}H_{1}$, then $H_{1}=_{p}H_{2}$.
(2) If $H^{+}_{1}(x)\leqslant H^{+}_{2}(x)$ and $H^{-}_{1}(x)\leqslant H^{-}_{2}(x)$, then $H_{1}(x)\subset_{a}H_{2}(x)$. If $H_{1}(x)\subset_{a}H_{2}(x)$ for all $x\in U$, then $H_{1}\subset_{a}H_{2}$. If $H_{1}\subset_{a}H_{2}$ and $H_{2}\subset_{a}H_{1}$, then $H_{1}=_{a}H_{2}$.
(3) If $mean[H_{1}(x)]\leqslant mean[H_{2}(x)]$, then $H_{1}(x)\subset_{m}H_{2}(x)$. If $H_{1}(x)\subset_{m}H_{2}(x)$ for all $x\in U$, then $H_{1}\subset_{m}H_{2}$. If $H_{1}\subset_{m}H_{2}$ and $H_{2}\subset_{m}H_{1}$, then $H_{1}=_{m}H_{2}$.
(4) Let $H_{1}(x)=V=\{v_{1},v_{2},\cdots,v_{k}\}$ and $H_{2}(x)=W=\{w_{1},w_{2},\cdots,w_{l}\}$ be two descending sequences. If $k\geqslant l$ and $w_{i}\geqslant v_{i}$ for $1\leqslant i\leqslant l$, then $H_{1}(x)\subset_{s}H_{2}(x)$. If $H_{1}(x)\subset_{s}H_{2}(x)$ for all $x\in U$, then $H_{1}\subset_{s}H_{2}$. If $H_{1}\subset_{s}H_{2}$ and $H_{2}\subset_{s}H_{1}$, then $H_{1}=_{s}H_{2}$.
(5) Let $H_{1}(x)=V=\{v_{1},v_{2},\cdots,v_{k}\}$ and $H_{2}(x)=W=\{w_{1},w_{2},\cdots,w_{l}\}$ be two descending sequences. If $k<l$ and $w_{i}\geqslant v_{i}$ for $1\leqslant i\leqslant k$, then $H_{1}(x)\subset_{t}H_{2}(x)$. If $H_{1}(x)\subset_{t}H_{2}(x)$ for all $x\in U$, then $H_{1}\subset_{t}H_{2}$. It is obvious that $H_{1}\subset_{t}H_{2}$ and $H_{2}\subset_{t}H_{1}$ cannot hold simultaneously.
(6) If $H_{1}^{+}(x)\leqslant H_{2}^{-}(x)$, then $H_{1}(x)\subset_{n}H_{2}(x)$. If $H_{1}(x)\subset_{n}H_{2}(x)$ for all $x\in U$, then $H_{1}\subset_{n}H_{2}$. If $H_{1}\subset_{n}H_{2}$ and $H_{2}\subset_{n}H_{1}$, then $H_{1}=_{n}H_{2}$.
Symbols. (1) Let $W=\{w_{1},w_{2},\cdots,w_{l}\}$ and $W^{\prime}=\{w_{1}^{\prime},w_{2}^{\prime},\cdots,w_{l}^{\prime}\}$ be two descending sequences and $|W|=|W^{\prime}|$. If $w_{1}^{\prime}\geqslant w_{1}$, $w_{2}^{\prime}\geqslant w_{2}$$\cdots$,$w_{l}^{\prime}\geqslant w_{l}$, then the relationship of $W$ and $W^{\prime}$ is denoted as $W\preccurlyeq W^{\prime}$.
(2) Let $q$ be a positive integer. To take $q$ larger numbers of $W$ and construct a best $q$-subsequence of $W$, denoted as $W(q)$, i.e., $|W(q)|=q$ and $w\leqslant min(W(q))$ for $w\in W-W(q)$.
Remark 2.8.
(1) If $W$ is a subsequence of $H_{2}(x)$, $H_{1}(x)\preccurlyeq W$ and $|H_{1}(x)|<|H_{2}(x)|$, then $H_{1}(x)\subset_{t}H_{2}(x)$, vice versa.
(2) If $W$ is a subsequence of $H_{2}(x)$, $H_{1}(x)\preccurlyeq W$ and $|H_{1}(x)|=|H_{2}(x)|$, then $H_{1}(x)\subset_{s}H_{2}(x)$.
(3) Let $q=min\{|H_{1}(x)|,|H_{2}(x)|\}$, $H_{1}(x)(q)$ and $H_{2}(x)(q)$ be the best $q$-subsequences of $H_{1}(x)$ and $H_{2}(x)$, respectively. Thereafter, $H_{1}(x)(q)\preccurlyeq H_{2}(x)(q)$ if and only if one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.
(4) Let $K=H_{1}(x)\sqcap H_{2}(x)$, if $max(H_{1}(x)-K)\leqslant min(K)\leqslant max(K)\leqslant min(H_{2}(x)-K)$, then one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.
Proposition 2.9.
The following statements hold for $A,B\in HF(U)$,
(1) If $A\subset_{a}B$, then $A\subset_{p}B$.
(2) If $A\subset_{s}B$, then $A\subset_{p}B$.
(3) If $A\subset_{s}B$, then $A\subset_{a}B$.
(4) If $A\subset_{s}B$, then $A\subset_{m}B$.
(5) If $A\subset_{t}B$, then $A\subset_{p}B$.
(6) If $A\subset_{n}B$, then $A\subset_{p}B$.
(7) If $A\subset_{n}B$, then $A\subset_{a}B$.
(8) If $A\subset_{n}B$, then $A\subset_{m}B$.
(9) If $A\subset_{n}B$, then one of $A(x)\subset_{s}B(x)$ and $A(x)\subset_{t}B(x)$ holds for all $x\in U$.
Proof (1)-(8) are obvious.
(9) If $A\subset_{n}B$, then $A^{+}(x)\leqslant B^{-}(x)$ for all $x\in U$. If $|A(x)|\geqslant|B(x)|$, then $A(x)\subset_{s}B(x)$. If $|A(x)|<|B(x)|$, then $A(x)\subset_{t}B(x)$$\blacksquare$
Proposition 2.10.
The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$ and $B\subset_{p}C$, then $A\subset_{p}C$.
(2) If $A\subset_{a}B$ and $B\subset_{a}C$, then $A\subset_{a}C$.
(3) If $A\subset_{m}B$ and $B\subset_{m}C$, then $A\subset_{m}C$.
(4) If $A\subset_{s}B$ and $B\subset_{s}C$, then $A\subset_{s}C$.
(5) If $A\subset_{t}B$ and $B\subset_{t}C$, then $A\subset_{t}C$.
(6) If $A\subset_{n}B$ and $B\subset_{n}C$, then $A\subset_{n}C$.
Proof (1) and (2) are obvious.
(3) $A\subset_{m}B$ and $B\subset_{m}C$ imply that $mean[A(x)]\leqslant mean[B(x)]\leqslant mean[C(x)]$ for all $x\in U$, where $mean[\cdot]$ is the mean value operator. Then $A\subset_{m}C$.
(4) $A\subset_{s}B$ and $B\subset_{s}C$ imply that $|A(x)|\geqslant|B(x)|\geqslant|C(x)|$ for all $x\in U$.
Let $C(x)=\{h^{C}_{1},h^{C}_{2},\cdots,h^{C}_{|C(x)|}\}$, $B(x)=\{h^{B}_{1},h^{B}_{2},\cdots,h^{B}_{|B(x)|}\}$ and $A(x)=\{h^{A}_{1},h^{A}_{2},\cdots,h^{A}_{|A(x)|}\}$ be three descending sequences, since $A(x)\subset_{s}B(x)$ and $B(x)\subset_{s}C(x)$, then $h^{C}_{1}\geqslant h^{B}_{1}\geqslant h^{A}_{1}$, $h^{C}_{2}\geqslant h^{B}_{2}\geqslant h^{A}_{2}$$\cdots$, and $h^{C}_{|C(x)|}\geqslant h^{B}_{|C(x)|}\geqslant h^{A}_{|C(x)|}$. Hence, $A\subset_{s}C$.
(5) $A\subset_{t}B$ and $B\subset_{t}C$ imply that $|A(x)|<|B(x)|<|C(x)|$ for all $x\in U$.
Let $C(x)=\{h^{C}_{1},h^{C}_{2},\cdots,h^{C}_{|C(x)|}\}$, $B(x)=\{h^{B}_{1},h^{B}_{2},\cdots,h^{B}_{|B(x)|}\}$ and $A(x)=\{h^{A}_{1},h^{A}_{2},\cdots,h^{A}_{|A(x)|}\}$ be three descending sequences, since $A(x)\subset_{t}B(x)$ and $B(x)\subset_{t}C(x)$, then $h^{C}_{1}\geqslant h^{B}_{1}\geqslant h^{A}_{1}$, $h^{C}_{2}\geqslant h^{B}_{2}\geqslant h^{A}_{2}$$\cdots$, and $h^{C}_{|A(x)|}\geqslant h^{B}_{|A(x)|}\geqslant h^{A}_{|A(x)|}$. Hence, $A\subset_{t}C$.
(6) $A\subset_{n}B$ and $B\subset_{n}C$ imply that $A^{+}(x)\leqslant B^{-}(x)\leqslant B^{+}(x)\leqslant C^{-}(x)$ for all $x\in U$, then $A\subset_{n}C$$\blacksquare$
Let $\bar{A}$ and $\bar{B}$ be two classical sets, we know that $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}$, $\bar{A}\sqcap\bar{B}\sqsubset\bar{B}$, $\bar{A}\sqsubset\bar{A}\sqcup\bar{B}$, $\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$, and $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$. However, these rules are not necessarily true for two hesitant fuzzy sets $A$ and $B$ (shown as Propositions 2.11-2.18).
Proposition 2.11.
The following statements hold for $A,B\in HF(U)$,
(1) $A\cap B\subset_{p}A$.
(2) $A\cap B\subset_{p}B$.
(3) $A\subset_{p}A\cup B$.
(4) $B\subset_{p}A\cup B$.
(5) $A\cap B\subset_{p}A\cup B$.
Proof (1) $min\{A^{+}(x),B^{+}(x)\}\leqslant A^{+}(x)$.
(2) $min\{A^{+}(x),B^{+}(x)\}\leqslant B^{+}(x)$.
(3) $A^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$.
(4) $B^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$.
(5) $min\{A^{+}(x),B^{+}(x)\}\leqslant max\{A^{+}(x),B^{+}(x)\}$$\blacksquare$
Proposition 2.12.
The following statements hold for $A,B\in HF(U)$,
(1) $A\cap B\subset_{a}A$.
(2) $A\cap B\subset_{a}B$.
(3) $A\subset_{a}A\cup B$.
(4) $B\subset_{a}A\cup B$.
(5) $A\cap B\subset_{a}A\cup B$.
Proof (1) $min\{A^{+}(x),B^{+}(x)\}\leqslant A^{+}(x)$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant A^{-}(x)$.
(2) $min\{A^{+}(x),B^{+}(x)\}\leqslant B^{+}(x)$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant B^{-}(x)$.
(3) $A^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $A^{-}(x)\leqslant max\{A^{-}(x),B^{-}(x)\}$.
(4) $B^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $B^{-}(x)\leqslant max\{A^{-}(x),B^{-}(x)\}$.
(5) $min\{A^{+}(x),B^{+}(x)\}\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant max\{A^{-}(x),B^{-}(x)\}$$\blacksquare$
Proposition 2.13.
The following statements hold for $x\in U$ and $A,B\in HF(U)$,
(1) At least one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
(2) At least one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
(3) $A\cap B\subset_{m}A\cup B$.
(4) If $A\subset_{m}B$, then $A\cap B\subset_{m}B$.
(5) If $A\subset_{m}B$, then $A\subset_{m}A\cup B$.
Proof For $x\in U$, denote $k_{1}=|A(x)|$, $k_{2}=|B(x)|$, $\bar{a}=mean[A(x)]$, and $\bar{b}=mean[B(x)]$, where $mean[\cdot]$ is the mean operator. Then $\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}$.
(1) For each $x\in U$, we can divide the relationships of $A^{+}(x)$ and $B^{+}(x)$ into two cases (c1) and (c2) as follows,
(c1) $A^{+}(x)=B^{+}(x)$.
$\sum\limits_{h\in(A\cap B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})$. $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=max(\bar{a},\bar{b})$. Then, one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
(c2) $A^{+}(x)\neq B^{+}(x)$.
$A^{+}(x)\neq B^{+}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=min(V)$, then $h^{\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$.
Since $h^{\prime}>max\{h:h\in(A\cap B)(x)\}$ and $|(A\cap B)(x)|+|V|=k_{1}+k_{2}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=min(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
Then, we can obtain $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=max(\bar{a},\bar{b})$. Then, one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
(2) For each $x\in U$, we can divide the relationships of $A^{-}(x)$ and $B^{-}(x)$ into two cases (c1) and (c2) as follows,
(c1) $A^{-}(x)=B^{-}(x)$.
$\sum\limits_{h\in(A\cup B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})$.
$\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=min(\bar{a},\bar{b})$. Then, one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
(c2) $A^{-}(x)\neq B^{-}(x)$.
$A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$. $h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$.
Since $h^{\prime}=max(V)$ and $|(A\cup B)(x)|+|V|=k_{1}+k_{2}$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$. Since $h^{\prime}<min\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=min(\bar{a},\bar{b})$. Then, one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
(3) For each $x\in U$, we can divide the relationships of $A^{-}(x)$, $A^{+}(x)$, $B^{-}(x)$ and $B^{+}(x)$ into four cases (c1), (c2), (c3) and (c4) as follows,
(c1) $A^{-}(x)=B^{-}(x)$ and $A^{+}(x)=B^{+}(x)$.
In this case, $(A\cap B)(x)=(A\cup B)(x)$, thus $mean[(A\cap B)(x)]=mean[(A\cup B)(x)]$, then $(A\cap B)(x)\subset_{m}(A\cup B)(x)$ is obvious.
(c2) $A^{-}(x)=B^{-}(x)$ and $A^{+}(x)\neq B^{+}(x)$.
In this case, $|(A\cap B)(x)|<k_{1}+k_{2}$ and $|(A\cup B)(x)|=k_{1}+k_{2}$. Let $V=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=min(V)$. $h^{\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$.
Since $|(A\cap B)(x)|+|V|=|(A\cup B)(x)|=k_{1}+k_{2}$ and $h^{\prime}>max\{h:h\in(A\cap B)(x)\}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=min(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})$.
Based on the results above, then $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})=mean[(A\cup B)(x)]$, i.e., $(A\cap B)(x)\subset_{m}(A\cup B)(x)$.
(c3) $A^{-}(x)\neq B^{-}(x)$ and $A^{+}(x)=B^{+}(x)$.
In this case, $|(A\cap B)(x)|=k_{1}+k_{2}$ and $|(A\cup B)(x)|<k_{1}+k_{2}$. Let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$. $h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$.
Since $|(A\cup B)(x)|+|V|=|(A\cap B)(x)|=k_{1}+k_{2}$ and $h^{\prime}<min\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=max(V)$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})$.
Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})=mean[(A\cap B)(x)]$, i.e., $(A\cap B)(x)\subset_{m}(A\cup B)(x)$.
(c4) $A^{-}(x)\neq B^{-}(x)$ and $A^{+}(x)\neq B^{+}(x)$.
Based on the condition $A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$, then $h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$, and $|V|+|(A\cup B)(x)|=k_{1}+k_{2}$.
Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
Based on the condition $A^{+}(x)\neq B^{+}(x)$, let $W=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime\prime}=min(W)$, then $h^{\prime\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$, and $|W|+|(A\cap B)(x)|=k_{1}+k_{2}$.
Based on the results above, then $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime\prime}|W|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})<mean[(A\cup B)(x)]$, i.e., $(A\cap B)(x)\subset_{m}(A\cup B)(x)$.
To sum up, $A\cap B\subset_{m}A\cup B$.
(4) If $A\subset_{m}B$, by (1), then $A\cap B\subset_{m}B$ is obtained.
(5) If $A\subset_{m}B$, by (2), then $A\subset_{m}A\cup B$ is obtained.$\blacksquare$
Example 2.14.
Let $U=\{x,y,z\}$, $A=\frac{\{0.1,0.8\}}{x}+\frac{\{0.1,0.8\}}{y}+\frac{\{0.7,0.9\}}{z}$, $B=\frac{\{0.7,0.9\}}{x}+\frac{\{0.1,0.9\}}{y}+\frac{\{0.1,0.8\}}{z}$.
$(A\cap B)(x)=\{0.1,0.7,0.8\}$, $mean[(A\cap B)(x)]=0.53$. $mean[A(x)]=0.45$. $(A\cap B)(x)\not\subset_{m}A(x)$.
$(A\cup B)(y)=\{0.1,0.1,0.8,0.9\}$, $mean[(A\cup B)(y)]=0.475$. $mean[B(y)]=0.5$. $B(y)\not\subset_{m}(A\cup B)(y)$.
$(A\cap B)(y)\subset_{m}A(y)$ and $(A\cap B)(y)\subset_{m}B(y)$.
$A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$.
For random two hesitant fuzzy sets $A$ and $B$, Example 2.14 shows that $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ do not hold simultaneously, $A(y)\subset_{m}(A\cup B)(y)$ and $B(y)\subset_{m}(A\cup B)(y)$ do not hold simultaneously.
$(A\cap B)(x)\subset_{m}B(x)$ and $(A\cap B)(x)\not\subset_{m}A(x)$. Furthermore, $(A\cap B)(z)\subset_{m}A(z)$ and $(A\cap B)(z)\not\subset_{m}B(z)$. It means that (1) of Proposition 2.13 cannot be written as that one of $A\cap B\subset_{m}A$ and $A\cap B\subset_{m}B$ holds.
Proposition 2.15.
The following statements hold for $x\in U$ and $A,B\in HF(U)$,
(1) At least one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
(2) At least one of $B(x)\subset_{s}(A\cup B)(x)$ and $B(x)\subset_{t}(A\cup B)(x)$ holds.
(3) At least one of $(A\cap B)(x)\subset_{s}(A\cup B)(x)$ and $(A\cap B)(x)\subset_{t}(A\cup B)(x)$ holds.
Proof (1) For each $x\in U$, if $A^{-}(x)\geqslant B^{-}(x)$, then $h\in(A\cup B)(x)$ for all $h\in A(x)$, i.e., $A(x)$ is a subsequence of $(A\cup B)(x)$. $A(x)\preccurlyeq A(x)$ and $|A(x)|\leqslant|(A\cup B)(x)|$, by Remark 2.8 (1) and (2), then one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
If $A^{-}(x)<B^{-}(x)$, let $K=A(x)\sqcap(A\cup B)(x)$, then $max(A(x)-K)<min(K)\leqslant max(K)<max((A\cup B)(x)-K)$, by Remark 2.8 (4), then one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
(2) It can be proved by the similar method of (1).
(3) Let $K=(A\cap B)(x)\sqcap(A\cup B)(x)$, then $max((A\cap B)(x)-K)<min(K)\leqslant max(K)<min((A\cup B)(x)-K)$, by Remark 2.8 (4), one of $(A\cap B)(x)\subset_{s}(A\cup B)(x)$ and $(A\cap B)(x)\subset_{t}(A\cup B)(x)$ holds.$\blacksquare$
If two inclusion relationships $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ in Definition 2.7 are combined in $H_{1}(x)\subset_{sot}H_{2}(x)$ ($\subset_{sot}$ means $\subset_{s}$ or $\subset_{t}$), then Proposition 2.15 can be described as follows:
Proposition ${}^{\prime}$ The following statements hold for $A,B\in HF(U)$,
(1) $A\subset_{sot}A\cup B$.
(2) $B\subset_{sot}A\cup B$.
(3) $A\cap B\subset_{sot}A\cup B$.
Proposition 2.16.
The following statements hold for $x\in U$ and $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$, then at least one of $A(x)\subset_{s}(A\cap B)(x)$ and $A(x)\subset_{t}(A\cap B)(x)$ holds.
(2) If $A\subset_{a}B$, then at least one of $A(x)\subset_{s}(A\cap B)(x)$ and $A(x)\subset_{t}(A\cap B)(x)$ holds.
(3) If $A\subset_{s}B$, then at least one of $A(x)\subset_{s}(A\cap B)(x)$ and $A(x)\subset_{t}(A\cap B)(x)$ holds.
(4) If $A\subset_{t}B$, then at least one of $A(x)\subset_{s}(A\cap B)(x)$ and $A(x)\subset_{t}(A\cap B)(x)$ holds.
(5) If $A\subset_{n}B$, then at least one of $A(x)\subset_{s}(A\cap B)(x)$ and $A(x)\subset_{t}(A\cap B)(x)$ holds.
(6) For a random $C\in HF(U)$, if $A\subset_{s}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(7) For a random $C\in HF(U)$, if $A\subset_{t}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(8) For a random $C\in HF(U)$, if $A\subset_{n}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
Proof (1) $A\subset_{p}B$ implies that $A^{+}(x)\leqslant B^{+}(x)$ for all $x\in U$. So we can deduce that if $h\in A(x)$ then $h\in(A\cap B)(x)$, i.e., $A(x)$ is a subsequence of $(A\cap B)(x)$. $A(x)\preccurlyeq A(x)$ and $|A(x)|\leqslant|(A\cap B)(x)|$, by Remark 2.8 (1) and (2), then one of $A(x)\subset_{s}(A\cap B)(x)$ and $A(x)\subset_{t}(A\cap B)(x)$ holds.
(2) $A\subset_{a}B$ implies $A\subset_{p}B$, by the result of (1), then (2) holds.
(3) $A\subset_{s}B$ implies $A\subset_{p}B$, by the result of (1), then (3) holds.
(4) $A\subset_{t}B$ implies $A\subset_{p}B$, by the result of (1), then (4) holds.
(5) $A\subset_{n}B$ implies $A\subset_{p}B$, by the result of (1), then (5) holds.
(6) We prove this proposition after dividing the relationships of $h\in C(x)$ and the interval $[B^{-}(x),B^{+}(x)]$ into two cases (i) and (ii) for each $x\in U$.
(i) $h\not\in[B^{-}(x),B^{+}(x)]$ for all $h\in C(x)$.
This case includes two sub-cases, $B^{+}(x)<C^{-}(x)$ and $C^{+}(x)<B^{-}(x)$.
If $B^{+}(x)<C^{-}(x)$, then $A(x)\subset_{s}B(x)\subset_{n}C(x)=(B\cup C)(x)$, then $A^{+}(x)\leqslant B^{+}(x)<C^{-}(x)$, i.e., $A(x)\subset_{n}C(x)=(B\cup C)(x)$. $A(x)\subset_{n}(B\cup C)(x)$, by Proposition 2.9 (9), then one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
If $C^{+}(x)<B^{-}(x)$, then $A(x)\subset_{s}B(x)=(B\cup C)(x)$.
(ii) $h\in[B^{-}(x),B^{+}(x)]$ for some $h\in C(x)$.
This case includes two sub-cases, $C^{-}(x)\leqslant B^{-}(x)$ and $B^{-}(x)<C^{-}(x)$. Suppose $A(x)=\{o_{1},o_{2},\cdots,o_{a}\}=O$, $B(x)=\{v_{1},v_{2},\cdots,v_{b}\}=V$ and $C(x)=\{w_{1},w_{2},\cdots,w_{c}\}=W$, and $O$, $V$ and $W$ are descending sequences.
When $C^{-}(x)\leqslant B^{-}(x)$.
It is obvious that $h\in(B\cup C)(x)$ for all $h\in B(x)$. We can divide the elements of $A(x)$ into two descending sequences $\{o_{1},o_{2},\cdots,o_{b}\}$ and $\{o_{b+1},o_{b+2},\cdots,o_{a}\}$.
Since $A(x)\subset_{s}B(x)$, for $o_{i}\in\{o_{1},o_{2},\cdots,o_{b}\}$, we have $v_{i}\geqslant o_{i}$ and $v_{i}\in(B\cup C)(x)$, where $i=1,2,\cdots,b$.
For $o_{j}\in\{o_{b+1},o_{b+2},\cdots,o_{a}\}\sqsubset A(x)$ and all $h\in(B\cup C)(x)$, we have $h\geqslant B^{-}(x)=v_{b}\geqslant o_{b}\geqslant o_{b+1}\geqslant o_{j}$, where $j=b+1,b+2,\cdots,a$.
Based on the illustrations of sub-case $C^{-}(x)\leqslant B^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
When $B^{-}(x)<C^{-}(x)$.
Suppose $v_{m-1}\geqslant w_{c}=C^{-}(x)$ and $v_{m}<w_{c}$. For all $h\in(B\cup C)(x)$, $h\geqslant w_{c}=C^{-}(x)$. We can divide the elements of $A(x)$ into two descending sequences $\{o_{1},o_{2},\cdots,o_{m-1}\}$ and $\{o_{m},o_{m+1},\cdots,o_{a}\}$.
For $o_{i}\in\{o_{1},o_{2},\cdots,o_{m-1}\}$, we have $v_{i}\geqslant o_{i}$ and $v_{i}\in(B\cup C)(x)$, where $i=1,2,\cdots,m-1$.
For $o_{j}\in\{o_{m},o_{m+1},\cdots,o_{a}\}\sqsubset A(x)$ and all $h\in(B\cup C)(x)$, we have $h\geqslant C^{-}(x)=w_{c}>v_{m}\geqslant o_{m}\geqslant o_{j}$, where $j=m,m+1,\cdots,a$.
Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(7) We prove this proposition after dividing the relationships of $h\in C(x)$ and the interval $[B^{-}(x),B^{+}(x)]$ into two cases (i) and (ii) for each $x\in U$.
(i) $h\not\in[B^{-}(x),B^{+}(x)]$ for all $h\in C(x)$.
This case includes two sub-cases, $B^{+}(x)<C^{-}(x)$ and $C^{+}(x)<B^{-}(x)$.
When $B^{+}(x)<C^{-}(x)$, we have $A(x)\subset_{t}B(x)\subset_{n}C(x)=(B\cup C)(x)$, then $A^{+}(x)\leqslant B^{+}(x)<C^{-}(x)$, i.e., $A(x)\subset_{n}C(x)=(B\cup C)(x)$. $A(x)\subset_{n}(B\cup C)(x)$, by Proposition 2.9 (9), then one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
When $C^{+}(x)<B^{-}(x)$, $A(x)\subset_{t}B(x)=(B\cup C)(x)$.
(ii) $h\in[B^{-}(x),B^{+}(x)]$ for some $h\in C(x)$.
This case includes two sub-cases, $C^{-}(x)\leqslant B^{-}(x)$ and $B^{-}(x)<C^{-}(x)$.
When $C^{-}(x)\leqslant B^{-}(x)$.
We have $h\in(B\cup C)(x)$ for all $h\in B(x)$, i.e., $B(x)\sqsubset(B\cup C)(x)$. Since $A(x)\subset_{t}B(x)$, by Remark 2.8 (1), there is a subsequence of $B(x)$, denoted as $\overline{B(x)}$, satisfying $A(x)\preccurlyeq\overline{B(x)}$. Since $B(x)\sqsubset(B\cup C)(x)$, then there is a subsequence of $(B\cup C)(x)$, denoted as $\overline{(B\cup C)(x)}$, satisfying $A(x)\preccurlyeq\overline{(B\cup C)(x)}$. $|A(x)|<|B(x)|\leqslant|(B\cup C)(x)|$ is obvious. By Remark 2.8 (1), $A(x)\subset_{t}(B\cup C)(x)$ holds.
When $B^{-}(x)<C^{-}(x)$.
Suppose $A(x)=\{o_{1},o_{2},\cdots,o_{a}\}=O$, $B(x)=\{v_{1},v_{2},\cdots,v_{b}\}=V$ and $C(x)=\{w_{1},w_{2},\cdots,w_{c}\}=W$, and $O$, $V$ and $W$ are descending sequences.
Suppose $v_{m-1}\geqslant w_{c}$ and $v_{m}<w_{c}=C^{-}(x)$. We can divide the elements of $A(x)$ into two descending sequences $\{o_{1},o_{2},\cdots,o_{m-1}\}$ and $\{o_{m},o_{m+1},\cdots,o_{a}\}$.
Since $A(x)\subset_{t}B(x)$ and $v_{m-1}\geqslant w_{c}=C^{-}(x)$, then $\{o_{1},o_{2},\cdots,o_{m-1}\}\preccurlyeq\{v_{1},v_{2},\cdots,v_{m-1}\}$ for $\{o_{1},o_{2},\cdots,o_{m-1}\}\sqsubset A(x)$ and $\{v_{1},v_{2},\cdots,v_{m-1}\}\sqsubset(B\cup C)(x)$.
Since $o_{m}\leqslant v_{m}<w_{c}=C^{-}(x)$, for $o_{j}\in\{o_{m},o_{m+1},\cdots,o_{a}\}\sqsubset A(x)$ and all $h\in(B\cup C)(x)$, it implies $h>o_{j}$, where $j=m,m+1,\cdots,a$.
Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$ is obtained, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(8) By Proposition 2.9 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for all $x\in U$. By the results of (6) and (7), (8) holds.$\blacksquare$
Proposition ${}^{\prime}$ The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$, then $A\subset_{sot}A\cap B$.
(2) If $A\subset_{a}B$, then $A\subset_{sot}A\cap B$.
(3) If $A\subset_{sot}B$, then $A\subset_{sot}A\cap B$.
(4) If $A\subset_{n}B$, then $A\subset_{sot}A\cap B$.
(5) If $A\subset_{sot}B$, then $A\subset_{sot}B\cup C$ for a random $C\in HF(U)$.
(6) If $A\subset_{n}B$, then $A\subset_{sot}B\cup C$ for a random $C\in HF(U)$.
Example 2.17.
Let $U=\{x,y\}$, $A=\frac{\{0.1,0.2,0.5,0.6,0.9\}}{x}+\frac{\{0.1,0.7\}}{y}$, $B=\frac{\{0.05,0.3,0.4,0.7,0.8\}}{x}+\frac{\{0.8,0.9,0.9\}}{y}$.
$(A\cap B)(x)=\{0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8\}$, then $(A\cap B)(x)\not\subset_{s}A(x)$, $(A\cap B)(x)\not\subset_{s}B(x)$, $(A\cap B)(x)\not\subset_{t}A(x)$, and $(A\cap B)(x)\not\subset_{t}B(x)$.
$(A\cup B)(y)=\{0.8,0.9,0.9\}$, then $A(y)\subset_{t}(A\cup B)(y)$ and $B(y)\subset_{s}(A\cup B)(y)$.
Proposition 2.18.
The following statements hold for $A,B\in HF(U)$,
(1) If $A\subset_{n}B$, $A\cap B\subset_{n}B$.
(2) If $A\subset_{n}B$, $A\subset_{n}A\cup B$.
(3) If $A\subset_{n}B$, $A\cap B\subset_{n}A\cup B$.
Proof Since $A\subset_{n}B$, then $A^{+}(x)\leqslant B^{-}(x)$ for all $x\in U$, then (1), (2) and (3) are obvious.$\blacksquare$
Example 2.19.
Let $U=\{x\}$, $A=\frac{\{0.1,0.3,0.5\}}{x}$, $B=\frac{\{0.2,0.4,0.6\}}{x}$.
$(A\cap B)(x)=\{0.1,0.2,0.3,0.4,0.5\}$, $(A\cup B)(x)=\{0.2,0.3,0.4,0.5,0.6\}$. $A\cap B\not\subset_{n}A$, $A\cap B\not\subset_{n}B$, $A\not\subset_{n}A\cup B$, $B\not\subset_{n}A\cup B$, $A\cap B\not\subset_{n}A\cup B$.
For three classical sets $\bar{A}$, $\bar{B}$ and $\bar{C}$, if $\bar{A}\sqsubset\bar{B}$, then $\bar{A}\sqcap\bar{C}\sqsubset\bar{B}\sqcap\bar{C}$ and $\bar{A}\sqcup\bar{C}\sqsubset\bar{B}\sqcup\bar{C}$. These rules in the cases of three hesitant fuzzy sets $A$, $B$ and $C$ are investigated, shown as Proposition 2.20.
Proposition 2.20.
The following statements hold for $x\in U$ and $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$, $A\cap C\subset_{p}B\cap C$.
(2) If $A\subset_{p}B$, $A\cup C\subset_{p}B\cup C$.
(3) If $A\subset_{a}B$, $A\cap C\subset_{a}B\cap C$.
(4) If $A\subset_{a}B$, $A\cup C\subset_{a}B\cup C$.
(5) If $A\subset_{s}B$, $A\cap C\subset_{a}B\cap C$.
(6) If $A\subset_{s}B$, $A\cup C\subset_{a}B\cup C$.
(7) If $A\subset_{t}B$, $A\cap C\subset_{p}B\cap C$.
(8) If $A\subset_{t}B$, $A\cup C\subset_{p}B\cup C$.
(9) If $A\subset_{n}B$, $A\cap C\subset_{a}B\cap C$.
(10) If $A\subset_{n}B$, $A\cup C\subset_{a}B\cup C$.
(11) If $A\subset_{s}B$, then at least one of $(A\cup C)(x)\subset_{s}(B\cup C)(x)$ and $(A\cup C)(x)\subset_{t}(B\cup C)(x)$ holds.
(12) If $A\subset_{t}B$, then at least one of $(A\cup C)(x)\subset_{s}(B\cup C)(x)$ and $(A\cup C)(x)\subset_{t}(B\cup C)(x)$ holds.
(13) If $A\subset_{n}B$, then at least one of $(A\cup C)(x)\subset_{s}(B\cup C)(x)$ and $(A\cup C)(x)\subset_{t}(B\cup C)(x)$ holds.
Proof (1)-(10) are obvious.
(11) For a random $x\in U$, suppose $A(x)=\{o_{1},o_{2},\cdots,o_{a}\}=O$, $B(x)=\{v_{1},v_{2},\cdots,v_{b}\}=V$ and $C(x)=\{w_{1},w_{2},\cdots,w_{c}\}=W$, and $O$, $V$ and $W$ are descending sequences. Since $A\subset_{s}B$, then $b\leqslant a$, and $o_{i}\leqslant v_{i}$ for $1\leqslant i\leqslant b$.
To sort and investigate the two sequences $W^{1}=\{o_{1},o_{2},\cdots,o_{b}\}\sqcup W$ and $W^{2}=\{v_{1},v_{2},\cdots,v_{b}\}\sqcup W$, it is obvious that $W^{1}\preccurlyeq W^{2}$.
$v_{b}=B^{-}(x)\leqslant(B\cup C)^{-}(x)$, then $(B\cup C)(x)$ is a subsequence of $W^{2}$. Let $q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$. $w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
By Remark 2.8 (3), one of $(A\cup C)(x)\subset_{s}(B\cup C)(x)$ and $(A\cup C)(x)\subset_{t}(B\cup C)(x)$ holds.
(12) For a random $x\in U$, suppose $A(x)=\{o_{1},o_{2},\cdots,o_{a}\}=O$, $B(x)=\{v_{1},v_{2},\cdots,v_{b}\}=V$ and $C(x)=\{w_{1},w_{2},\cdots,w_{c}\}=W$, and $O$, $V$ and $W$ are descending sequences. Since $A\subset_{t}B$, then $a<b$, and $o_{i}\leqslant v_{i}$ for $1\leqslant i\leqslant a$.
To sort and investigate the two sequences $W^{1}=\{o_{1},o_{2},\cdots,o_{a}\}\sqcup W$ and $W^{2}=\{v_{1},v_{2},\cdots,v_{a}\}\sqcup W$, it is obvious that $W^{1}\preccurlyeq W^{2}$.
$o_{a}=A^{-}(x)\leqslant(A\cup C)^{-}(x)$, then $(A\cup C)(x)$ is a subsequence of $W^{1}$. Let $q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$. $w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
By Remark 2.8 (3), one of $(A\cup C)(x)\subset_{s}(B\cup C)(x)$ and $(A\cup C)(x)\subset_{t}(B\cup C)(x)$ holds.
(13) By Proposition 2.9 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for $x\in U$. By the results of (11) and (12), then (13) holds.$\blacksquare$
(11) and (12) of Proposition 2.20 can be combined as (1) of Proposition ${}^{\prime}$. (13) of Proposition 2.20 can be written into (2) of Proposition ${}^{\prime}$.
Proposition ${}^{\prime}$ The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{sot}B$, then $A\cup C\subset_{sot}B\cup C$.
(2) If $A\subset_{n}B$, then $A\cup C\subset_{sot}B\cup C$.
Example 2.21.
Let $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7},x_{8}\}$, $A=\frac{\{0.2,0.4\}}{x_{1}}+\frac{\{0.2,0.5\}}{x_{2}}+\frac{\{0.3,0.5\}}{x_{3}}+\frac{\{0.5,0.6\}}{x_{4}}+\frac{\{0.6,0.7\}}{x_{5}}+\frac{\{0.3,0.4\}}{x_{6}}+\frac{\{0.3,0.4\}}{x_{7}}+\frac{\{0.3,0.4\}}{x_{8}}$, $B=\frac{\{0.1,0.1,0.5\}}{x_{1}}+\frac{\{0.1,0.8\}}{x_{2}}+\frac{\{0.4,0.41\}}{x_{3}}+\frac{\{0.6,0.7\}}{x_{4}}+\frac{\{0.8,0.9\}}{x_{5}}+\frac{\{0.1,0.1,0.3,0.5\}}{x_{6}}+\frac{\{0.1,0.1,0.3,0.5\}}{x_{7}}+\frac{\{0.5,0.8\}}{x_{8}}$, $C=\frac{\{0.1,0.2\}}{x_{1}}+\frac{\{0.5\}}{x_{2}}+\frac{\{0.45,0.45\}}{x_{3}}+\frac{\{0.7,0.8\}}{x_{4}}+\frac{\{0.1,0.7\}}{x_{5}}+\frac{\{0.3,0.4\}}{x_{6}}+\frac{\{0.1,0.4\}}{x_{7}}+\frac{\{0.8,0.9\}}{x_{8}}$.
(1) $A(x_{1})\subset_{p}B(x_{1})$, $(A\cap C)(x_{1})=\{0.1,0.2,0.2\}$, $(B\cap C)(x_{1})=\{0.1,0.1,0.1,0.2\}$. $(A\cap C)(x_{1})\not\subset_{m}(B\cap C)(x_{1})$. $(A\cap C)(x_{1})\not\subset_{t}(B\cap C)(x_{1})$.
$A(x_{2})\subset_{p}B(x_{2})$, $(A\cap C)(x_{2})=\{0.2,0.5,0.5\}$, $(B\cap C)(x_{2})=\{0.1,0.5\}$, $(A\cap C)(x_{2})\not\subset_{a}(B\cap C)(x_{2})$.
$A(x_{1})\subset_{p}B(x_{1})$, $(A\cup C)(x_{1})=\{0.2,0.2,0.4\}$, $(B\cup C)(x_{1})=\{0.1,0.1,0.1,0.2,0.5\}$. $(A\cup C)(x_{1})\not\subset_{m}(B\cup C)(x_{1})$. $(A\cup C)(x_{1})\not\subset_{a}(B\cup C)(x_{1})$. $(A\cup C)(x_{1})\not\subset_{t}(B\cup C)(x_{1})$.
(2) $A(x_{5})\subset_{a}B(x_{5})$, $(A\cap C)(x_{5})\not\subset_{m}(B\cap C)(x_{5})$. $(A\cap C)(x_{5})\not\subset_{t}(B\cap C)(x_{5})$.
$A(x_{8})\subset_{a}B(x_{8})$, $(A\cup C)(x_{8})\not\subset_{m}(B\cup C)(x_{8})$.
$A(x_{5})\subset_{a}B(x_{5})$, $(A\cup C)(x_{5})\not\subset_{t}(B\cup C)(x_{5})$.
(3) $A(x_{3})\subset_{m}B(x_{3})$, $(A\cap C)(x_{3})=\{0.3,0.45,0.45\}$, $(B\cap C)(x_{3})=\{0.4,0.41\}$. $(A\cap C)(x_{3})\not\subset_{p}(B\cap C)(x_{3})$.
$A(x_{2})\subset_{m}B(x_{2})$, $(A\cap C)(x_{2})=\{0.2,0.5,0.5\}$, $(B\cap C)(x_{2})=\{0.1,0.5\}$. $(A\cap C)(x_{2})\not\subset_{m}(B\cap C)(x_{2})$.
$A(x_{3})\subset_{m}B(x_{3})$, $(A\cup C)(x_{3})=\{0.45,0.45,0.5\}$, $(B\cup C)(x_{3})=\{0.45,0.45\}$. $(A\cup C)(x_{3})\not\subset_{p}(B\cup C)(x_{3})$.
$A(x_{4})\subset_{m}B(x_{4})$, $(A\cup C)(x_{4})=\{0.7,0.8\}$, $(B\cup C)(x_{4})=\{0.7,0.7,0.8\}$. $(A\cup C)(x_{4})\not\subset_{m}(B\cup C)(x_{4})$.
(4) $A(x_{5})\subset_{s}B(x_{5})$, $(A\cap C)(x_{5})=\{0.1,0.6,0.7,0.7\}$, $(B\cap C)(x_{5})=\{0.1,0.7\}$. $(A\cap C)(x_{5})\not\subset_{m}(B\cap C)(x_{5})$. $(A\cap C)(x_{5})\not\subset_{t}(B\cap C)(x_{5})$.
$A(x_{4})\subset_{s}B(x_{4})$, $(A\cup C)(x_{4})\not\subset_{m}(B\cup C)(x_{4})$.
$A(x_{5})\subset_{s}B(x_{5})$, $(A\cup C)(x_{5})\not\subset_{t}(B\cup C)(x_{5})$.
(5) $A(x_{6})\subset_{t}B(x_{6})$, $(A\cap C)(x_{6})=\{0.3,0.3,0.4,0.4\}$, $(B\cap C)(x_{6})=\{0.1,0.1,0.3,0.3,0.4\}$. $(A\cap C)(x_{6})\not\subset_{m}(B\cap C)(x_{6})$. $(A\cap C)(x_{6})\not\subset_{a}(B\cap C)(x_{6})$. $(A\cap C)(x_{6})\not\subset_{t}(B\cap C)(x_{6})$.
$A(x_{7})\subset_{t}B(x_{7})$, $(A\cup C)(x_{7})=\{0.3,0.4,0.4\}$, $(B\cup C)(x_{7})=\{0.1,0.1,0.1,0.3,0.4,0.5\}$. $(A\cup C)(x_{7})\not\subset_{m}(B\cup C)(x_{7})$. $(A\cup C)(x_{7})\not\subset_{a}(B\cup C)(x_{7})$.
$A(x_{6})\subset_{t}B(x_{6})$, $(A\cup C)(x_{6})\not\subset_{t}(B\cup C)(x_{6})$.
(6) $A(x_{5})\subset_{n}B(x_{5})$, $(A\cap C)(x_{5})\not\subset_{m}(B\cap C)(x_{5})$. $(A\cap C)(x_{5})\not\subset_{t}(B\cap C)(x_{5})$.
$A(x_{8})\subset_{n}B(x_{8})$, $(A\cup C)(x_{8})=\{0.8,0.9\}$, $(B\cup C)(x_{8})=\{0.8,0.8,0.9\}$. $(A\cup C)(x_{8})\not\subset_{m}(B\cup C)(x_{8})$.
$A(x_{5})\subset_{n}B(x_{5})$, $(A\cup C)(x_{5})\not\subset_{t}(B\cup C)(x_{5})$.
For three classical sets $\bar{A}$, $\bar{B}$ and $\bar{C}$, $\bar{A}\sqsubset\bar{B}$ and $\bar{A}\sqsubset\bar{C}$ if and only if $\bar{A}\sqsubset\bar{B}\sqcap\bar{C}$. This rule in the cases of three hesitant fuzzy sets $A$, $B$ and $C$ is investigated, shown as Proposition 2.22.
Proposition 2.22.
The following statements hold for $A,B,C\in HF(U)$,
(1) $A\subset_{p}B$ and $A\subset_{p}C$ if and only if $A\subset_{p}B\cap C$.
(2) $A\subset_{a}B$ and $A\subset_{a}C$ if and only if $A\subset_{a}B\cap C$.
(3) $A\subset_{t}B$ and $A\subset_{t}C$, then $A\subset_{t}B\cap C$.
(4) $A\subset_{n}B$ and $A\subset_{n}C$ if and only if $A\subset_{n}B\cap C$.
Proof (1) For all $x\in U$, $A^{+}(x)\leqslant B^{+}(x)$ and $A^{+}(x)\leqslant C^{+}(x)$ if and only if $A^{+}(x)\leqslant\min\{B^{+}(x),C^{+}(x)\}$.
(2) For all $x\in U$, $A^{-}(x)\leqslant B^{-}(x)$ and $A^{-}(x)\leqslant C^{-}(x)$ if and only if $A^{-}(x)\leqslant\min\{B^{-}(x),C^{-}(x)\}$. To combine the result of (1), then (2) holds.
(3) For a random $x\in U$, the relationships of $h\in C(x)$ and interval $[B^{-}(x),B^{+}(x)]$ are divided into two cases (i) and (ii) as follows,
(i) $h\not\in[B^{-}(x),B^{+}(x)]$ for all $h\in C(x)$.
This case includes two sub-cases, $B^{+}(x)<C^{-}(x)$ and $C^{+}(x)<B^{-}(x)$. When $B^{+}(x)<C^{-}(x)$, $A(x)\subset_{t}B(x)=(B\cap C)(x)$. When $C^{+}(x)<B^{-}(x)$, $A(x)\subset_{t}C(x)=(B\cap C)(x)$.
(ii) $h\in[B^{-}(x),B^{+}(x)]$ for some $h\in C(x)$.
This case includes two sub-cases, $B^{+}(x)\leqslant C^{+}(x)$ and $C^{+}(x)<B^{+}(x)$.
When $B^{+}(x)\leqslant C^{+}(x)$, $h\in(B\cap C)(x)$ for all $h\in B(x)$, i.e., $B(x)\sqsubset(B\cap C)(x)$. Since $A(x)\subset_{t}B(x)$, by Remark 2.8 (1), there is a subsequence of $B(x)$, denoted as $\overline{B(x)}$, satisfying $A(x)\preccurlyeq\overline{B(x)}$. Since $B(x)\sqsubset(B\cap C)(x)$, thus $\overline{B(x)}$ is a subsequence of $(B\cap C)(x)$, then $A(x)\subset_{t}(B\cap C)(x)$.
For another sub-case $C^{+}(x)<B^{+}(x)$, $h\in(B\cap C)(x)$ for all $h\in C(x)$, i.e., $C(x)\sqsubset(B\cap C)(x)$. Since $A(x)\subset_{t}C(x)$, by Remark 2.8 (1), there is a subsequence of $C(x)$, denoted as $\overline{C(x)}$, satisfying $A(x)\preccurlyeq\overline{C(x)}$. Since $C(x)\sqsubset(B\cap C)(x)$, thus $\overline{C(x)}$ is a subsequence of $(B\cap C)(x)$, then $A(x)\subset_{t}(B\cap C)(x)$.
To sum up, $A(x)\subset_{t}(B\cap C)(x)$ for random $x\in U$, then $A\subset_{t}B\cap C$.
(4) For all $x\in U$, $A^{+}(x)\leqslant B^{-}(x)$ and $A^{+}(x)\leqslant C^{-}(x)$ if and only if $A^{+}(x)\leqslant\min\{B^{-}(x),C^{-}(x)\}$$\blacksquare$
Suppose $A(x)=\{0.3,0.5,0.7\}$, $B(x)=\{0.8,0.9\}$, $C(x)=\{0.6,0.8,0.9\}$. $A(x)\subset_{t}(B\cap C)(x)$, but $A(x)\not\subset_{t}B(x)$. The converse of Proposition 2.22 (3) does not hold.
Definition 2.23.
TV For $x\in U$ and $H\in HF(U)$, the complement of $H$ is denoted as $H^{c}$, where
$H^{c}(x)=\sqcup_{\gamma\in H(x)}\{1-\gamma\}$.
Theorem 2.24.
TV ; XX The following statements hold for $A,B,C\in HF(U)$,
(1) $(A^{c})^{c}=A$.
(2) $(A\cap B)^{c}=A^{c}\cup B^{c}$.
(3) $(A\cup B)^{c}=A^{c}\cap B^{c}$.
(4) $A\cap B=B\cap A$, $A\cup B=B\cup A$.
(5) $(A\cap B)\cap C=A\cap(B\cap C)$, $(A\cup B)\cup C=A\cup(B\cup C)$.
Proposition 2.25.
The following statements hold for $A,B\in HF(U)$,
(1) $A\subset_{a}B$, then $B^{c}\subset_{a}A^{c}$.
(2) $A\subset_{m}B$, then $B^{c}\subset_{m}A^{c}$.
(3) $A\subset_{s}B$, then $B^{c}\subset_{sot}A^{c}$.
(4) $A\subset_{n}B$, then $B^{c}\subset_{n}A^{c}$.
Proof (1) For all $x\in U$, since $A\subset_{a}B$, then $A^{+}(x)\leqslant B^{+}(x)$ and $A^{-}(x)\leqslant B^{-}(x)$. Then $(B^{c})^{-}(x)=1-B^{+}(x)\leqslant 1-A^{+}(x)=(A^{c})^{-}(x)$ and $(B^{c})^{+}(x)=1-B^{-}(x)\leqslant 1-A^{-}(x)=(A^{c})^{+}(x)$. Then $B^{c}\subset_{a}A^{c}$.
(2) For each $x\in U$, and each $H\in HF(U)$, $mean[H^{c}(x)]=(|H(x)|\times 1-|H(x)|\times mean[H(x)])/|H^{c}(x)|$, since $|H(x)|=|H^{c}(x)|$, then $mean[H^{c}(x)]=1-mean[H(x)]$. $A\subset_{m}B$, thus $mean[A(x)]\leqslant mean[B(x)]$, then $1-mean[B(x)]\leqslant 1-mean[A(x)]$, it implies $B^{c}(x)\subset_{m}A^{c}(x)$.
(3) For each $x\in U$, we can sort the elements of $A(x)$ and $B(x)$ at first. Let $A(x)=\{h_{1},h_{2},\cdots,h_{l}\}$ and $B(x)=\{h^{\prime}_{1},h^{\prime}_{2},\cdots,h^{\prime}_{k}\}$ be two descending sequences.
Since $A\subset_{s}B$, then $h^{\prime}_{1}\geqslant h_{1}$, $h^{\prime}_{2}\geqslant h_{2}$, $\cdots$, $h^{\prime}_{k}\geqslant h_{k}$ and $k\leqslant l$. Then, $1-h^{\prime}_{1}\leqslant 1-h_{1}$, $1-h^{\prime}_{2}\leqslant 1-h_{2}$, $\cdots$, $1-h^{\prime}_{k}\leqslant 1-h_{k}$. There is a subsequence of $A^{c}(x)$, denoted as $\overline{A^{c}(x)}$, satisfying $B^{c}(x)\preccurlyeq\overline{A^{c}(x)}$.
Since $|A^{c}(x)|=|A(x)|\geqslant|B(x)|=|B^{c}(x)|$, by Remark 2.8 (1) and (2), then $B^{c}(x)\subset_{sot}A^{c}(x)$.
(4) For all $x\in U$, since $A\subset_{n}B$, then $A^{+}(x)\leqslant B^{-}(x)$. $(A^{c})^{-}(x)=1-A^{+}(x)\geqslant 1-B^{-}(x)=(B^{c})^{+}(x)$, then $B^{c}(x)\subset_{n}A^{c}(x)$$\blacksquare$
Example 2.26.
Let $U=\{x,y\}$, $A=\frac{0.4,0.4}{x}+\frac{0.2,0.25}{y}$, $B=\frac{0.1,0.1,0.41}{x}+\frac{0.1,0.2,0.3}{y}$.
(1) $A(x)\subset_{p}B(x)$. $A^{c}(x)=\{0.6,0.6\}$, $B^{c}(x)=\{0.9,0.9,0.59\}$. $B^{c}(x)\not\subset_{p}A^{c}(x)$, $B^{c}(x)\not\subset_{m}A^{c}(x)$.
(2) $A(y)\subset_{t}B(y)$. $A^{c}(y)=\{0.8,0.75\}$, $B^{c}(y)=\{0.9,0.8,0.7\}$. $B^{c}(y)\not\subset_{p}A^{c}(y)$, $B^{c}(y)\not\subset_{a}A^{c}(y)$, $B^{c}(y)\not\subset_{m}A^{c}(y)$, $B^{c}(y)\not\subset_{s}A^{c}(y)$, $B^{c}(y)\not\subset_{t}A^{c}(y)$, $B^{c}(y)\not\subset_{n}A^{c}(y)$.
It should be noted that
$A\subset_{p}B\nRightarrow B^{c}\subset_{p}A^{c}.$
2.3 Equivalence relationships of hesitant fuzzy sets
Theorem 2.27.
The following statements hold for $x\in U$ and $A,B\in HF(U)$,
(1) $A=B$, then $A=_{p}B$, $A=_{a}B$ and $A=_{m}B$.
(2) $A=_{s}B$Leftrightarrow $A=B$.
(3) $A=_{s}B$, then $A=_{p}B$, $A=_{a}B$ and $A=_{m}B$.
(4) If $A=_{n}B$, then $h^{\prime}=h^{\prime\prime}$ for all $h^{\prime}\in A(x)$ and all $h^{\prime\prime}\in B(x)$.
(5) If $A=_{n}B$, then $A=_{p}B$, $A=_{a}B$ and $A=_{m}B$.
Proof (1) Obvious.
(2) It is obvious that $A=B$ implies $A=_{s}B$.
On the other hand, for each $x\in U$, we can sort $A(x)$ and $B(x)$ to obtain two descending sequences $A(x)=\{h_{1},h_{2},\cdots,h_{l}\}$ and $B(x)=\{g_{1},g_{2},\cdots,g_{k}\}$. Since $A=_{s}B$, i.e., $A(x)\subset_{s}B(x)$ and $B(x)\subset_{s}A(x)$. Then $|A(x)|=|B(x)|$, $h_{i}\leqslant g_{i}$ and $g_{i}\leqslant h_{i}$ for $1\leqslant i\leqslant|A(x)|$. It means that $A=B$.
(3) By (2) and (1), then (3) is obvious.
(4) For each $x\in U$, if $A=_{n}B$, i.e., $A(x)\subset_{n}B(x)$ and $B(x)\subset_{n}A(x)$, then $B^{+}(x)\geqslant B^{-}(x)\geqslant A^{+}(x)\geqslant A^{-}(x)\geqslant B^{+}(x)$, then $h^{\prime}=h^{\prime\prime}$ for all $h^{\prime}\in A(x)$ and all $h^{\prime\prime}\in B(x)$.
(5) By (4), (5) is obvious.$\blacksquare$
Theorem 2.27 indicates that if one of $A=_{p}B$, $A=_{a}B$ and $A=_{m}B$ does not hold then $A=_{n}B$ and $A=B$ ($A=_{s}B$) do not hold.
Theorem 2.28.
The following statements hold for $A,B,C\in HF(U)$,
(1) $A\cap A=_{p}A$, $A\cup A=_{p}A$.
(2) $A\cap A=_{a}A$, $A\cup A=_{a}A$.
(3) $A\cap A=_{m}A$, $A\cup A=_{m}A$.
(4) $(A\cup B)\cap A=_{p}A$, $(A\cap B)\cup A=_{p}A$.
(5) $(A\cup B)\cap A=_{a}A$, $(A\cap B)\cup A=_{a}A$.
(6) $(A\cup B)\cap C=_{p}(C\cap A)\cup(C\cap B)$, $(A\cap B)\cup C=_{p}(C\cup A)\cap(C\cup B)$.
(7) $(A\cup B)\cap C=_{a}(C\cap A)\cup(C\cap B)$, $(A\cap B)\cup C=_{a}(C\cup A)\cap(C\cup B)$.
Proof (1)-(3) are obvious.
Let $A^{\prime}$, $A^{\prime\prime}$, $B^{\prime}$, $B^{\prime\prime}$, $C^{\prime}$ and $C^{\prime\prime}$ be six fuzzy sets. For all $x\in U$, let $A^{\prime}(x)=A^{+}(x)$, $A^{\prime\prime}(x)=A^{-}(x)$, $B^{\prime}(x)=B^{+}(x)$, $B^{\prime\prime}(x)=B^{-}(x)$, $C^{\prime}(x)=C^{+}(x)$ and $C^{\prime\prime}(x)=C^{-}(x)$.
For the proofs of (4)-(7), we give a simple example as follows. By the results of fuzzy sets ZADE , $(A^{\prime}(x)\cup B^{\prime}(x))\cap A^{\prime}(x)=A^{\prime}(x)$ and $(A^{\prime\prime}(x)\cup B^{\prime\prime}(x))\cap A^{\prime\prime}(x)=A^{\prime\prime}(x)$, i.e., $(A^{+}(x)\cup B^{+}(x))\cap A^{+}(x)=A^{+}(x)$ and $(A^{-}(x)\cup B^{-}(x))\cap A^{-}(x)=A^{-}(x)$, then $(A\cup B)\cap A=_{a}A$.
(4)-(7) are easy to be proved by the results of fuzzy sets in ZADE $\blacksquare$
Example 2.29.
Let $U=\{x\}$, $A=\frac{0.1,0.2,0.3}{x}$, $B=\frac{0.3,0.4,0.5}{x}$, $C=\frac{0.3,0.45,0.5}{x}$.
$(A\cap B)(x)=\{0.1,0.2,0.3,0.3\}$, $((A\cap B)\cup A)(x)=\{0.1,0,1,0.2,0.2,0.3,0.3,0.3\}\neq_{m}A(x)$.
$(A\cup B)(x)=\{0.3,0.3,0.4,0.5\}$, $((A\cup B)\cap A)(x)=\{0.1,0.2,0.3,0.3,0.3\}\neq_{m}A(x)$.
$((A\cup B)\cap C)(x)=\{0.3,0.3,0.3,0.4,0.45,0.5\}$, $((A\cap C)\cup(B\cap C))(x)=\{0.3,0.3,0.3,0.3,0.4,0.45,0.5\}$, $((A\cup B)\cap C)(x)\neq_{m}((A\cap C)\cup(B\cap C))(x)$.
$((A\cap B)\cup C)(x)=\{0.3,0.3,0.3,0.45,0.5\}$, $((A\cup C)\cap(B\cup C))(x)=\{0.3,0.3,0.3,0.3,0.4,0.45,0.45,0.5,0.5\}$, $((A\cap B)\cup C)(x)\neq_{m}((A\cup C)\cap(B\cup C))(x)$.
In addition, $0.4\not\in((A\cap B)\cup C)(x)$, however, $0.4\in((A\cup C)\cap(B\cup C))(x)$.
Inspired by Theorem 2.27 and Example 2.29, it should be noted that
$\begin{cases}(A\cap B)\cup A\not=A,\\ (A\cup B)\cap A\not=A,\\ (A\cup B)\cap C\not=(A\cap C)\cup(B\cap C),\\ (A\cap B)\cup C\not=(A\cup C)\cap(B\cup C),\\ (A\cap B)\cup A\neq_{n}A,\\ (A\cup B)\cap A\neq_{n}A,\\ (A\cup B)\cap C\neq_{n}(A\cap C)\cup(B\cap C),\\ (A\cap B)\cup C\neq_{n}(A\cup C)\cap(B\cup C).\\ \end{cases}$
For the $\subset_{p}$ and $=_{p}$ relationships, the upper bound with respect to an object $x\in U$ and a hesitant fuzzy set over $U$ is the unique thing that needs to be investigated. Hence, many inclusion relationships and equivalence relationships of fuzzy sets also hold on hesitant fuzzy sets as the forms of $\subset_{p}$ and $=_{p}$ relationships. However, $A\subset_{p}B\nRightarrow B^{c}\subset_{p}A^{c}$, then the $\subset_{p}$ and $=_{p}$ relationships are not consistent with the inclusion relationships and equivalence relationships of fuzzy sets.
Section 2.2 and 2.3 reveal that some common inclusion relationships and equivalence relationships do not hold on hesitant fuzzy sets but hold on classical sets.
3 Foundations of families of hesitant fuzzy sets
Let $\mathscr{H}_{1}$ and $\mathscr{H}_{2}$ be two families of hesitant fuzzy sets. If $H_{i}\in\mathscr{H}_{2}$ holds for all $H_{i}\in\mathscr{H}_{1}$, we say that $\mathscr{H}_{1}$ is the classical subset of $\mathscr{H}_{2}$, denoted as $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$.
Theorem 3.30.
Let $\mathscr{H}\sqsubset HF(U)$ be a family of hesitant fuzzy sets over $U$. The following statements hold,
(1) $\bigcap\{H:H\in\mathscr{H}\}\subset_{p}\bigcup\{H:H\in\mathscr{H}\}$.
(2) $\bigcap\{H:H\in\mathscr{H}\}\subset_{a}\bigcup\{H:H\in\mathscr{H}\}$.
(3) $\bigcap\{H:H\in\mathscr{H}\}\subset_{m}\bigcup\{H:H\in\mathscr{H}\}$.
(4) $\bigcap\{H:H\in\mathscr{H}\}\subset_{sot}\bigcup\{H:H\in\mathscr{H}\}$.
Proof (1) For each $x\in U$, $inf\{H^{+}(x):H\in\mathscr{H}\}\leqslant sup\{H^{+}(x):H\in\mathscr{H}\}$.
(2) For each $x\in U$, $inf\{H^{-}(x):H\in\mathscr{H}\}\leqslant sup\{H^{-}(x):H\in\mathscr{H}\}$. To combine the result of (1), then (2) holds.
(3) For each $x\in U$, let $H_{inf}^{+}(x)=inf\{H^{+}(x):H\in\mathscr{H}\}$, $H_{sup}^{-}(x)=sup\{H^{-}(x):H\in\mathscr{H}\}$.
If $H_{inf}^{+}(x)\leqslant H_{sup}^{-}(x)$, ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{m}(\bigcup\{H:H\in\mathscr{H}\})(x)$ is obvious.
If $H_{sup}^{-}(x)<H_{inf}^{+}(x)$, we construct a set of numbers $V=\{h:h\in H(x),H\in\mathscr{H},H_{sup}^{-}(x)\leqslant h\leqslant H_{inf}^{+}(x)\}$.
If ($(\bigcap\{H:H\in\mathscr{H}\})(x)-V\neq\emptyset$, then $w<inf(V)$ for all $w\in(\bigcap\{H:H\in\mathscr{H}\})(x)-V$. If ($(\bigcup\{H:H\in\mathscr{H}\})(x)-V\neq\emptyset$, then $w^{\prime}>sup(V)$ for all $w^{\prime}\in(\bigcup\{H:H\in\mathscr{H}\})(x)-V$. Then, $mean[(\bigcap\{H:H\in\mathscr{H}\})(x)]\leqslant mean[V]\leqslant mean[(\bigcup\{H:H\in\mathscr{H}\})(x)]$, i.e., ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{m}(\bigcup\{H:H\in\mathscr{H}\})(x)$.
(4) For each $x\in U$, $H_{inf}^{+}(x)$ and $H_{sup}^{-}(x)$ are constructed as (2) above.
If $H_{inf}^{+}(x)\leqslant H_{sup}^{-}(x)$, by Proposition 2.9 (9), ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}\})(x)$ is obvious.
If $H_{sup}^{-}(x)<H_{inf}^{+}(x)$, let $V=(\bigcap\{H:H\in\mathscr{H}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}\})(x)$, $min((\bigcup\{H:H\in\mathscr{H}\})(x)-V)>H_{inf}^{+}(x)>H_{sup}^{-}(x)>max((\bigcap\{H:H\in\mathscr{H}\})(x)-V)$, by Remark 2.8 (4), ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}\})(x)$$\blacksquare$
Theorem 3.31.
Let $A\in HF(U)$ and $\mathscr{H}\sqsubset HF(U)$ be a family of hesitant fuzzy sets over $U$. The following statements hold,
(1) $A\subset_{p}H$ for all $H\in\mathscr{H}$ if and only if $A\subset_{p}\bigcap\{H:H\in\mathscr{H}\}$.
(2) If $A\subset_{p}H_{\alpha}$ for a $H_{\alpha}\in\mathscr{H}$, then $A\subset_{p}\bigcup\{H:H\in\mathscr{H}\}$.
(3) $A\subset_{a}H$ for all $H\in\mathscr{H}$ if and only if $A\subset_{a}\bigcap\{H:H\in\mathscr{H}\}$.
(4) If $A\subset_{a}H_{\alpha}$ for a $H_{\alpha}\in\mathscr{H}$, then $A\subset_{a}\bigcup\{H:H\in\mathscr{H}\}$.
(5) If $A\subset_{t}H$ for all $H\in\mathscr{H}$, then $A\subset_{t}\bigcap\{H:H\in\mathscr{H}\}$.
(6) If $A\subset_{t}H_{\alpha}$ for a $H_{\alpha}\in\mathscr{H}$, and $|A(x)|<|H(x)|$ for all $H\in\mathscr{H}$ and $x\in U$, then $A\subset_{t}\bigcup\{H:H\in\mathscr{H}\}$.
(7) $A\subset_{n}H$ for all $H\in\mathscr{H}$ if and only if $A\subset_{n}\bigcap\{H:H\in\mathscr{H}\}$.
(8) If $A\subset_{n}H_{\alpha}$ for a $H_{\alpha}\in\mathscr{H}$, then $A\subset_{n}\bigcup\{H:H\in\mathscr{H}\}$.
Proof (1) For each $x\in U$, $A^{+}(x)\leqslant inf\{H^{+}(x):H\in\mathscr{H}\}$ if and only if $A^{+}(x)\leqslant H^{+}(x)$ for all $H\in\mathscr{H}$.
(2) $A\subset_{p}H_{\alpha}\in\mathscr{H}$, $A^{+}(x)\leqslant H_{\alpha}^{+}(x)\leqslant sup\{H^{+}(x):H\in\mathscr{H}\}$ for all $x\in U$.
(3) For each $x\in U$, $A^{-}(x)\leqslant inf\{H^{-}(x):H\in\mathscr{H}\}$ if and only if $A^{-}(x)\leqslant H^{-}(x)$ for all $H\in\mathscr{H}$. To combine the result of (1), then (3) holds.
(4) $A\subset_{a}H_{\alpha}\in\mathscr{H}$, $A^{+}(x)\leqslant H_{\alpha}^{+}(x)\leqslant sup\{H^{+}(x):H\in\mathscr{H}\}$ and $A^{-}(x)\leqslant H_{\alpha}^{-}(x)\leqslant sup\{H^{-}(x):H\in\mathscr{H}\}$ for all $x\in U$.
(5) For each $x\in U$, $A\subset_{t}H$ for all $H\in\mathscr{H}$, then $|A(x)|<|H(x)|$ and $|A(x)|<|(\bigcap\{H:H\in\mathscr{H}\})(x)|$ are obvious.
$H_{\gamma}(x)$ can be sorted into a descending sequence $H_{\gamma}(x)=\{h_{\gamma}^{1},h_{\gamma}^{2},\cdots\}$ for each $H_{\gamma}\in\mathscr{H}$ and $\gamma\in\Gamma$, where $\Gamma$ is an order set and $|\Gamma|=|\mathscr{H}|$.
Let $\overline{H}=\{\overline{h}^{1},\overline{h}^{2},\cdots,\overline{h}^{|A(x)|}\}$ be a subsequence of ($(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H},{\gamma}\in\Gamma\})(x)$, where $\overline{h}^{j}=inf\{h_{\gamma}^{j}:h_{\gamma}^{j}\in H_{\gamma}(x),\gamma\in\Gamma\}$ for $1\leqslant j\leqslant|A(x)|$.
$A(x)$ can be sorted into a descending sequence $A(x)=\{v_{1},v_{2},\cdots,v_{|A(x)|}\}$. Since $h_{\gamma}^{j}\geqslant v_{j}$ for all $\gamma\in\Gamma$ and $1\leqslant j\leqslant|A(x)|$, then $\overline{h}^{j}=inf\{h_{\gamma}^{j}:h_{\gamma}^{j}\in H_{\gamma}(x),\gamma\in\Gamma\}\geqslant v_{j}$ for $1\leqslant j\leqslant|A(x)|$, i.e., $A(x)\preccurlyeq\overline{H}$.
$|A(x)|<|(\bigcap\{H:H\in\mathscr{H}\})(x)|$, by Remark 2.8 (1), then $A(x)\subset_{t}(\bigcap\{H:H\in\mathscr{H}\})(x)$.
(6) $|A(x)|<|H(x)|$ for all $H\in\mathscr{H}$ and $x\in U$. Let $H_{*}=\{h_{*}^{1},h_{*}^{2},\cdots,h_{*}^{|A(x)|}\}$ be a subsequence of ($(\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H},{\gamma}\in\Gamma\})(x)$, where $h_{*}^{j}=sup\{h_{\gamma}^{j}:h_{\gamma}^{j}\in H_{\gamma}(x),\gamma\in\Gamma\}$ for $1\leqslant j\leqslant|A(x)|$.
For two descending sequences $A(x)=\{v_{1},v_{2},\cdots,v_{|A(x)|}\}$ and $H_{\alpha}(x)=\{h_{\alpha}^{1},h_{\alpha}^{2},\cdots\}$, since $A\subset_{t}H_{\alpha}$, then $v_{j}\leqslant h_{\alpha}^{j}\leqslant h_{*}^{j}=sup\{h_{\gamma}^{j}:h_{\gamma}^{j}\in H_{\gamma}(x),\gamma\in\Gamma\}$ for $1\leqslant j\leqslant|A(x)|$, i.e., $A(x)\preccurlyeq H_{*}$.
$|A(x)|<|(\bigcup\{H:H\in\mathscr{H}\})(x)|$, by Remark 2.8 (1), then $A(x)\subset_{t}(\bigcup\{H:H\in\mathscr{H}\})(x)$.
(7) For each $x\in U$, $A^{+}(x)\leqslant inf\{H^{-}(x):H\in\mathscr{H}\}$ if and only if $A^{+}(x)\leqslant H^{-}(x)$ for all $H\in\mathscr{H}$.
(8) $A\subset_{n}H_{\alpha}\in\mathscr{H}$, then $A^{+}(x)\leqslant H_{\alpha}^{-}(x)\leqslant sup\{H^{-}(x):H\in\mathscr{H}\}$ for all $x\in U$$\blacksquare$
Theorem 3.32.
Let $\mathscr{H}_{1},\mathscr{H}_{2}\sqsubset HF(U)$ be two families of hesitant fuzzy sets over $U$. The following statements hold,
(1) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\bigcap\{H:H\in\mathscr{H}_{2}\}\subset_{p}\bigcap\{H:H\in\mathscr{H}_{1}\}$.
(2) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\bigcup\{H:H\in\mathscr{H}_{1}\}\subset_{p}\bigcup\{H:H\in\mathscr{H}_{2}\}$.
(3) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\bigcap\{H:H\in\mathscr{H}_{1}\}\subset_{p}\bigcup\{H:H\in\mathscr{H}_{2}\}$.
(4) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\bigcap\{H:H\in\mathscr{H}_{2}\}\subset_{a}\bigcap\{H:H\in\mathscr{H}_{1}\}$.
(5) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\bigcup\{H:H\in\mathscr{H}_{1}\}\subset_{a}\bigcup\{H:H\in\mathscr{H}_{2}\}$.
(6) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\bigcap\{H:H\in\mathscr{H}_{1}\}\subset_{a}\bigcup\{H:H\in\mathscr{H}_{2}\}$.
(7) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\bigcup\{H:H\in\mathscr{H}_{1}\}\subset_{sot}\bigcup\{H:H\in\mathscr{H}_{2}\}$.
(8) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\bigcap\{H:H\in\mathscr{H}_{1}\}\subset_{sot}\bigcup\{H:H\in\mathscr{H}_{2}\}$.
Proof (1) For each $x\in U$, if $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\{H^{+}(x):H\in\mathscr{H}_{1}\}\sqsubset\{H^{+}(x):H\in\mathscr{H}_{2}\}$. $inf\{H^{+}(x):H\in\mathscr{H}_{1}\}\geqslant inf\{H^{+}(x):H\in\mathscr{H}_{2}\}$, then ($(\bigcap\{H:H\in\mathscr{H}_{2}\})(x)\subset_{p}(\bigcap\{H:H\in\mathscr{H}_{1}\})(x)$.
(2) For each $x\in U$, if $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\{H^{+}(x):H\in\mathscr{H}_{1}\}\sqsubset\{H^{+}(x):H\in\mathscr{H}_{2}\}$. $sup\{H^{+}(x):H\in\mathscr{H}_{1}\}\leqslant sup\{H^{+}(x):H\in\mathscr{H}_{2}\}$, then ($(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\subset_{p}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.
(3) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, by Theorem 3.30 (1) and Theorem 3.32 (2), $\bigcap\{H:H\in\mathscr{H}_{1}\}\subset_{p}\bigcup\{H:H\in\mathscr{H}_{1}\}\subset_{p}\bigcup\{H:H\in\mathscr{H}_{2}\}$.
(4) For each $x\in U$, if $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\{H^{-}(x):H\in\mathscr{H}_{1}\}\sqsubset\{H^{-}(x):H\in\mathscr{H}_{2}\}$. $inf\{H^{-}(x):H\in\mathscr{H}_{1}\}\geqslant inf\{H^{-}(x):H\in\mathscr{H}_{2}\}$. To combine the result of (1), then (4) holds.
(5) For each $x\in U$, if $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $\{H^{-}(x):H\in\mathscr{H}_{1}\}\sqsubset\{H^{-}(x):H\in\mathscr{H}_{2}\}$. $sup\{H^{-}(x):H\in\mathscr{H}_{1}\}\leqslant sup\{H^{-}(x):H\in\mathscr{H}_{2}\}$. To combine the result of (2), then (5) holds.
(6) If $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, by Theorem 3.30 (2) and Theorem 3.32 (5), $\bigcap\{H:H\in\mathscr{H}_{1}\}\subset_{a}\bigcup\{H:H\in\mathscr{H}_{1}\}\subset_{a}\bigcup\{H:H\in\mathscr{H}_{2}\}$.
(7) For each $x\in U$, let $sup_{1}^{-}=sup\{H^{-}(x):H\in\mathscr{H}_{1}\}=(\bigcup\{H:H\in\mathscr{H}_{1}\})^{-}(x)$, $sup_{2}^{-}=sup\{H^{-}(x):H\in\mathscr{H}_{2}\}=(\bigcup\{H:H\in\mathscr{H}_{2}\})^{-}(x)$.
Let $V=(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$. Since $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $sup_{1}^{-}\leqslant sup_{2}^{-}$. If $h\in(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)$ and $sup_{2}^{-}\leqslant h$, then $h\in(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$. It means that $max((\bigcup\{H:H\in\mathscr{H}_{1}\})(x)-V)<sup_{2}^{-}$. On the other hand, $sup_{2}^{-}<min((\bigcup\{H:H\in\mathscr{H}_{2}\})(x)-V)$, by Remark 2.8 (4), then ($(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.
(8) For each $x\in U$, let $inf_{1}^{+}=inf\{H^{+}(x):H\in\mathscr{H}_{1}\}$, $sup_{2}^{-}=sup\{H^{-}(x):H\in\mathscr{H}_{2}\}$.
If $inf_{1}^{+}\leqslant sup_{2}^{-}$, by Proposition 2.9 (9), ($(\bigcap\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$ is obvious.
If $sup_{2}^{-}<inf_{1}^{+}$. Let $V=(\bigcap\{H:H\in\mathscr{H}_{1}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.
Base on the condition $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, if $h\in(\bigcap\{H:H\in\mathscr{H}_{1}\})(x)$ and $sup_{2}^{-}\leqslant h$, then $h\in(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$. It means that $max((\bigcap\{H:H\in\mathscr{H}_{1}\})(x)-V)<sup_{2}^{-}$. On the other hand, $sup_{2}^{-}<min((\bigcup\{H:H\in\mathscr{H}_{2}\})(x)-V)$, by Remark 2.8 (4), then ($(\bigcap\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$$\blacksquare$
Theorem 3.33.
Let $\mathscr{H}_{1},\mathscr{H}_{2}\sqsubset HF(U)$ be two families of hesitant fuzzy sets over $U$. The following statements hold,
(1) There is a $H_{\gamma^{*}}\in\mathscr{H}_{2}$ satisfied $H_{\alpha}\subset_{p}H_{\gamma^{*}}$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{p}\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
(2) There is a $H_{\alpha^{*}}\in\mathscr{H}_{1}$ satisfied $H_{\alpha^{*}}\subset_{p}H_{\gamma}$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{p}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
(3) If $H_{\alpha}\subset_{p}H_{\gamma}$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{p}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
(4) There is a $H_{\gamma^{*}}\in\mathscr{H}_{2}$ satisfied $H_{\alpha}\subset_{a}H_{\gamma^{*}}$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{a}\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
(5) There is a $H_{\alpha^{*}}\in\mathscr{H}_{1}$ satisfied $H_{\alpha^{*}}\subset_{a}H_{\gamma}$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{a}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
(6) If $H_{\alpha}\subset_{a}H_{\gamma}$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{a}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
(7) There is a $H_{\gamma^{*}}\in\mathscr{H}_{2}$ satisfied $H_{\alpha}\subset_{n}H_{\gamma^{*}}$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{n}\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
(8) There is a $H_{\alpha^{*}}\in\mathscr{H}_{1}$ satisfied $H_{\alpha^{*}}\subset_{n}H_{\gamma}$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{n}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
(9) If $H_{\alpha}\subset_{n}H_{\gamma}$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{n}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
Proof (1) For each $x\in U$, there is $H_{\gamma^{*}}\in\mathscr{H}_{2}$ such that $H_{\alpha}^{+}(x)\leqslant H_{\gamma^{*}}^{+}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma^{*}}^{+}(x)\leqslant sup\{H_{\gamma}^{+}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. Then ($(\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{p}(\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
(2) For each $x\in U$, $H_{\alpha^{*}}^{+}(x)\leqslant H_{\gamma}^{+}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $inf\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\alpha^{*}}^{+}(x)\leqslant inf\{H_{\gamma}^{+}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. Then ($(\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{p}(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
(3) For each $x\in U$, $H_{\alpha}^{+}(x)\leqslant H_{\gamma}^{+}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, thus $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma}^{+}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant inf\{H_{\gamma}^{+}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. Then ($(\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{p}(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
(4) For each $x\in U$, there is $H_{\gamma^{*}}\in\mathscr{H}_{2}$ such that $H_{\alpha}^{-}(x)\leqslant H_{\gamma^{*}}^{-}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $sup\{H_{\alpha}^{-}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma^{*}}^{-}(x)\leqslant sup\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. To combine the result of (1), then (4) holds.
(5) For each $x\in U$, $H_{\alpha^{*}}^{-}(x)\leqslant H_{\gamma}^{-}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $inf\{H_{\alpha}^{-}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\alpha^{*}}^{-}(x)\leqslant inf\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. To combine the result of (2), then (5) holds.
(6) For each $x\in U$, $H_{\alpha}^{-}(x)\leqslant H_{\gamma}^{-}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, thus $sup\{H_{\alpha}^{-}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma}^{-}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $sup\{H_{\alpha}^{-}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant inf\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. To combine the result of (3), then (6) holds.
(7) For each $x\in U$, there is $H_{\gamma^{*}}\in\mathscr{H}_{2}$ such that $H_{\alpha}^{+}(x)\leqslant H_{\gamma^{*}}^{-}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma^{*}}^{-}(x)\leqslant sup\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. Then ($(\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{n}(\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
(8) For each $x\in U$, $H_{\alpha^{*}}^{+}(x)\leqslant H_{\gamma}^{-}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $inf\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\alpha^{*}}^{+}(x)\leqslant inf\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. Then ($(\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{n}(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
(9) For each $x\in U$, $H_{\alpha}^{+}(x)\leqslant H_{\gamma}^{-}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, thus $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma}^{-}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant inf\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$. Then ($(\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{n}(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$$\blacksquare$
4 Parameter reductions of hesitant fuzzy information systems
In this section, we investigate some propositions with regard to parameter reductions in hesitant fuzzy information systems, and then present an algorithm for parameter reductions of hesitant fuzzy information systems.
4.1 Parameter reductions based on $\beta$-partition in hesitant fuzzy information systems
In the classical information systems, if an object has the characteristic with respect to attribute $e$, then $F(e)(x)$ is denoted as 1, otherwise, $F(e)(x)$ is denoted as 0. Inspired by the idea of classical information systems, if $\beta(e)\subset_{p}F(e)(x)$ then $x$ is considered to have the characteristic with respect to attribute $e$ based on $p\beta$-partition in hesitant fuzzy information systems. $\beta=\sum\limits_{e_{i}\in E}\frac{\beta(e_{i})}{e_{i}}$ can be considered as an indicator to investigate whether object $x$ has characteristic $e_{i}\in E$, where $\beta(e_{i})$ is a hesitant fuzzy membership degree of $e_{i}$ in indicator $\beta$.
As shown in Example 4.44, $\beta(e_{1})\not\subset_{p}F(e_{1})(x)$ for $x\in\{x_{1},x_{2},x_{3}\}$, $\beta(e_{1})\subset_{p}F(e_{1})(y)$ for $y\in\{x_{4},x_{5},x_{6}\}$. A family of $U$ can be denoted as $N_{e_{1}}^{p\beta}=\{\{x:\beta(e_{1})\not\subset_{p}F(e_{1})(x),x\in U\},\{y:\beta(e_{1})\subset_{p}F(e_{1})(y),y\in U\}\}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$. $N_{e_{1}}^{p\beta}(x_{1})=N_{e_{1}}^{p\beta}(x_{2})=N_{e_{1}}^{p\beta}(x_{3})=\{x_{1},x_{2},x_{3}\}$, $N_{e_{1}}^{p\beta}(x_{4})=N_{e_{1}}^{p\beta}(x_{5})=N_{e_{1}}^{p\beta}(x_{6})=\{x_{4},x_{5},x_{6}\}$. $N_{e_{1}}^{p\beta}$ is a $p\beta$-partition of $U$ with respect to $e_{1}$.
Definition 4.34.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$.
(1) Let $N_{e}^{p\beta}=\{\{x:\beta(e)\subset_{p}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{p}F(e)(y),y\in U\}\}$ be a $p\beta$-partition of $U$ with respect to $e$, $N_{e}^{p\beta}(x)$ be an element of $N_{e}^{p\beta}$ which contains $x$, $N_{A}^{p\beta}(x)=\sqcap\{N_{e}^{p\beta}(x):e\in A\}$, and $Cov^{p\beta}(A)=\{N_{A}^{p\beta}(x):x\in U\}$.
(2) Let $N_{e}^{a\beta}=\{\{x:\beta(e)\subset_{a}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{a}F(e)(y),y\in U\}\}$ be a $a\beta$-partition of $U$ with respect to $e$, $N_{e}^{a\beta}(x)$ be an element of $N_{e}^{a\beta}$ which contains $x$, $N_{A}^{a\beta}(x)=\sqcap\{N_{e}^{a\beta}(x):e\in A\}$, and $Cov^{a\beta}(A)=\{N_{A}^{a\beta}(x):x\in U\}$.
(3) Let $N_{e}^{m\beta}=\{\{x:\beta(e)\subset_{m}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{m}F(e)(y),y\in U\}\}$ be a $m\beta$-partition of $U$ with respect to $e$, $N_{e}^{m\beta}(x)$ be an element of $N_{e}^{m\beta}$ which contains $x$, $N_{A}^{m\beta}(x)=\sqcap\{N_{e}^{m\beta}(x):e\in A\}$, and $Cov^{m\beta}(A)=\{N_{A}^{m\beta}(x):x\in U\}$.
(4) Let $N_{e}^{s\beta}=\{\{x:\beta(e)\subset_{s}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{s}F(e)(y),y\in U\}\}$ be a $s\beta$-partition of $U$ with respect to $e$, $N_{e}^{s\beta}(x)$ be an element of $N_{e}^{s\beta}$ which contains $x$, $N_{A}^{s\beta}(x)=\sqcap\{N_{e}^{s\beta}(x):e\in A\}$, and $Cov^{s\beta}(A)=\{N_{A}^{s\beta}(x):x\in U\}$.
(5) Let $N_{e}^{t\beta}=\{\{x:\beta(e)\subset_{t}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{t}F(e)(y),y\in U\}\}$ be a $t\beta$-partition of $U$ with respect to $e$, $N_{e}^{t\beta}(x)$ be an element of $N_{e}^{t\beta}$ which contains $x$, $N_{A}^{t\beta}(x)=\sqcap\{N_{e}^{t\beta}(x):e\in A\}$, and $Cov^{t\beta}(A)=\{N_{A}^{t\beta}(x):x\in U\}$.
(6) Let $N_{e}^{n\beta}=\{\{x:\beta(e)\subset_{n}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{n}F(e)(y),y\in U\}\}$ be a $n\beta$-partition of $U$ with respect to $e$, $N_{e}^{n\beta}(x)$ be an element of $N_{e}^{n\beta}$ which contains $x$, $N_{A}^{n\beta}(x)=\sqcap\{N_{e}^{n\beta}(x):e\in A\}$, and $Cov^{n\beta}(A)=\{N_{A}^{n\beta}(x):x\in U\}$.
Proposition 4.35.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $e\in E$. The following statements hold,
(1) Suppose $N_{e}^{p\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
(2) Suppose $N_{e}^{a\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
(3) Suppose $N_{e}^{m\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
(4) Suppose $N_{e}^{s\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
(5) Suppose $N_{e}^{t\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
(6) Suppose $N_{e}^{n\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
Proof (1)-(6) are obvious.
Proposition 4.36.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $e\in E$. The following statements hold,
(1) If $y\in N_{e}^{p\beta}(x)$, then $N_{e}^{p\beta}(y)=N_{e}^{p\beta}(x)$.
(2) If $y\in N_{e}^{a\beta}(x)$, then $N_{e}^{a\beta}(y)=N_{e}^{a\beta}(x)$.
(3) If $y\in N_{e}^{m\beta}(x)$, then $N_{e}^{m\beta}(y)=N_{e}^{m\beta}(x)$.
(4) If $y\in N_{e}^{s\beta}(x)$, then $N_{e}^{s\beta}(y)=N_{e}^{s\beta}(x)$.
(5) If $y\in N_{e}^{t\beta}(x)$, then $N_{e}^{t\beta}(y)=N_{e}^{t\beta}(x)$.
(6) If $y\in N_{e}^{n\beta}(x)$, then $N_{e}^{n\beta}(y)=N_{e}^{n\beta}(x)$.
Proof (1)-(6) are obvious.
Definition 4.37.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $a\in A\sqsubset E$.
(1) If $Cov^{p\beta}(A)=Cov^{p\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $p\beta$-partition, denoted as $a\in R^{p\beta}(A)$; if $Cov^{p\beta}(A)\neq Cov^{p\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $p\beta$-partition, denoted as $a\in Core^{p\beta}(A)$.
(2) If $Cov^{a\beta}(A)=Cov^{a\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $a\beta$-partition, denoted as $a\in R^{a\beta}(A)$; if $Cov^{a\beta}(A)\neq Cov^{a\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $a\beta$-partition, denoted as $a\in Core^{a\beta}(A)$.
(3) If $Cov^{m\beta}(A)=Cov^{m\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $m\beta$-partition, denoted as $a\in R^{m\beta}(A)$; if $Cov^{m\beta}(A)\neq Cov^{m\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $m\beta$-partition, denoted as $a\in Core^{m\beta}(A)$.
(4) If $Cov^{s\beta}(A)=Cov^{s\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $s\beta$-partition, denoted as $a\in R^{s\beta}(A)$; if $Cov^{s\beta}(A)\neq Cov^{s\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $s\beta$-partition, denoted as $a\in Core^{s\beta}(A)$.
(5) If $Cov^{t\beta}(A)=Cov^{t\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $t\beta$-partition, denoted as $a\in R^{t\beta}(A)$; if $Cov^{t\beta}(A)\neq Cov^{t\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $t\beta$-partition, denoted as $a\in Core^{t\beta}(A)$.
(6) If $Cov^{n\beta}(A)=Cov^{n\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $n\beta$-partition, denoted as $a\in R^{n\beta}(A)$; if $Cov^{n\beta}(A)\neq Cov^{n\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $n\beta$-partition, denoted as $a\in Core^{n\beta}(A)$.
Theorem 4.38.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$. The following statements hold,
(1) $\sqcup\{C:C\in Cov^{p\beta}(A)\}=U$.
(2) $\sqcup\{C:C\in Cov^{a\beta}(A)\}=U$.
(3) $\sqcup\{C:C\in Cov^{m\beta}(A)\}=U$.
(4) $\sqcup\{C:C\in Cov^{s\beta}(A)\}=U$.
(5) $\sqcup\{C:C\in Cov^{t\beta}(A)\}=U$.
(6) $\sqcup\{C:C\in Cov^{n\beta}(A)\}=U$.
Proof (1) For all $x\in U$, $x\in N_{e}^{p\beta}(x)$, then $x\in\sqcap\{N_{e}^{p\beta}(x):e\in A\}=N_{A}^{p\beta}(x)\in Cov^{p\beta}(A)$. Then, $U\sqsubset\sqcup\{C:C\in Cov^{p\beta}(A)\}$.
On the other hand, $C\sqsubset U$ for all $C\in Cov^{p\beta}(A)$, then $\sqcup\{C:C\in Cov^{p\beta}(A)\}\sqsubset U$.
Hence, $\sqcup\{C:C\in Cov^{p\beta}(A)\}=U$.
The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
Theorem 4.39.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$. The following statements hold,
(1) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{p\beta}(A)$.
(2) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{a\beta}(A)$.
(3) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{m\beta}(A)$.
(4) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{s\beta}(A)$.
(5) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{t\beta}(A)$.
(6) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{n\beta}(A)$.
Proof (1) If there are $C_{i},C_{j}\in Cov^{p\beta}(A)$ such that $C_{i}\sqcap C_{j}\neq\emptyset$, then there is $x\in C_{i}\sqcap C_{j}$.
Since $C_{i}\neq C_{j}$, then one of $C_{i}-C_{j}\neq\emptyset$ and $C_{j}-C_{i}\neq\emptyset$ holds. These two cases can derive the following two cases, case 1, $y\in C_{i}$ and $y\not\in C_{j}$; and case 2, $y\not\in C_{i}$ and $y\in C_{j}$. The proof of case 1 is shown as follows,
(i) Since $x,y\in C_{i}\in Cov^{p\beta}(A)$, then $\beta(e)\subset_{p}F(e)(x)$ if and only if $\beta(e)\subset_{p}F(e)(y)$ for $e\in\overline{A}\sqsubset A$; and $\beta(e^{\prime})\not\subset_{p}F(e^{\prime})(x)$ if and only if $\beta(e^{\prime})\not\subset_{p}F(e^{\prime})(y)$ for others $e^{\prime}\in A-\overline{A}$.
(ii) Since $x\in C_{j}$, $y\not\in C_{j}$ and $C_{j}\in Cov^{p\beta}(A)$, then there is at least one $e^{*}\in A$ such that $\beta(e^{*})\subset_{p}F(e^{*})(x)$ and $\beta(e^{*})\not\subset_{p}F(e^{*})(y)$, or $\beta(e^{*})\not\subset_{p}F(e^{*})(x)$ and $\beta(e^{*})\subset_{p}F(e^{*})(y)$.
There is a contradiction in (i) and (ii). Hence, $C_{i}\sqcap C_{j}=\emptyset$.
The proof of case 2 ($y\not\in C_{i}$ and $y\in C_{j}$) is similar to the proof of case 1 ($y\in C_{i}$ and $y\not\in C_{j}$), and case 2 also can obtain $C_{i}\sqcap C_{j}=\emptyset$. Then, $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{p\beta}(A)$.
The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
Theorem 4.40.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$, $x,y\in U$. The following statements hold,
(1) If $y\in N_{A}^{p\beta}(x)$, then $N_{A}^{p\beta}(x)=N_{A}^{p\beta}(y)$.
(2) If $y\in N_{A}^{a\beta}(x)$, then $N_{A}^{a\beta}(x)=N_{A}^{a\beta}(y)$.
(3) If $y\in N_{A}^{m\beta}(x)$, then $N_{A}^{m\beta}(x)=N_{A}^{m\beta}(y)$.
(4) If $y\in N_{A}^{s\beta}(x)$, then $N_{A}^{s\beta}(x)=N_{A}^{s\beta}(y)$.
(5) If $y\in N_{A}^{t\beta}(x)$, then $N_{A}^{t\beta}(x)=N_{A}^{t\beta}(y)$.
(6) If $y\in N_{A}^{n\beta}(x)$, then $N_{A}^{n\beta}(x)=N_{A}^{n\beta}(y)$.
Proof (1) Suppose $N_{A}^{p\beta}(x)\neq N_{A}^{p\beta}(y)$, since $N_{A}^{p\beta}(x),N_{A}^{p\beta}(y)\in Cov^{p\beta}(A)$, let $N_{A}^{p\beta}(x)=C_{i}$ and $N_{A}^{p\beta}(y)=C_{j}$. By Theorem 4.39, $C_{i}\sqcap C_{j}=\emptyset$.
On the other hand, $y\in N_{A}^{p\beta}(x)$ and $y\in N_{A}^{p\beta}(y)$, then $\emptyset\neq\{y\}\sqsubset C_{i}\sqcap C_{j}$. There is a contradiction.
Hence, $N_{A}^{p\beta}(x)=N_{A}^{p\beta}(y)$.
The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
Theorem 4.41.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $a\in A\sqsubset E$. The following statements hold,
(1) $a\in R^{p\beta}(A)$ if and only if $N_{a}^{p\beta}(x)=\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}$ for all $x\in U$.
(2) $a\in R^{a\beta}(A)$ if and only if $N_{a}^{a\beta}(x)=\sqcup\{N_{A-\{a\}}^{a\beta}(y):y\in N_{a}^{a\beta}(x)\}$ for all $x\in U$.
(3) $a\in R^{m\beta}(A)$ if and only if $N_{a}^{m\beta}(x)=\sqcup\{N_{A-\{a\}}^{m\beta}(y):y\in N_{a}^{m\beta}(x)\}$ for all $x\in U$.
(4) $a\in R^{s\beta}(A)$ if and only if $N_{a}^{s\beta}(x)=\sqcup\{N_{A-\{a\}}^{s\beta}(y):y\in N_{a}^{s\beta}(x)\}$ for all $x\in U$.
(5) $a\in R^{t\beta}(A)$ if and only if $N_{a}^{t\beta}(x)=\sqcup\{N_{A-\{a\}}^{t\beta}(y):y\in N_{a}^{t\beta}(x)\}$ for all $x\in U$.
(6) $a\in R^{n\beta}(A)$ if and only if $N_{a}^{n\beta}(x)=\sqcup\{N_{A-\{a\}}^{n\beta}(y):y\in N_{a}^{n\beta}(x)\}$ for all $x\in U$.
Proof (1) The hesitant fuzzy information system $(U,F,A)$ has one more parameter $a$ than $(U,F,A-\{a\})$. Then $N_{A}^{p\beta}(x)=\sqcap\{N_{e}^{p\beta}(x):e\in A\}=N_{A-\{a\}}^{p\beta}(x)\sqcap N_{a}^{p\beta}(x)=\sqcap\{N_{e}^{p\beta}(x):e\in A-\{a\}\}\sqcap N_{a}^{p\beta}(x)$.
The proof of necessity.
$a\in R^{p\beta}(A)$, then $N_{A}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)$ for all $x\in U$. $N_{A}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)\sqcap N_{a}^{p\beta}(x)$, then $N_{A-\{a\}}^{p\beta}(x)\sqsubset N_{a}^{p\beta}(x)$ for all $x\in U$.
By Proposition 4.36, if $y\in N_{a}^{p\beta}(x)$, then $N_{a}^{p\beta}(x)=N_{a}^{p\beta}(y)$.
If $y\in N_{a}^{p\beta}(x)$, then $N_{A-\{a\}}^{p\beta}(y)\sqsubset N_{a}^{p\beta}(y)=N_{a}^{p\beta}(x)$, then $\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}\sqsubset N_{a}^{p\beta}(x)$.
Since $y\in N_{A-\{a\}}^{p\beta}(y)$, then $N_{a}^{p\beta}(x)\sqsubset\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}$ holds.
Then $N_{a}^{p\beta}(x)=\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}$.
The proof of sufficiency.
If $N_{a}^{p\beta}(x)=\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}$, since $x\in N_{a}^{p\beta}(x)$, then $N_{A-\{a\}}^{p\beta}(x)\sqsubset N_{a}^{p\beta}(x)$. Then $N_{A}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)\sqcap N_{a}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)$, i.e., $a\in R^{p\beta}(A)$.
The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
Theorem 4.42.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $a,b\in A\sqsubset E$. The following statements hold,
(1) If $N_{a}^{p\beta}=N_{b}^{p\beta}$, then $a\in R^{p\beta}(A)$ ($b\in R^{p\beta}(A)$).
(2) If $N_{a}^{a\beta}=N_{b}^{a\beta}$, then $a\in R^{a\beta}(A)$ ($b\in R^{a\beta}(A)$).
(3) If $N_{a}^{m\beta}=N_{b}^{m\beta}$, then $a\in R^{m\beta}(A)$ ($b\in R^{m\beta}(A)$).
(4) If $N_{a}^{s\beta}=N_{b}^{s\beta}$, then $a\in R^{s\beta}(A)$ ($b\in R^{s\beta}(A)$).
(5) If $N_{a}^{t\beta}=N_{b}^{t\beta}$, then $a\in R^{t\beta}(A)$ ($b\in R^{t\beta}(A)$).
(6) If $N_{a}^{n\beta}=N_{b}^{n\beta}$, then $a\in R^{n\beta}(A)$ ($b\in R^{n\beta}(A)$).
Proof (1) $N_{a}^{p\beta}=N_{b}^{p\beta}$, i.e., $N_{a}^{p\beta}(x)=N_{b}^{p\beta}(x)$ for all $x\in U$, then $N_{A}^{p\beta}(x)=N_{A-\{a,b\}}^{p\beta}(x)\sqcap N_{b}^{p\beta}(x)\sqcap N_{a}^{p\beta}(x)=N_{A-\{a,b\}}^{p\beta}(x)\sqcap N_{b}^{p\beta}(x)\sqcap N_{b}^{p\beta}(x)=N_{A-\{a,b\}}^{p\beta}(x)\sqcap N_{b}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)$, i.e., $a\in R^{p\beta}(A)$.
$b\in R^{p\beta}(A)$ can be proved by the same manner of the case $a\in R^{p\beta}(A)$.
The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
$a\in R^{p\beta}(A)$ or $b\in R^{p\beta}(A)$ holds, however, $\{a,b\}\sqsubset R^{p\beta}(A)$ is not determinate. The situations of (2)-(6) are the same as the situations of (1).
Theorem 4.43.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $a,b\in A\sqsubset E$. The following statements hold,
(1) If $b\in R^{p\beta}(A-\{a\})$, then $b\in R^{p\beta}(A)$.
(2) If $b\in R^{a\beta}(A-\{a\})$, then $b\in R^{a\beta}(A)$.
(3) If $b\in R^{m\beta}(A-\{a\})$, then $b\in R^{m\beta}(A)$.
(4) If $b\in R^{s\beta}(A-\{a\})$, then $b\in R^{s\beta}(A)$.
(5) If $b\in R^{t\beta}(A-\{a\})$, then $b\in R^{t\beta}(A)$.
(6) If $b\in R^{n\beta}(A-\{a\})$, then $b\in R^{n\beta}(A)$.
Proof (1) Since $b\in R^{p\beta}(A-\{a\})$, by Theorem 4.41, $N_{b}^{p\beta}(x)=\sqcup\{N_{A-\{a,b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}$. Since $N_{A-\{b\}}^{p\beta}(y)=N_{A-\{a,b\}}^{p\beta}(y)\sqcap N_{a}^{p\beta}(y)\sqsubset N_{A-\{a,b\}}^{p\beta}(y)$, then $N_{b}^{p\beta}(x)=\sqcup\{N_{A-\{a,b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}\sqsupset\sqcup\{N_{A-\{b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}$.
Since $y\in N_{A-\{b\}}^{p\beta}(y)$, then $N_{b}^{p\beta}(x)\sqsubset\sqcup\{N_{A-\{b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}$ holds.
Based on the results above, then $N_{b}^{p\beta}(x)=\sqcup\{N_{A-\{b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}$, by Theorem 4.41, $b\in R^{p\beta}(A)$.
The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
Table 1. A triple $(U,F,E)$.
$U$ $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$
$x_{1}$ 0.5, 0.3, 0.2 0.7, 0.6, 0.5 0.8, 0.7 0.7, 0.2 0.6, 0.5, 0.3, 0.3 0.6, 0.3, 0.3 0.8, 0.8, 0.6, 0.5
$x_{2}$ 0.4, 0.4, 0.3 0.7, 0.7 0.7, 0.6, 0.5 0.7, 0.5, 0.5 0.7, 0.5, 0.3, 0.3 0.6, 0.3, 0.3 0.6, 0.4, 0.3
$x_{3}$ 0.3, 0.2, 0.1 0.7, 0.6 0.6, 0.5 0.7, 0.2 0.6, 0.5, 0.3, 0.1 0.6, 0.3 0.3, 0.3
$x_{4}$ 0.8, 0.7, 0.7 0.4, 0.3 0.3, 0.3, 0.2 0.9, 0.7 0.5, 0.5, 0.2 0.9, 0.8, 0.7, 0.6 0.9, 0.7
$x_{5}$ 0.8, 0.7, 0.6 0.3, 0.3, 0.2 0.3, 0.3, 0.1 0.8, 0.8, 0.7 0.6, 0.2, 0.2 0.7, 0.6, 0.5, 0.3 0.5, 0.3
$x_{6}$ 0.7, 0.7, 0.6 0.8, 0.6, 0.5 0.2, 0.1, 0.1 0.8, 0.7, 0.7 0.6, 0.2 0.6, 0.5, 0.3, 0.1 0.5, 0.3, 0.1, 0.1
Example 4.44.
$U=\{x_{1},x_{2},\cdots,x_{6}\}$, $E=\{e_{1},e_{2},\cdots,e_{7}\}$, a triple $(U,F,E)$ is shown in Table 1. Let $\beta=\frac{0.6,0.5,0.3}{e_{1}}+\frac{0.5,0.5,0.3}{e_{2}}+\frac{0.4,0.3,0.3}{e_{3}}+\frac{0.7,0.6,0.5}{e_{4}}+\frac{0.6,0.3,0.2}{e_{5}}+\frac{0.6,0.3,0.1}{e_{6}}+\frac{0.6,0.3,0.2}{e_{7}}$.
(1) For the Theorem 4.43 (1), $a\in R^{p\beta}(A)$ and $b\in R^{p\beta}(A)$, however $b\not\in R^{p\beta}(A-\{a\})$.
$N_{e_{1}}^{p\beta}=N_{e_{3}}^{p\beta}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$, $N_{e_{2}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{6}\},\{x_{4},x_{5}\}\}$.
Let $a=e_{1}$, $b=e_{3}$ and $A=\{e_{1},e_{2},e_{3}\}$. $Cov^{p\beta}(A)=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5}\},\{x_{6}\}\}=Cov^{p\beta}(A-\{a\})=Cov^{p\beta}(A-\{b\})$, i.e., $a\in R^{p\beta}(A)$ and $b\in R^{p\beta}(A)$.
$Cov^{p\beta}(A-\{a,b\})=\{\{x_{1},x_{2},x_{3},x_{6}\},\{x_{4},x_{5}\}\}$. $Cov^{p\beta}(A-\{a,b\})\neq Cov^{p\beta}(A-\{a\})$, i.e., $b\not\in R^{p\beta}(A-\{a\})$.
(2) The cases of Theorem 4.43 (2), (3), (4) and (6) are similar to the case of Theorem 4.43 (1).
$a\in R^{p\beta}(A)$ and $b\in R^{p\beta}(A)$, $b\not\in R^{p\beta}(A-\{a\})$.
With the similar ways above, $Cov^{a\beta}(A)=Cov^{a\beta}(A-\{a\})=Cov^{a\beta}(A-\{b\})$, $Cov^{a\beta}(A-\{a,b\})\neq Cov^{a\beta}(A-\{a\})$, i.e., $a\in R^{a\beta}(A)$ and $b\in R^{a\beta}(A)$, $b\not\in R^{a\beta}(A-\{a\})$.
Furthermore, $a\in R^{m\beta}(A)$ and $b\in R^{m\beta}(A)$, $b\not\in R^{m\beta}(A-\{a\})$; $a\in R^{s\beta}(A)$ and $b\in R^{s\beta}(A)$, $b\not\in R^{s\beta}(A-\{a\})$; and $a\in R^{n\beta}(A)$ and $b\in R^{n\beta}(A)$, $b\not\in R^{n\beta}(A-\{a\})$.
(3) For the Theorem 4.43 (5), $a\in R^{t\beta}(A)$ and $b\in R^{t\beta}(A)$, however $b\not\in R^{t\beta}(A-\{a\})$.
$N_{e_{5}}^{t\beta}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}=N_{e_{6}}^{t\beta}$, $N_{e_{7}}^{t\beta}=\{\{x_{1}\},\{x_{2},x_{3},x_{4},x_{5},x_{6}\}\}$.
Let $a=e_{5}$, $b=e_{6}$, $A=\{e_{5},e_{6},e_{7}\}$. $Cov^{t\beta}(A)=Cov^{t\beta}(A-\{a\})=Cov^{t\beta}(A-\{b\})=\{\{x_{1}\},\{x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$. $Cov^{t\beta}(A-\{a,b\})=\{\{x_{1}\},\{x_{2},x_{3},x_{4},x_{5},x_{6}\}\neq Cov^{t\beta}(A-\{a\})$.
$a\in R^{t\beta}(A)$ and $b\in R^{t\beta}(A)$, $b\not\in R^{t\beta}(A-\{a\})$.
Let $\mathscr{B}^{p}=\{B^{p}:Cov^{p\beta}(A)=Cov^{p\beta}(A-B^{p})\}$ and $\mathcal{I}^{p}=\{I^{p}:I^{p}=A-B^{p},B^{p}\in\mathscr{B}^{p}\}$ be two families over $A$ in $S_{p}=(U,F,A)_{\beta}$. The similar ways to define ($\mathscr{B}^{a},\mathcal{I}^{a}$), ($\mathscr{B}^{m},\mathcal{I}^{m}$), ($\mathscr{B}^{s},\mathcal{I}^{s}$), ($\mathscr{B}^{t},\mathcal{I}^{t}$) and ($\mathscr{B}^{n},\mathcal{I}^{n}$) in $S_{a}=(U,F,A)_{\beta}$, $S_{m}=(U,F,A)_{\beta}$, $S_{s}=(U,F,A)_{\beta}$, $S_{t}=(U,F,A)_{\beta}$ and $S_{n}=(U,F,A)_{\beta}$, respectively.
Theorem 4.45.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$. The following statements hold,
(1) $Core^{p\beta}(A)=\sqcap_{I^{p}\in\mathcal{I}^{p}}I^{p}$.
(2) $Core^{a\beta}(A)=\sqcap_{I^{a}\in\mathcal{I}^{a}}I^{a}$.
(3) $Core^{m\beta}(A)=\sqcap_{I^{m}\in\mathcal{I}^{m}}I^{m}$.
(4) $Core^{s\beta}(A)=\sqcap_{I^{s}\in\mathcal{I}^{s}}I^{s}$.
(5) $Core^{t\beta}(A)=\sqcap_{I^{t}\in\mathcal{I}^{t}}I^{t}$.
(6) $Core^{n\beta}(A)=\sqcap_{I^{n}\in\mathcal{I}^{n}}I^{n}$.
Proof (1) For all $B^{p}\in\mathscr{B}^{p}$, $Core^{p\beta}(A)\sqsubset A-B^{p}$, then $Core^{p\beta}(A)\sqsubset\sqcap_{B^{p}\in\mathscr{B}^{p}}(A-B^{p})=\sqcap_{I^{p}\in\mathcal{I}^{p}}I^{p}$.
For all $e\not\in Core^{p\beta}(A)$, $\{e\}\in\mathscr{B}^{p}$, then $(Core^{p\beta}(A))^{c}\sqsubset\sqcup_{B^{p}\in\mathscr{B}^{p}}B^{p}$. Then $Core^{p\beta}(A)\sqsupset(\sqcup_{B^{p}\in\mathscr{B}^{p}}B^{p})^{c}=\sqcap_{B^{p}\in\mathscr{B}^{p}}(B^{p})^{c}=\sqcap_{B^{p}\in\mathscr{B}^{p}}(A-B^{p})=\sqcap_{I^{p}\in\mathcal{I}^{p}}I^{p}$.
Hence, $Core^{p\beta}(A)=\sqcap_{I^{p}\in\mathcal{I}^{p}}I^{p}$.
The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
Theorem 4.46.
Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$. The following statements hold,
(1) $e\in Core^{p\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{p_{1}}=(U,F,A)_{\beta}$ and $S_{p_{2}}=(U,F,A-\{e\})_{\beta}$.
(2) $e\in Core^{a\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{a_{1}}=(U,F,A)_{\beta}$ and $S_{a_{2}}=(U,F,A-\{e\})_{\beta}$.
(3) $e\in Core^{m\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{m_{1}}=(U,F,A)_{\beta}$ and $S_{m_{2}}=(U,F,A-\{e\})_{\beta}$.
(4) $e\in Core^{s\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{s_{1}}=(U,F,A)_{\beta}$ and $S_{s_{2}}=(U,F,A-\{e\})_{\beta}$.
(5) $e\in Core^{t\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{t_{1}}=(U,F,A)_{\beta}$ and $S_{t_{2}}=(U,F,A-\{e\})_{\beta}$.
(6) $e\in Core^{n\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{n_{1}}=(U,F,A)_{\beta}$ and $S_{n_{2}}=(U,F,A-\{e\})_{\beta}$.
Proof (1) To denote $A^{\prime}=A-\{e\}$.
The proof of necessity.
$e\in Core^{p\beta}(A)$, then $Cov^{p\beta}(A)\neq Cov^{p\beta}(A-\{e\})=Cov^{p\beta}(A^{\prime})$. There is $x_{0}\in U$ satisfying $N_{A}^{p\beta}(x_{0})\neq N_{A^{\prime}}^{p\beta}(x_{0})$. Then there is $y_{0}$ satisfying $y_{0}\in N_{A}^{p\beta}(x_{0})$ and $y_{0}\not\in N_{A^{\prime}}^{p\beta}(x_{0})$, or another case $y_{0}\not\in N_{A}^{p\beta}(x_{0})$ and $y_{0}\in N_{A^{\prime}}^{p\beta}(x_{0})$. There is a pair of $x_{0},y_{0}\in U$ that their relations are different in $S_{p_{1}}=(U,F,A)_{\beta}$ and $S_{p_{2}}=(U,F,A-\{e\})_{\beta}$.
The proof of sufficiency.
For any $x,y\in U$, if the relations of $x$ and $y$ are different in $S_{p_{1}}=(U,F,A)_{\beta}$ and $S_{p_{2}}=(U,F,A-\{e\})_{\beta}$, i.e., $y\in N_{A}^{p\beta}(x)$ and $y\not\in N_{A^{\prime}}^{p\beta}(x)$, or another case $y\not\in N_{A}^{p\beta}(x)$ and $y\in N_{A^{\prime}}^{p\beta}(x)$. In either case, we can deduce $N_{A}^{p\beta}(x)\neq N_{A^{\prime}}^{p\beta}(x)$, i.e., $Cov^{p\beta}(A)\neq Cov^{p\beta}(A-\{e\})$, then $e\in Core^{p\beta}(A)$.
The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
4.2 Example and algorithm of parameter reduction in hesitant fuzzy information systems
In this subsection, we provide an example and an algorithm of parameter reduction based on the $p\beta$-partition in hesitant fuzzy information systems. The parameter reductions based on other kinds of $\beta$-partitions can be calculated with the similar method of $p\beta$-partition.
Example 4.47.
$(U,F,E)$ and $\beta$ are shown in the Example 4.44. Let $A=E$, we can thus calculate the families $\mathscr{B}^{p}$ and $\mathcal{I}^{p}$ of hesitant fuzzy information system $(U,F,A)$ based on $p\beta$-partition.
We show the processes of calculating the $\mathscr{B}^{p}$ and $\mathcal{I}^{p}$ in the following steps:
Step 1. Calculate $N_{e_{i}}^{p\beta}$ for every $e_{i}\in A$.
$N_{e_{1}}^{p\beta}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$,
$N_{e_{2}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{6}\},\{x_{4},x_{5}\}\}$,
$N_{e_{3}}^{p\beta}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$,
$N_{e_{4}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\},\emptyset\}$,
$N_{e_{5}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{5},x_{6}\},\{x_{4}\}\}$,
$N_{e_{6}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\},\emptyset\}$,
$N_{e_{7}}^{p\beta}=\{\{x_{1},x_{2},x_{4}\},\{x_{3},x_{5},x_{6}\}\}$.
Step 2. Obtain an initial $\mathscr{B}^{p}$, if $e_{i}\in R^{p\beta}(A)$, then $\{e_{i}\}\in\mathscr{B}^{p}$. $|B^{p}|=1$ for every $B^{p}$ in the initial $\mathscr{B}^{p}$.
In this example, the initial $\mathscr{B}^{p}=\{\{e_{1}\},\{e_{3}\},\{e_{4}\},\newline \{e_{5}\},\{e_{6}\}\}$.
Step 3. Select each $B^{p}_{i}\in\mathscr{B}^{p}$ one by one, if $e_{i}\in R^{p\beta}(A-B^{p}_{i})$ for an $e_{i}\in A-B^{p}_{i}$, then remove $B^{p}_{i}$ from $\mathscr{B}^{p}$ and add $B^{p}_{i}\sqcup\{e_{i}\}$ to $\mathscr{B}^{p}$. Here, $e_{i}$ may not be the only one that satisfies $e\in R^{p\beta}(A-B^{p}_{i})$ in some cases. Thus, if $e_{j}\in R^{p\beta}(A-B^{p}_{i})$ for $e_{j}\in A-B^{p}_{i}$, then $B^{p}_{i}\sqcup\{e_{j}\}$ is also added to $\mathscr{B}^{p}$.
For $B^{p}_{1}=\{e_{1}\}$, $e_{4}\in R^{p\beta}(A-B^{p}_{1})$, $e_{5}\in R^{p\beta}(A-B^{p}_{1})$ and $e_{6}\in R^{p\beta}(A-B^{p}_{1})$, then $\{e_{1}\}$ is removed from $\mathscr{B}^{p}$, and $B^{p}_{1}\sqcup\{e_{4}\}=\{e_{1},e_{4}\}$, $B^{p}_{1}\sqcup\{e_{5}\}=\{e_{1},e_{5}\}$ and $B^{p}_{1}\sqcup\{e_{6}\}=\{e_{1},e_{6}\}$ are added to $\mathscr{B}^{p}$. Based on the results above, then $\mathscr{B}^{p}=\{\{e_{1},e_{4}\},\{e_{1},e_{5}\},\{e_{1},e_{6}\},\{e_{3}\},\{e_{4}\},\{e_{6}\}\}$.
To deal with other $B^{p}\in\mathscr{B}^{p}$, then we obtain $\mathscr{B}^{p}=\{\{e_{1},e_{4}\},\{e_{1},e_{5}\},\{e_{1},e_{6}\},\{e_{3},e_{4}\},\{e_{3},e_{5}\},\{e_{3},e_{6}\},\newline \{e_{4},e_{5}\},\{e_{4},e_{6}\},\{e_{5},e_{6}\}\}$.
Step 4. Repeat the similar operations of Step 3 until $e\not\in R^{p\beta}(A-B^{p})$ for every $B^{p}\in\mathscr{B}^{p}$ and every $e\in A-B^{p}$. Then, the final $\mathscr{B}^{p}$ is obtained as follows:
$\mathscr{B}^{p}=\{\{e_{1},e_{4},e_{5},e_{6}\},\{e_{3},e_{4},e_{5},e_{6}\}\}$.
Step 5. Calculate $\mathcal{I}^{p}$, $\mathcal{I}^{p}=\{A-B^{p}:B^{p}\in\mathscr{B}^{p}\}$.
$\mathcal{I}^{p}=\{\{e_{2},e_{3},e_{7}\},\{e_{1},e_{2},e_{7}\}\}$.
$Cov^{p\beta}(A)=Cov^{p\beta}(I^{p})$ for all $I^{p}\in\mathcal{I}^{p}$.
A detailed algorithm of parameter reduction of hesitant fuzzy information systems based on the $p\beta$-partition is shown as follows,
Algorithm 1: Parameter reduction of hesitant fuzzy information systems based on $p\beta$-partition.
Input: Hesitant fuzzy information system $S_{p}=(U,F,A)_{\beta}$.
Output: $\mathscr{B}^{p}$ and $\mathcal{I}^{p}$.
1. Calculate $N_{e_{i}}^{p\beta}$ for every $e_{i}\in A$.
2. Obtain an initial $\mathscr{B}^{p}$, if $e_{i}\in R^{p\beta}(A)$, then $\{e_{i}\}\in\mathscr{B}^{p}$.
3. While There is $B^{p}\in\mathscr{B}^{p}$ and $e\in A-B^{p}$ such that $e\in R^{p\beta}(A-B^{p})$ do
4. repeat
5. $\mathscr{B}^{p}\leftarrow\mathscr{B}^{p}-\{B^{p}\}$.
6. While There is $e\in A-B^{p}$ that satisfied $e\in R^{p\beta}(A-B^{p})$ do
7. repeat
8. $\mathscr{B}^{p}\leftarrow\mathscr{B}^{p}\sqcup\{B^{p}\sqcup\{e\}\}$.
9. until all the $B^{p}\sqcup\{e\}$ are added to $\mathscr{B}^{p}$ for some $e\in A-B^{p}$ that satisfy $e\in R^{p\beta}(A-B^{p})$.
10. end while
11. until $e\not\in R^{p\beta}(A-B^{p})$ for every $B^{p}\in\mathscr{B}^{p}$ and every $e\in A-B^{p}$.
12. end while
13. $\mathcal{I}^{p}\leftarrow\{A-B^{p}:B^{p}\in\mathscr{B}^{p}\}$.
5 Application example of foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems
Some complex systems, such as large wind turbines and shield tunneling machines, have a few failed samples and lots of factors that may affect system failure. These kinds of systems make it difficult to make system diagnosis or decision of system replacement. We propose a multi-strength intelligent classifier to give an alternative method to deal with the cases stated above.
We assumed that $|F(e)(x_{i})|=|F(e)(x_{j})|$ for all $e\in E$ and $x_{i},x_{j}\in U$, then the relationship “$\subset_{t}$” will not be considered in Section 5.
5.1 Knowledge bases and similar degree functions
Definition 5.48.
PALA ; PASA A pair $(U,\mathcal{R})$ is called a knowledge base, where $\mathcal{R}\neq\emptyset$ and $\mathcal{R}$ is a family of equivalence relations on $U$.
As shown in Example 4.44, let $A=\{e_{1},e_{2},e_{3},e_{4}\}$, $Cov^{p\beta}(A)=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5}\},\{x_{6}\}\}$ can be considered as a knowledge, denoted as $(U,R_{(A,p\beta)})$. Furthermore, $R_{(A,p\beta)}(x_{1},x_{2})=1$ or $x_{2}\in[x_{1}]_{R_{(A,p\beta)}}$; $R_{(A,p\beta)}(x_{1},x_{4})=0$ or $x_{4}\not\in[x_{1}]_{R_{(A,p\beta)}}$. $Cov^{a\beta}(A)$, $Cov^{m\beta}(A)$, $Cov^{s\beta}(A)$ and $Cov^{n\beta}(A)$ are knowledge generated by $R_{(A,a\beta)}$, $R_{(A,m\beta)}$, $R_{(A,s\beta)}$ and $R_{(A,n\beta)}$, respectively.
Let *=$\mathcal{R}^{*}=\{R_{(A,p\beta)},R_{(A,a\beta)},R_{(A,m\beta)},R_{(A,s\beta)},\newline R_{(A,n\beta)}\}$, then $(U,\mathcal{R}^{*})$ is a knowledge base.
Definition 5.49.
Let $(U,F,A)_{\beta}$ be a hesitant fuzzy information system, $x_{i},x_{j}\in U$, $e\in A\sqsubset E$, and $\neg A=\{\neg e_{i}:e_{i}\in A\}$,
(1) Attribute recognition functions:
$\xi^{p\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{p}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
$\xi^{a\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{a}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
$\xi^{m\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{m}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
$\xi^{s\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{s}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
$\xi^{n\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{n}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
(2) Similar degree functions:
$Sim^{p\beta}(x_{i},x_{j})=\frac{|\xi^{p\beta}(x_{i})\sqcap\xi^{p\beta}(x_{j})|}{|A|}$.
$Sim^{a\beta}(x_{i},x_{j})=\frac{|\xi^{a\beta}(x_{i})\sqcap\xi^{a\beta}(x_{j})|}{|A|}$.
$Sim^{m\beta}(x_{i},x_{j})=\frac{|\xi^{m\beta}(x_{i})\sqcap\xi^{m\beta}(x_{j})|}{|A|}$.
$Sim^{s\beta}(x_{i},x_{j})=\frac{|\xi^{s\beta}(x_{i})\sqcap\xi^{s\beta}(x_{j})|}{|A|}$.
$Sim^{n\beta}(x_{i},x_{j})=\frac{|\xi^{n\beta}(x_{i})\sqcap\xi^{n\beta}(x_{j})|}{|A|}$.
As shown in Example 4.44, $\beta(e_{1})\not\subset_{p}F(e_{1})(x_{1})$; and $\beta(e_{i})\subset_{p}F(e_{i})(x_{1})$ for $e_{i}\in E-\{e_{1}\}$. It means that the object $x_{1}$ has no characteristic concerning $e_{1}$ based on $p\beta$-partition. Then, $\xi^{p\beta}(x_{1})=\{\neg e_{1},e_{2},e_{3},e_{4},e_{5},e_{6},e_{7}\}$. In the same manner, $\xi^{p\beta}(x_{5})=\{e_{1},\neg e_{2},\neg e_{3},e_{4},e_{5},e_{6},\neg e_{7}\}$. Furthermore, $Sim^{p\beta}(x_{1},x_{5})=\frac{|\xi^{p\beta}(x_{1})\sqcap\xi^{p\beta}(x_{5})|}{|E|}=\frac{|\{e_{4},e_{5},e_{6}\}|}{|E|}=\frac{3}{7}$.
Theorem 5.50.
For $X\sqsubset U$, the following statements hold in the knowledge base $(U,\mathcal{R}^{*})$.
(1) $Sim^{p\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,p\beta)})$.
(2) $Sim^{a\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,a\beta)})$.
(3) $Sim^{m\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,m\beta)})$.
(4) $Sim^{s\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,s\beta)})$.
(5) $Sim^{n\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,n\beta)})$.
Proof (1) For $X\in(U,R_{(A,p\beta)})=Cov^{p\beta}(A)$ and $x\in X$, since $\{x,y\}\sqsubset X=\sqcap\{N_{e}^{p\beta}(x):e\in A\}$, then $y\in N_{e}^{p\beta}(x)$ for all $e\in A$. Therefore, we can obtain that $\xi^{p\beta}(x)=\xi^{p\beta}(y)$, then $Sim^{p\beta}(x,y)=1$.
The proofs of (2)-(5) are similar to the proof of (1)$\blacksquare$
5.2 Multi-strength intelligent classifier for health status diagnosis of complex systems
This subsection presents a multi-strength intelligent classifier in knowledge bases to diagnose the health status of complex systems. In the following, the health state partitions of the complex system, the termination training conditions of the intelligent classifier, and the classification way of the intelligent classifier will be introduced.
5.2.1 Partitions of health states
In order to use the results of the previous section to diagnose the health status of the system, we divide the health level of the system into five levels according to Definition 2.7, shown in Table 2. For example, $x\in U_{n}$ means that it is necessary credible that $x$ is a healthy equipment; $y\in U_{f}$ means that it is not credible that $y$ is a healthy equipment, i.e., $y$ is a failed equipment.
Table 2. Credible degrees that the equipment is healthy
Kinds of Credible degrees of Health Sample
partition systems’ health states grade set
$n\beta$-partition necessary credible G1 $U_{n}$
$s\beta$-partition strong credible G2 $U_{s}$
$m\beta$-partition middle credible G3 $U_{m}$
$a\beta$-partition middle credible G3 $U_{a}$
$p\beta$-partition weak credible G4 $U_{p}$
other cases not credible G5 $U_{f}$
5.2.2 Conditions under which the multi-strength intelligent classifier terminates training
If $Sim^{n\beta}(x_{i},x_{j})\geqslant PT$ ($Sim^{n\beta}(x_{i},x_{j})<NT$), then $x_{i}$ and $x_{j}$ are (not) considered as two samples in the same category based on $n\beta$-partition, where PT is a positive threshold and NT is a negative threshold.
$\frac{1}{C_{|U^{*}_{n}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{C_{|U^{*}_{s}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{s}}Sim^{s\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{|U^{*}_{s}|\times|U^{*}_{n}|}\sum\limits_{x_{i}\in U^{*}_{s},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})<NT,\\ \frac{1}{C_{|U^{*}_{m}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{m}}Sim^{m\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{|U^{*}_{m}|\times|U^{*}_{s}|}\sum\limits_{x_{i}\in U^{*}_{m},x_{j}\in U^{*}_{s}}Sim^{s\beta}(x_{i},x_{j})<NT,\\ \frac{1}{C_{|U^{*}_{a}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{a}}Sim^{a\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{|U^{*}_{a}|\times|U^{*}_{s}|}\sum\limits_{x_{i}\in U^{*}_{a},x_{j}\in U^{*}_{s}}Sim^{s\beta}(x_{i},x_{j})<NT,\\ \frac{1}{C_{|U^{*}_{p}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{p}}Sim^{p\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{|U_{1}|\times|U^{*}_{f}|}\sum\limits_{x_{i}\in U_{1},x_{j}\in U^{*}_{f}}Sim^{p\beta}(x_{i},x_{j})<NT,\\ \frac{1}{|U_{2}|\times|U^{*}_{f}|}\sum\limits_{x_{i}\in U_{2},x_{j}\in U^{*}_{f}}Sim^{m\beta}(x_{i},x_{j})<NT.\\ \end{aligned}\right.$ (1)
Based on the Proposition 2.9, $\subset_{n}\Rightarrow\subset_{s}\Rightarrow\subset_{a}\Rightarrow\subset_{p}$ and $\subset_{n}\Rightarrow\subset_{s}\Rightarrow\subset_{m}$, then we set $U_{1}=U^{*}_{n}\sqcup U^{*}_{s}\sqcup U^{*}_{a}\sqcup U^{*}_{p}$ and $U_{2}=U^{*}_{n}\sqcup U^{*}_{s}\sqcup U^{*}_{m}$ to distinguish $U^{*}_{f}$, where $U^{*}_{n}=U_{n}\sqcap U_{tain}$, $U^{*}_{s}=U_{s}\sqcap U_{tain}$, $U^{*}_{m}=U_{m}\sqcap U_{tain}$, $U^{*}_{a}=U_{a}\sqcap U_{tain}$, $U^{*}_{p}=U_{p}\sqcap U_{tain}$ and $U^{*}_{f}=U_{f}\sqcap U_{tain}$. The conditions for the intelligent classifier to terminate the training are shown in the inequality array (1) above, where $C_{N}^{2}=N\times(N-1)/2$.
The last two inequalities of the inequality array (1) are used to distinguish the failure samples of the healthy samples, the samples in $U_{1}$ and $U^{*}_{f}$ share fewer of the same attributes based on $p\beta$-partition, and the samples in $U_{2}$ and $U^{*}_{f}$ share fewer of the same attributes based on $m\beta$-partition. That is the last two inequalities of the inequality array (1) restrict samples from having many features of both healthy and failed samples and prevent the same sample from being classified as both healthy and failed by the classifier.
The inequality $\frac{1}{|U^{*}_{s}|\times|U^{*}_{n}|}\sum\limits_{x_{i}\in U^{*}_{s},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})<NT$ restricts samples $x_{i}\in U_{s}$ and $x_{j}\in U_{n}$ from having many features of both G1 and G2 health grades and prevents a sample $x_{i}\in U_{s}$ from being classified as $x_{i}\in U_{n}$, i.e., it prevents the classifier from classifying samples of low-level health states into a set of high-level health states. The other inequalities with a condition of “$<NT$” in array (1) serve a similar purpose.
5.2.3 The classification way of the multi-strength intelligent classifier
When the training of a multi-strength intelligent classifier is finished, it is used to classify other samples. Let $Cl(x)=\{Cl_{n}(x),Cl_{s}(x),Cl_{m}(x),Cl_{a}(x),Cl_{p}(x),\newline Cl_{f}(x)\}$, where
{$Cl_{n}(x)=\frac{1}{|U^{*}_{n}|}\sum\limits_{x_{i}\in U^{*}_{n}}Sim^{n\beta}(x,x_{i}),\\ Cl_{s}(x)=\frac{1}{|U^{*}_{s}|}\sum\limits_{x_{i}\in U^{*}_{s}}Sim^{s\beta}(x,x_{i}),\\ Cl_{m}(x)=\frac{1}{|U^{*}_{m}|}\sum\limits_{x_{i}\in U^{*}_{m}}Sim^{m\beta}(x,x_{i}),\\ Cl_{a}(x)=\frac{1}{|U^{*}_{a}|}\sum\limits_{x_{i}\in U^{*}_{a}}Sim^{a\beta}(x,x_{i}),\\ Cl_{p}(x)=\frac{1}{|U^{*}_{p}|}\sum\limits_{x_{i}\in U^{*}_{p}}Sim^{p\beta}(x,x_{i}),\\ Cl_{f}(x)=\frac{1}{|U^{*}_{f}|}\sum\limits_{x_{i}\in U^{*}_{f}}Sim^{p\beta}(x,x_{i}).\\ \end{aligned}\right.$ (2)
The values in $Cl(x)$ have three cases (i), (ii) and (iii) as follows.
(i) One of the values in $Cl(x)$ is greater than or equal to $PT$, then $x$ is assigned to the class corresponding to the maximum value in $Cl(x)$.
(ii) More than one value in $Cl(x)$ is greater than or equal to $PT$, then $x$ is assigned to more advanced health states. For example, $Cl_{n}(x)\geqslant PT$ and $Cl_{s}(x)\geqslant PT$, then $x$ is assigned to $U_{n}$. By Proposition 2.9, $\subset_{n}\Rightarrow\subset_{s}$, then case (ii) exists. On the other hand, the inequality $\frac{1}{|U^{*}_{s}|\times|U^{*}_{n}|}\sum\limits_{x_{i}\in U^{*}_{s},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})<NT$ prevents the sample $x\in U_{s}$ from being classified as $x\in U_{n}$. Hence, $x\in U_{s}$ cannot obtain $Cl_{n}(x)\geqslant PT$, but $x\in U_{n}$ can obtain $Cl_{s}(x)\geqslant PT$. If $Cl_{n}(x)\geqslant PT$ and $Cl_{s}(x)\geqslant PT$, then $x\in U_{n}$.
(iii) No value in $Cl(x)$ is greater than or equal to $PT$, the classifier is invalidated for $x$. For small sample size and high dimension data, it is difficult for the classifier to learn all the features of the data, case (iii) may be present in a new monitoring sample.
5.2.4 The algorithm of the multi-strength intelligent classifier
The algorithm for training the multi-strength intelligent classifier is shown as Algorithm 2, where $AcT$ is the threshold of the correct classification rate of the testing samples. And, referring back the assumption of Section 5, $|F(e)(x)|=|F(e)(y)|$ for all $x,y\in U$.
Algorithm 2: Multi-strength intelligent classifier.
Input: $PT$, $NT$, $AcT$, sample sets $U_{n}$, $U_{s}$, $U_{a}$, $U_{m}$, $U_{p}$ and $U_{f}$, and hesitant fuzzy information system $(U,F,E)$, where $U=U_{n}\sqcup U_{s}\sqcup U_{a}\sqcup U_{m}\sqcup U_{p}\sqcup U_{f}$.
Output: $U^{*}_{tain}$, $A^{*}$ and $\beta^{*}$.
1. $\beta(e)\leftarrow\{0,0,\cdots,0\}$, where $|\beta(e)|=|F(e)(x)|$ for each $e\in E$.
2. $AC\leftarrow 0$.
3. While $AC<AcT$ do
4. repeat
5. Two-thirds of the samples in $U$ are randomly selected as training samples $U_{tain}$ and the other third as testing samples $U_{test}$.
6. $Sim^{n\beta}(x_{i},x_{j})\leftarrow 0$ for all $x_{i},x_{j}\in U_{n}$.
7. While One of the inequalities in the inequality array (1) does not hold do
8. repeat
9. The genetic algorithm is used to iterate an indicator $\beta$.
10. Apply Algorithm 1 to calculate $\mathcal{I}^{p}$, $\mathcal{I}^{a}$, $\mathcal{I}^{m}$, $\mathcal{I}^{s}$ and $\mathcal{I}^{n}$.
11. $A\leftarrow I_{p}\sqcup I_{a}\sqcup I_{m}\sqcup I_{s}\sqcup I_{n}$, where $I_{p},I_{a},I_{m},I_{s}$ and $I_{n}$ are the random elements in $\mathcal{I}^{p}$, $\mathcal{I}^{a}$, $\mathcal{I}^{m}$, $\mathcal{I}^{s}$ and $\mathcal{I}^{n}$, respectively.
12. Calculate the factors in the array (1) ($\frac{1}{C_{|U^{*}_{n}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})$, etc), and judge whether all the inequalities in the inequality array (1) hold.
13. until All the inequalities in the inequality array (1) hold.
14. end while
15. Calculate $Cl(x)$ for $x\in U_{test}$ and classify $x$.
16. Calculate $AC$, where $AC=\frac{N_{0}}{|U_{tset}|}$ and $N_{0}$ is the number of testing samples that were correctly classified.
17. until $AC\geqslant AcT$.
18. end while
19. $U^{*}_{tain}\leftarrow U_{tain}$, $A^{*}\leftarrow A$ and $\beta^{*}\leftarrow\beta$.
When $U^{*}_{tain}$, $A^{*}$ and $\beta^{*}$ are obtained, a multi-strength intelligent classifier $(U^{*}_{tain},\mathcal{R}^{*},\{Sim\})$ is obtained, where *=$\mathcal{R}^{*}=\{R_{(A,p\beta)},R_{(A,a\beta)},R_{(A,m\beta)},R_{(A,s\beta)},\newline R_{(A,n\beta)}\}$ is illustrated below Definition 5.48, $\{Sim\}$ is a family of similar degree functions and $\{Sim\}=\{Sim^{p\beta}(\cdot),Sim^{a\beta}(\cdot),Sim^{m\beta}(\cdot),Sim^{s\beta}(\cdot),Sim^{n\beta}(\cdot)\}$ that is illustrated in the Definition 5.49.
5.2.5 An example of the multi-strength intelligent classifier application
Monitoring data of large wind turbines of the same model with different health states are used to do case analysis as follows. For the raw data of devices, let $U=\{x_{1},x_{2},\cdots,x_{|U|}\}$ be a sample set and $data(e)(x_{i})$ be the raw data of $x_{i}\in U$ with respect to the attribute $e$. Let $data(e)=\sqcup_{x_{i}\in U}data(e)(x_{i})$ and $h=\frac{d-min(data(e))}{max(data(e))-min(data(e))}$, where $d\in data(e)(x_{i})$ and $h\in F(e)(x_{i})$. Then, we can convert the raw data into the data conformed to the value range of the hesitant fuzzy membership degree.
Example 5.51.
$U=\{x_{1},x_{2},\cdots,x_{17}\}$ is a sample set. $E=\{e_{1},e_{2},\cdots,e_{26}\}$ is an attribute set. $PT$, $NT$ and $AcT$ are set to 85%, 50% and 95%, respectively.
The samples are classified based on different health states $U_{n}=\{x_{1},x_{6},x_{13}\}$, $U_{s}=\{x_{2},x_{12},x_{14}\}$, $U_{m}=U_{a}=\{x_{3},x_{5},x_{15}\}$ (the samples $x_{3}$, $x_{5}$ and $x_{15}$ are in a same health grade G3), $U_{p}=\{x_{7},x_{8},x_{11},x_{16}\}$ and $U_{f}=\{x_{4},x_{9},x_{10},x_{17}\}$. About two-thirds of the samples from various health states are taken as training samples, and the remaining samples are used as testing samples.
In the last iteration, $U^{*}_{tain}=\{x_{2},x_{3},x_{4},x_{6},x_{7},x_{9},\newline x_{11},x_{13},x_{14},x_{15},x_{16},x_{17}\}$, $U_{test}=\{x_{1},x_{5},x_{8},x_{10},x_{12}\}$ and $A^{*}=\{e_{1},e_{2},e_{6},e_{7},e_{8},e_{11},e_{14},e_{17},e_{18},e_{19},e_{21},e_{26}\}$, shown as the Table 3 in Appendix. In the following, we demonstrate some numerical calculations of Algorithm 2 with the results of the last iteration.
In $(U^{*}_{tain},\mathcal{R}^{*},\{Sim\})$, $\xi^{n\beta}(x_{6})=(\neg e_{1},e_{2},e_{6},e_{7},e_{8},\newline \neg e_{11},\neg e_{14},\neg e_{17},e_{18},e_{19},\neg e_{21},\neg e_{26})$ and $\xi^{n\beta}(x_{13})=(\neg e_{1},e_{2},e_{6},e_{7},e_{8},\neg e_{11},\neg e_{14},e_{17},e_{18},e_{19},\neg e_{21},\neg e_{26})$. Then, $Sim^{n\beta}(x_{6},x_{13})=\frac{|\xi^{n\beta}(x_{6})\sqcap\xi^{n\beta}(x_{13})|}{|A^{*}|}=\frac{11}{12}$. Then, $\frac{1}{C_{|U^{*}_{n}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})=\frac{11}{12}>PN=85\%$.
In the same manner, to calculate the other factors in the inequality array (1), we can verify that $U^{*}_{tain}$, $A^{*}$ and $\beta^{*}$ satisfy the other conditions of the inequality array (1). Then, we classify the samples in $U_{test}$ to calculate the classification accuracy.
For $x_{1}\in U_{test}$, $Cl(x_{1})=\{Cl_{n}(x_{1}),Cl_{s}(x_{1}),Cl_{m}(x_{1}),\newline Cl_{a}(x_{1}),Cl_{p}(x_{1}),Cl_{f}(x_{1})\}=\{0.9583,0.9583,0.7083,\newline 0.7083,0.4722,0.0556\}$, $x_{1}$ is assigned to $U_{n}$ (health grade is G1) and classified correctly. In the same manner, we can obtain $x_{5}\in U_{m}=U_{a}$ (G3), $x_{8}\in U_{p}$ (G4), $x_{10}\in U_{f}$ (G5), and $x_{12}\in U_{s}$ (G2). All five testing samples are correctly classified to their classes, then $AC=\frac{5}{|U_{test}|}=100\%>95\%$, the intelligent classifier passes the test. And the intelligent classifier $(U^{*}_{tain},\mathcal{R}^{*},\{Sim\})$ is stored in the knowledge base $(U^{*}_{tain},\mathcal{R}^{*})$.
When we obtain a new sample $x_{new}$, we can calculate $Cl(x_{new})$ and classify it by using the multi-strength intelligent classifier $(U^{*}_{tain},\mathcal{R}^{*},\{Sim\})$ like the way above.
Regarding attribute $e$, the larger the value of the hesitant fuzzy membership, the higher the level of the system health status, that is, the system health status is positively correlated with attribute $e$. Regarding attribute $e$, the smaller the value of hesitant fuzzy membership is, the higher the level of system health status is, that is, the system health status is negatively correlated with attribute $e$. The advantage of the proposed multi-strength intelligent classifier: training classifier does not require analyzing the positive or negative correlation between attributes and the health state of the system before.
If there are some attributes that are not positively or negatively correlated with the health state of the system, these attributes will affect the similarity functions of the classifier. The classifier can eliminate these attributes during the training processes to satisfy the constraints of the inequality array (1). This means that the classifier can reduce the dimensionality of high-dimensional data, removing attributes that are not relevant to the result. If the objects in $U$ are candidates for investment, the multi-strength intelligent classifier can evaluate the projects and select the optimal investment objects, and the classifier can help us make investment decisions.
In summary, this multi-strength intelligent classifier is a candidate tool for dimensionality reduction of high-dimensional data, classification, diagnosis, evaluation, decision-making, and so on.
6 Conclusion
The definition of inclusion relationship is one of the most foundational concepts of sets. Based on the hesitant fuzzy membership degree of discrete form, we propose several kinds of inclusion relationships for hesitant fuzzy sets. Following the definitions of several kinds of inclusion relationships, some foundational propositions of hesitant fuzzy sets and the families of hesitant fuzzy sets are investigated. Some rules hold in classical sets but not in hesitant fuzzy sets are also shown. In subsequent studies of hesitant fuzzy sets, researchers must avoid the intuitions of classical sets and cannot use these rules.
Furthermore, we investigate some foundational theorems of hesitant fuzzy information systems with respect to parameter reductions and present an example and an algorithm in hesitant fuzzy information systems to illustrate the processes of parameter reductions. Finally, we propose a multi-strength intelligent classifier to make health state diagnoses for complex systems.
In our future works, we will investigate more applications of the multi-strength intelligent classifiers, saturation reductions in hesitant fuzzy information systems, and foundational theories of knowledge bases with different strengths of knowledge, among other topics.
Acknowledgments
The authors thank the editors and the anonymous reviewers for their helpful comments and suggestions that have led to this improved version of the paper. The work was supported by the National Natural Science Foundations of China (No. 71904043).

### v4

@@
-Foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems and their applications for multi-strength intelligent classifiers
+Foundational theories of hesitant fuzzy sets and families of hesitant fuzzy sets
 Abstract
-Hesitant fuzzy sets are widely used in certain instances of uncertainty and hesitation.
-In sets, the inclusion relationship is an important and foundational definition.
-Thus, as a kind of set, hesitant fuzzy sets require an explicit definition of inclusion relationship.
-Based on the hesitant fuzzy membership degree of discrete form, several kinds of inclusion relationships for hesitant fuzzy sets are proposed in this work.
-Then, some foundational propositions of hesitant fuzzy sets are presented, along with propositions of families of hesitant fuzzy sets.
-Some foundational propositions of hesitant fuzzy information systems are proposed with respect to parameter reductions and an example and an algorithm are given to illustrate the processes of parameter reduction.
-Finally, a multi-strength intelligent classifier is proposed to make health state diagnoses for complex systems.
+Hesitant fuzzy sets find extensive application in specific scenarios involving uncertainty and hesitation.
+In the context of set theory, the concept of inclusion relationship holds significant importance as a fundamental definition.
+Consequently, as a type of sets, hesitant fuzzy sets necessitate a clear and explicit definition of the inclusion relationship.
+Based on the discrete form of hesitant fuzzy membership degrees, this study proposes multiple types of inclusion relationships for hesitant fuzzy sets.
+Subsequently, this paper introduces foundational propositions related to hesitant fuzzy sets, as well as propositions concerning families of hesitant fuzzy sets.
 keywords:
-Hesitant fuzzy sets, hesitant fuzzy information systems, parameter reductions, multi-strength intelligent classifier
+Hesitant fuzzy sets , inclusion relationship , families of hesitant fuzzy sets
 1 Introduction
-Set theory established by Cantor FP (hereafter referred to as classical set) is the foundation of modern mathematics.
-Meanwhile, modern mathematics, such as group theory EJB ; DDJ , topology SWA , graph theory GHH and so on, are the foundations of computer science.
-After classical set theory was perfected, other kinds of sets were developed, such as fuzzy set ZADE , interval-valued fuzzy set GOMB , intuitionistic fuzzy set AKT and hesitant fuzzy set TV and so on.
-Several kinds of sets had developed their foundational theoretical systems except hesitant fuzzy set.
-Especially, one of the most foundational definitions, the definition of the inclusion relationship between two hesitant fuzzy sets, is not sufficiently clear.
-This article develops the foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems, and proposes a multi-strength intelligent classifier to make health state diagnoses for complex systems.
-For the classical set FP , a set describes the membership degree of an element in two cases, namely, 0% and 100%.
-The relationship between an element and a set is determined depending on the element in or not in the set.
-However, there are uncertainties in the real world.
-To deal with the uncertain cases, Zadeh proposed the concept of fuzzy set ZADE wherein the membership degree of an element of a set is a value in $[0,1]$.
-The membership degree of a fuzzy set represents the probability that an element belongs to a set TZH ; YSB and is no longer just 0% and 100%.
-Subsequently, Gorzalczany GOMB proposed the concept of interval-valued fuzzy set in which the membership degree of an element of a set is an interval included in $[0,1]$.
-Interval-valued fuzzy set can deal with cases wherein we cannot determine the exact probability of an element belonging to a set, but we can determine an interval that the probability must appear in this interval BSO ; FKM .
-Furthermore, Atanassov AKT proposed the concept of intuitionistic fuzzy set wherein the membership degree of an element of a set is a bibasic array.
-Intuitionistic fuzzy set can describe things with a dual character, one number of the bibasic array expresses the positive aspect and the other number expresses the negative aspect KRA ; PLDY .
-Torra TV proposed the concept of hesitant fuzzy set wherein the membership degree of an element of a set is a multidimensional array.
-Hesitant fuzzy sets focus on cases of uncertainty and hesitation, which are common in the real world.
-For instance, about the data sampling for equipment monitoring and decision scoring by an expert team, the feedback of such cases always comes in the form of arrays, and the multiple different values can make us indecisive.
-Thus, hesitant fuzzy sets with a membership degree of a discrete array are more suitable in describing the indecisiveness in decisions compared with the classical, fuzzy, interval-valued fuzzy, and intuitionistic fuzzy sets.
-Hesitant fuzzy sets are widely used in different fields, such as decision-making HJH ; LLYX ; LNZQY ; GXLY , risk evaluation HXZS , agglomerative hierarchical clustering JLGL , investment policy DYM , modeling for complex product manufacturing tasks BLYY , linguistic perceptual AWNJE , forecasting HLJW ; PMNME , and so on.
-However, there is no sufficiently clear definition of the inclusion relationship between two hesitant fuzzy sets.
-The definition of inclusion relationship is a significant foundation for sets, because if a set is a subset of another set and vice versa, then these two sets are equivalent.
-Based on the hesitant fuzzy membership degree of discrete form, we propose in this work several kinds of inclusion relationships for hesitant fuzzy sets and then present some foundational propositions of hesitant fuzzy sets and the families of hesitant fuzzy sets.
-It is worth noting that some foundational propositions of classical sets do not hold in hesitant fuzzy sets.
-In recent years, information systems with uncertain information have become a hot topic due to the rising instances of uncertain information in the real world.
-Plenty of studies have been conducted on this topic, such as fuzzy information system NXSZ , fuzzy incomplete information system WWZJ , interval-valued information system DJWZ ; SDZX , and intuitionistic fuzzy information system HBLH .
-Based on the foundational propositions of hesitant fuzzy sets, we then study some theories with respect to parameter reductions in hesitant fuzzy information systems.
-Analyzing high-dimensional data with small sample sizes is a challenge, the health status of a complex system with such data is difficult to diagnose.
-Based on the foundational propositions of hesitant fuzzy information systems, a multi-strength intelligent classifier in knowledge bases PALA ; PASA is proposed to make health state diagnoses for complex systems.
-This multi-strength intelligent classifier is a candidate tool for dimensionality reduction of high-dimensional data, classification, diagnosis, evaluation, decision-making, and so on.
-The remainder of the paper is constructed as follows.
-Section 2 presents several kinds of inclusion relationships of hesitant fuzzy sets, while some foundational propositions of hesitant fuzzy sets are discussed, showing some rules that hold in classical sets or hesitating fuzzy sets.
-Section 3 presents some propositions on families of hesitant fuzzy sets.
-Section 4 proposes some propositions with respect to parameter reductions in hesitant fuzzy information systems and provides an algorithm for parameter reductions.
-Section 5 proposes a multi-strength intelligent classifier in knowledge bases to make health state diagnoses for complex systems.
-Section 6 presents a conclusion.
+The foundation of modern mathematics lies in set theory established by Cantor [1] (referred to as classical sets hereafter).
+Moreover, modern branches of mathematics, including group theory [2, 3], topology [4], graph theory [5], among others, form the underpinnings of computer science.
+Subsequent to the refinement of classical set theory, various types of sets emerged, encompassing fuzzy sets [6, 7], interval-valued fuzzy sets [8], intuitionistic fuzzy sets [9], and hesitant fuzzy sets [10], among others.
+With the exception of hesitant fuzzy sets, several types of sets have established their foundational theoretical frameworks.
+Notably, the definition of the inclusion relationship between two hesitant fuzzy sets, which is one of the fundamental definitions, lacks adequate clarity.
+This article develops the foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems, and introduces a multi-strength intelligent classifier for diagnosing the health states of complex systems.
+In the case of the classical set [1], the membership degree of an element is described as either 0% or 100%, and the element’s relationship with the set is determined by whether it belongs to the set or not.
+Nevertheless, the real world is fraught with uncertainties.
+To address such uncertain scenarios, Zadeh introduced the concept of fuzzy sets [6], where the membership degree of an element in a set is a value ranging from 0 to 1.
+The membership degree of a fuzzy set denotes the probability of an element belonging to the set [11], thereby extending beyond the binary values of 0% and 100%.
+Later, Gorzalczany [8] introduced the notion of interval-valued fuzzy sets, where the membership degree of an element in a set is represented by an interval included in $[0,1]$.
+Interval-valued fuzzy sets can handle situations where the precise probability of an element belonging to a set cannot be determined, but an interval can be identified within which the probability must lie [12].
+Additionally, Atanassov [9] introduced the notion of intuitionistic fuzzy sets, where the membership degree of an element in a set is represented by a bipartite array.
+Intuitionistic fuzzy sets can capture entities with dual characteristics, with one value in the bipartite array representing the positive aspect and the other value representing the negative aspect [13, 14].
+Torra [10] introduced the concept of hesitant fuzzy sets, where the membership degree of an element in a set is represented by a multidimensional array.
+Hesitant fuzzy sets are designed to address situations of uncertainty and hesitation, which are prevalent in real-world scenarios.
+For example, in scenarios involving data sampling for equipment monitoring and decision scoring by expert teams, the feedback often manifests as arrays with multiple diverse values, leading to indecisiveness.
+Consequently, hesitant fuzzy sets with a membership degree represented by a discrete array are better suited for capturing the indecisiveness in decision-making compared to classical, fuzzy, interval-valued fuzzy, and intuitionistic fuzzy sets.
+Hesitant fuzzy sets find extensive applications in various domains, including decision-making [15, 16], attribute reduction [17], classification [18], linguistic perceptual [19], and forecasting [20], among others.
+Nevertheless, a sufficiently clear definition of the inclusion relationship between two hesitant fuzzy sets is currently lacking.
+The definition of the inclusion relationship serves as a crucial foundation for sets, as it establishes equivalence between two sets when one is a subset of the other and vice versa.
+Based on the discrete form of hesitant fuzzy membership degrees, this study introduces multiple types of inclusion relationships for hesitant fuzzy sets and subsequently presents foundational propositions concerning hesitant fuzzy sets and their families.
+It is important to highlight that some foundational propositions applicable to classical sets do not hold in the case of hesitant fuzzy sets.
+The remainder of this paper is organized as follows.
+Section 2 introduces various types of inclusion relationships among hesitant fuzzy sets, along with their foundational propositions, and shows whether certain rules that apply to classical sets also hold for hesitant fuzzy sets.
+In Section 3, we put forth propositions concerning families of hesitant fuzzy sets.
+Section 4 provides a conclusion.
 2 Foundations of hesitant fuzzy sets
-In this section, we briefly recall some basic concepts about hesitant fuzzy sets and then present several kinds of inclusion relationships of hesitant fuzzy sets that are suitable for hesitant fuzzy membership degree of discrete form.
-We focus on whether some rules of classical sets hold in hesitant fuzzy sets.
-In subsequent studies of hesitant fuzzy sets, researchers must avoid the intuitions of classical sets as well as the rules held in the classical sets but not in the hesitant fuzzy sets.
+In this section, we provide a brief overview of fundamental concepts related to hesitant fuzzy sets and introduce various types of inclusion relationships that are applicable to the discrete form of hesitant fuzzy membership degrees.
+Our focus is on examining whether certain rules that apply to classical sets also hold for hesitant fuzzy sets.
+In further research on hesitant fuzzy sets, it is crucial for researchers to refrain from relying on the intuitions derived from classical sets and to be mindful of the rules that are valid in classical sets but not in hesitant fuzzy sets.
 If two sets $A$ and $B$ are classical sets, $A\sqcap B$ and $A\sqcup B$ represent the intersection and union of $A$ and $B$, respectively.
@@
 Let $U$ be a universal set and $E$ be a set of parameters.
-Let $HF(U)$ be the set of all hesitant fuzzy sets defined over $U$.
 2.1 Reviews of fuzzy sets and hesitant fuzzy sets
 Definition 2.1.
-ZADE A fuzzy set $F$ on $U$ is a mapping $F:U\rightarrow[0,1]$.
+[6] A fuzzy set $F$ on $U$ is a mapping $F:U\rightarrow[0,1]$.
 Definition 2.2.
-ZADE $F_{1}$ and $F_{2}$ are two fuzzy sets on $U$, $F_{1}$ is a subset of $F_{2}$ if $F_{1}(x)\leqslant F_{2}(x)$ for all $x\in U$, denoted as $F_{1}\subset F_{2}$.
+[6] $F_{1}$ and $F_{2}$ are two fuzzy sets on $U$, $F_{1}$ is a subset of $F_{2}$ if $F_{1}(x)\leqslant F_{2}(x)$ for all $x\in U$, denoted as $F_{1}\subset F_{2}$.
 Definition 2.3.
-TV A hesitant fuzzy set on $U$ is defined as a function that when applied to $U$ returns a subset of $[0,1]$.
+[21] A hesitant fuzzy element is a non-empty, finite subset of $[0,1]$.
 Definition 2.4.
-TV For each $x\in U$, and a hesitant fuzzy set $H$, the lower bound and upper bound of $H(x)$ are defined
-lower bound $H^{-}(x)=min\ H(x)$,
-upper bound $H^{+}(x)=max\ H(x)$.
+[10] A hesitant fuzzy set on $U$ is defined as a function that when applied to $U$ returns a subset of $[0,1]$.
+In the following, $HF(U)$ denotes the set of all hesitant fuzzy sets defined over $U$.
 Definition 2.5.
-TV Given two hesitant fuzzy sets represented by their membership functions $H_{1}$ and $H_{2}$, their union and intersection are defined
-union $(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant max(H_{1}^{-}(x),H_{2}^{-}(x))\}$,
-intersection $(H_{1}\cap H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\leqslant min(H_{1}^{+}(x),H_{2}^{+}(x))\}$.
-Let $U=\{x,y\}$, there are three common expression forms of hesitant fuzzy set, such as $A=\{\frac{0.6,0.5,0.3}{x},\frac{0.5,0.3,0.2}{y}\}$, $A=\frac{0.6,0.5,0.3}{x}+\frac{0.5,0.3,0.2}{y}$ and $A=\{<x,(0.6,0.5,0.3)>,<y,(0.5,0.3,0.2)>\}$.
-$A=B$ means that $A$ and $B$ are perfectly consistent in this article, i.e., $B=\frac{0.3,0.6,0.5}{x}+\frac{0.2,0.5,0.3}{y}$.
-2.2 Inclusion relationships of hesitant fuzzy sets
-For two hesitant fuzzy sets $H_{1}$ and $H_{2}$, Babitha et al BJ introduced that $H_{1}$ is a hesitant fuzzy subset of $H_{2}$ if $H_{1}(x)\subset H_{2}(x)$ for all $x\in U$.
-However, what is the meaning of $H_{1}(x)\subset H_{2}(x)$?
-There is no detailed description in BJ .
-Carlos et al JCVT introduced a definition of inclusion relationship for hesitant fuzzy sets, i.e., $H_{1}\subset H_{2}$Leftrightarrow $H_{1}\cup H_{2}=H_{2}$.
-By the description in JCVT , $H_{1}(x)\subset H_{2}(x)\Leftrightarrow\{h:h\in H_{2}(x)\}=H_{2}(x)=(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant\max(H_{1}^{-}(x),H_{2}^{-}(x))\}$, i.e., $H_{1}(x)\subset H_{2}(x)$ implies $h<H_{2}^{-}(x)$ for all $h\in H_{1}(x)$.
-The definition of inclusion relationship for hesitant fuzzy sets in JCVT is not perfect enough for these aspects, shown as (i) and (ii),
+[10] For each $x\in U$, and a hesitant fuzzy set $H$, the lower bound and upper bound of $H(x)$ are defined
+lower bound $H^{-}(x)=inf\ H(x)$,
+upper bound $H^{+}(x)=sup\ H(x)$.
+Definition 2.6.
+[10] Given two hesitant fuzzy sets represented by their membership functions $H_{1}$ and $H_{2}$, their union and intersection are defined
+union $(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant sup(H_{1}^{-}(x),H_{2}^{-}(x))\}$,
+intersection $(H_{1}\cap H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\leqslant inf(H_{1}^{+}(x),H_{2}^{+}(x))\}$.
+Definition 2.7.
+[10] For $x\in U$ and $H\in HF(U)$, the complement of $H$ is denoted as $H^{c}$, where
+$H^{c}(x)=\sqcup_{\gamma\in H(x)}\{1-\gamma\}$.
+Theorem 2.8.
+[10, 21] The following statements hold for $A,B,C\in HF(U)$,
+(1) $(A^{c})^{c}=A$.
+(2) $(A\cap B)^{c}=A^{c}\cup B^{c}$.
+(3) $(A\cup B)^{c}=A^{c}\cap B^{c}$.
+(4) $A\cap B=B\cap A$, $A\cup B=B\cup A$.
+(5) $(A\cap B)\cap C=A\cap(B\cap C)$, $(A\cup B)\cup C=A\cup(B\cup C)$.
+Let $U=\{x,y\}$, $A(x)=\{0.6,0.5,0.3\}$ and $A(y)=\{0.5,0.3,0.2\}$. There are three common expression types for the hesitant fuzzy set $A$, shown as follows,
+[…]
+Furthermore, $A=B$ means that $A(x)$ and $B(x)$ are perfectly consistent for each $x\in U$.
+For example, $B=\frac{\{0.3,0.6,0.5\}}{x}+\frac{\{0.2,0.5,0.3\}}{y}$, then $B=A$; $C=\frac{\{0.3,0.3,0.6,0.5\}}{x}+\frac{\{0.2,0.5,0.3\}}{y}$, then $C\neq A$.
+2.2 The proposed inclusion definitions of hesitant fuzzy sets and their relationships
+For two hesitant fuzzy sets $H_{1}$ and $H_{2}$, Babitha et al [22] defined that $H_{1}$ is a hesitant fuzzy subset of $H_{2}$ if $H_{1}(x)\subset H_{2}(x)$ for all $x\in U$.
+However, the reference [22] lacks a detailed description of $H_{1}(x)\subset H_{2}(x)$.
+For two fuzzy sets $F_{1}$ and $F_{2}$, we have $F_{1}\subset F_{2}\Leftrightarrow F_{1}\cup F_{2}=F_{2}$.
+Carlos et al [23] introduced a definition of inclusion relationship for hesitant fuzzy sets, i.e., $H_{1}\subset H_{2}$Leftrightarrow $H_{1}\cup H_{2}=H_{2}$.
+This definition (in [23]) inherits the idea of fuzzy sets, which is convenient for scholars to refer to the existing theory of fuzzy sets to study hesitant fuzzy sets.
+However, we hold two different views on this definition (in [23]), shown as (i) and (ii),
 (i) The inclusion relationship of sets is also the ordering relationship of sets.
 A foundational and important proposition of inclusion relationship, $A\subset B$ and $B\subset A$ if and only if $A=B$, i.e., the equivalent sets contain each other.
-However, by the definition of inclusion relationship in JCVT , if $H_{1}(x)\subset H_{2}(x)$ and $H_{2}(x)\subset H_{1}(x)$, then a contradiction is produced, i.e., $H_{1}^{+}(x)<H_{2}^{-}(x)<H_{2}^{+}(x)<H_{1}^{-}(x)<H_{1}^{+}(x)$.
-(ii) The definition of inclusion relationship in JCVT is too specific to have a small range of applications.
+We assume that $H_{1}=H_{2}$ means that $H_{1}(x)$ and $H_{2}(x)$ are perfectly consistent for each $x\in U$.
+By the description in [23], $H_{1}(x)\subset H_{2}(x)\Leftrightarrow\{h:h\in H_{2}(x)\}=H_{2}(x)=(H_{1}\cup H_{2})(x)=\{h\in H_{1}(x)\sqcup H_{2}(x):h\geqslant sup(H_{1}^{-}(x),H_{2}^{-}(x))\}$, i.e., $H_{1}(x)\subset H_{2}(x)$ implies $h<H_{2}^{-}(x)$ for all $h\in H_{1}(x)$.
+If $H_{1}(x)\subset H_{2}(x)$ and $H_{2}(x)\subset H_{1}(x)$, then a contradiction is produced, i.e., $H_{1}^{+}(x)<H_{2}^{-}(x)\leqslant H_{2}^{+}(x)<H_{1}^{-}(x)\leqslant H_{1}^{+}(x)$.
+(ii) The definition of inclusion relationship in [23] is too specific to have a small range of applications.
 For example, let $H_{1}(x)=\{0.1,0.2,0.5\}$, $H_{2}(x)=\{0.6,0.7,0.9\}$ and $H_{3}(x)=\{0.1,0.5,0.8\}$, then $H_{1}(x)\subset H_{2}(x)$ is obvious.
 However, this definition cannot describe the relationship of $H_{2}(x)$ and $H_{3}(x)$ and is not applicable for many cases of hesitant fuzzy sets.
-In this work, we propose several kinds of inclusion relationships for hesitant fuzzy sets based on the strength of the information in hesitant fuzzy membership degrees.
-First, we present a detailed example to illustrate the Definition 2.7.
+This study introduces various types of inclusion relations for hesitant fuzzy sets determined by the strength of the information contained in hesitant fuzzy membership degrees.
+Initially, a comprehensive example is provided to clarify Definition 2.10.
 Example 2.9.
 Let $U=\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\}$ be a set of decision-making schemes and $H$ be an expert team that consists of three experts.
-$H(U)=\frac{0.9,0.2}{x_{1}}+\frac{0.6,0.6,0.5}{x_{2}}+\frac{0.7,0.5,0.5}{x_{3}}+\frac{0.8,0.6,0.5}{x_{4}}+\frac{0.9,0.3,0.1}{x_{5}}+\frac{0.9,0.8,0.7}{x_{6}}$ are the schemes’ scores that are given by experts, in which the scores for $x_{1}$ are 0.9 and 0.2 that are obtained through the evaluations made by experts.
+$H(U)=\frac{\{0.9,0.2\}}{x_{1}}+\frac{\{0.6,0.6,0.5\}}{x_{2}}+\frac{\{0.7,0.5,0.5\}}{x_{3}}+\frac{\{0.8,0.6,0.5\}}{x_{4}}+\frac{\{0.9,0.3,0.1\}}{x_{5}}+\frac{\{0.9,0.8,0.7\}}{x_{6}}$ are the estimated values of schemes provided by experts, in which the estimated values for $x_{1}$ are 0.9 and 0.2 that are obtained through the evaluations made by experts.
 One of the three experts fails to evaluate the scheme $x_{1}$.
-(1) Here, $x_{1}$ has a score of 0.9, which is greater than or equal to all the scores for $x_{2}$, it is possible that the scheme $x_{1}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{p}H(x_{1})$.
+(1) Here, $x_{1}$ has an estimated value of 0.9, which is greater than or equal to all the estimated values for $x_{2}$, it is possible that the scheme $x_{1}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{p}H(x_{1})$.
 (2) Here, $0.5670.55=mean[H(x_{1})]<mean[H(x_{2})]=$ 0.567, where $mean[\cdot]$ is the mean value operator.
-In comparing the mean values of scores, we find that the scheme $x_{2}$ is better than the scheme $x_{1}$, denoted as $H(x_{1})\subset_{m}H(x_{2})$.
-(3) On the one hand, the best score of $x_{3}$ is 0.7, which is greater than or equal to the best score of $x_{2}$.
-On the other hand, the worst score of $x_{3}$ is 0.5, which is greater than or equal to the worst score of $x_{2}$.
+In comparing the mean values of estimated values, we find that the scheme $x_{2}$ is better than the scheme $x_{1}$, denoted as $H(x_{1})\subset_{m}H(x_{2})$.
+(3) On the one hand, the best estimated value of $x_{3}$ is 0.7, which is greater than or equal to the best estimated value of $x_{2}$.
+On the other hand, the worst estimated value of $x_{3}$ is 0.5, which is greater than or equal to the worst estimated value of $x_{2}$.
 To compare the respective best and worst cases of schemes $x_{2}$ and $x_{3}$, it is acceptable that the scheme $x_{3}$ is better than the scheme $x_{2}$, denoted as $H(x_{2})\subset_{a}H(x_{3})$.
-(4) To compare the scores for schemes $x_{3}$ and $x_{4}$ one by one ($0.7\leqslant 0.8$; $0.5\leqslant 0.6$; and $0.5\leqslant 0.5$), it is strongly credible that the scheme $x_{4}$ is better than the scheme $x_{3}$, denoted as $H(x_{3})\subset_{s}H(x_{4})$.
-(5) We can obtain $H(x_{1})\subset_{s}H(x_{5})$ after truncating the tail score of $x_{5}$, i.e., deleting the score 0.1 of $x_{5}$, this case is denoted as $H(x_{1})\subset_{st}H(x_{5})$ and is recorded briefly as $H(x_{1})\subset_{t}H(x_{5})$.
-(6) The worst score of $x_{6}$ is greater than or equal to the best score of $x_{3}$.
+(4) To compare the estimated values for schemes $x_{3}$ and $x_{4}$ one by one ($0.7\leqslant 0.8$; $0.5\leqslant 0.6$; and $0.5\leqslant 0.5$), it is strongly credible that the scheme $x_{4}$ is better than the scheme $x_{3}$, denoted as $H(x_{3})\subset_{s}H(x_{4})$.
+(5) We can obtain $H(x_{1})\subset_{s}H(x_{5})$ after truncating the tail estimated value of $x_{5}$, i.e., deleting the estimated value 0.1 of $x_{5}$.
+This case is denoted as $H(x_{1})\subset_{st}H(x_{5})$ and is recorded briefly as $H(x_{1})\subset_{t}H(x_{5})$.
+(6) The worst estimated value of $x_{6}$ is greater than or equal to the best estimated value of $x_{3}$.
 Thus, it is necessary that the scheme $x_{6}$ is better than the scheme $x_{3}$, which is denoted as $H(x_{3})\subset_{n}H(x_{6})$.
-Comparing the mean value of scores is a common thought for decision-making; however, while doing so, some important information may be lost, such as the best and the worst scores.
+Comparing the mean value of estimated values is a common approach for decision-making; however, while doing so, some important information may be lost, such as the best and the worst estimated values.
 Definition 2.10.
@@
 If $w_{1}^{\prime}\geqslant w_{1}$, $w_{2}^{\prime}\geqslant w_{2}$$\cdots$,$w_{l}^{\prime}\geqslant w_{l}$, then the relationship of $W$ and $W^{\prime}$ is denoted as $W\preccurlyeq W^{\prime}$.
-(2) Let $q$ be a positive integer. To take $q$ larger numbers of $W$ and construct a best $q$-subsequence of $W$, denoted as $W(q)$, i.e., $|W(q)|=q$ and $w\leqslant min(W(q))$ for $w\in W-W(q)$.
+(2) Let $q$ be a positive integer.
+To take $q$ larger numbers of $W$ and construct a best $q$-subsequence of $W$, denoted as $W(q)$, i.e., $|W(q)|=q$ and $w\leqslant inf(W(q))$ for $w\in W-W(q)$.
+For example, let $W=\{0.9,0.8,0.7,0.65,0.6,0.5\}$, when $q=2$, $W(q)=\{0.9,0.8\}$; when $q=3$, $W(q)=\{0.9,0.8,0.7\}$.
 Remark 2.11.
@@
 (2) If $W$ is a subsequence of $H_{2}(x)$, $H_{1}(x)\preccurlyeq W$ and $|H_{1}(x)|=|H_{2}(x)|$, then $H_{1}(x)\subset_{s}H_{2}(x)$.
-(3) Let $q=min\{|H_{1}(x)|,|H_{2}(x)|\}$, $H_{1}(x)(q)$ and $H_{2}(x)(q)$ be the best $q$-subsequences of $H_{1}(x)$ and $H_{2}(x)$, respectively.
+(3) Let $q=inf\{|H_{1}(x)|,|H_{2}(x)|\}$, $H_{1}(x)(q)$ and $H_{2}(x)(q)$ be the best $q$-subsequences of $H_{1}(x)$ and $H_{2}(x)$, respectively.
 Thereafter, $H_{1}(x)(q)\preccurlyeq H_{2}(x)(q)$ if and only if one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.
-(4) Let $K=H_{1}(x)\sqcap H_{2}(x)$, if $max(H_{1}(x)-K)\leqslant min(K)\leqslant max(K)\leqslant min(H_{2}(x)-K)$, then one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.
+(4) Let $K=H_{1}(x)\sqcap H_{2}(x)$, if $sup(H_{1}(x)-K)\leqslant inf(K)\leqslant sup(K)\leqslant inf(H_{2}(x)-K)$, then one of $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ holds.
 Proposition 2.12.
@@
 (9) If $A\subset_{n}B$, then $A^{+}(x)\leqslant B^{-}(x)$ for all $x\in U$. If $|A(x)|\geqslant|B(x)|$, then $A(x)\subset_{s}B(x)$. If $|A(x)|<|B(x)|$, then $A(x)\subset_{t}B(x)$$\blacksquare$
+A brief summary for Proposition 2.12 is shown as follows,
+{``$\subset_{n}"\Rightarrow``\subset_{s}"\Rightarrow``\subset_{a}"\Rightarrow``\subset_{p}",\\ ``\subset_{n}"\Rightarrow``\subset_{s}"\Rightarrow``\subset_{m}",\\ ``\subset_{n}"\Rightarrow``\subset_{t}"\Rightarrow``\subset_{p}".\\ \end{cases}$
+2.3 Some propositions about the intersection and union of two hesitant fuzzy sets
+Let $\bar{A}$ and $\bar{B}$ be two classical sets, we know that $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}$, $\bar{A}\sqcap\bar{B}\sqsubset\bar{B}$, $\bar{A}\sqsubset\bar{A}\sqcup\bar{B}$, $\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$, and $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$.
+However, these rules are not necessarily true for two hesitant fuzzy sets $A$ and $B$ (shown as Propositions 2.13-2.20).
 Proposition 2.13.
-The following statements hold for $A,B,C\in HF(U)$,
-(1) If $A\subset_{p}B$ and $B\subset_{p}C$, then $A\subset_{p}C$.
-(2) If $A\subset_{a}B$ and $B\subset_{a}C$, then $A\subset_{a}C$.
-(3) If $A\subset_{m}B$ and $B\subset_{m}C$, then $A\subset_{m}C$.
-(4) If $A\subset_{s}B$ and $B\subset_{s}C$, then $A\subset_{s}C$.
-(5) If $A\subset_{t}B$ and $B\subset_{t}C$, then $A\subset_{t}C$.
-(6) If $A\subset_{n}B$ and $B\subset_{n}C$, then $A\subset_{n}C$.
+The following statements hold for $A_{1},A_{2}\in HF(U)$,
+(1) $A_{1}\cap A_{2}\subset_{p}A_{i}$ for $i=1,2$.
+(2) $A_{i}\subset_{p}A_{1}\cup A_{2}$ for $i=1,2$.
+(3) $A_{1}\cap A_{2}\subset_{p}A_{1}\cup A_{2}$.
 Proof (1) and (2) are obvious.
-(3) $A\subset_{m}B$ and $B\subset_{m}C$ imply that $mean[A(x)]\leqslant mean[B(x)]\leqslant mean[C(x)]$ for all $x\in U$, where $mean[\cdot]$ is the mean value operator. Then $A\subset_{m}C$.
-(4) $A\subset_{s}B$ and $B\subset_{s}C$ imply that $|A(x)|\geqslant|B(x)|\geqslant|C(x)|$ for all $x\in U$.
-Let $C(x)=\{h^{C}_{1},h^{C}_{2},\cdots,h^{C}_{|C(x)|}\}$, $B(x)=\{h^{B}_{1},h^{B}_{2},\cdots,h^{B}_{|B(x)|}\}$ and $A(x)=\{h^{A}_{1},h^{A}_{2},\cdots,h^{A}_{|A(x)|}\}$ be three descending sequences, since $A(x)\subset_{s}B(x)$ and $B(x)\subset_{s}C(x)$, then $h^{C}_{1}\geqslant h^{B}_{1}\geqslant h^{A}_{1}$, $h^{C}_{2}\geqslant h^{B}_{2}\geqslant h^{A}_{2}$$\cdots$, and $h^{C}_{|C(x)|}\geqslant h^{B}_{|C(x)|}\geqslant h^{A}_{|C(x)|}$.
-Hence, $A\subset_{s}C$.
-(5) $A\subset_{t}B$ and $B\subset_{t}C$ imply that $|A(x)|<|B(x)|<|C(x)|$ for all $x\in U$.
-Let $C(x)=\{h^{C}_{1},h^{C}_{2},\cdots,h^{C}_{|C(x)|}\}$, $B(x)=\{h^{B}_{1},h^{B}_{2},\cdots,h^{B}_{|B(x)|}\}$ and $A(x)=\{h^{A}_{1},h^{A}_{2},\cdots,h^{A}_{|A(x)|}\}$ be three descending sequences, since $A(x)\subset_{t}B(x)$ and $B(x)\subset_{t}C(x)$, then $h^{C}_{1}\geqslant h^{B}_{1}\geqslant h^{A}_{1}$, $h^{C}_{2}\geqslant h^{B}_{2}\geqslant h^{A}_{2}$$\cdots$, and $h^{C}_{|A(x)|}\geqslant h^{B}_{|A(x)|}\geqslant h^{A}_{|A(x)|}$.
-Hence, $A\subset_{t}C$.
-(6) $A\subset_{n}B$ and $B\subset_{n}C$ imply that $A^{+}(x)\leqslant B^{-}(x)\leqslant B^{+}(x)\leqslant C^{-}(x)$ for all $x\in U$, then $A\subset_{n}C$$\blacksquare$
-Let $\bar{A}$ and $\bar{B}$ be two classical sets, we know that $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}$, $\bar{A}\sqcap\bar{B}\sqsubset\bar{B}$, $\bar{A}\sqsubset\bar{A}\sqcup\bar{B}$, $\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$, and $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$.
-However, these rules are not necessarily true for two hesitant fuzzy sets $A$ and $B$ (shown as Propositions 2.11-2.18).
+(3) $inf\{A_{1}^{+}(x),A_{2}^{+}(x)\}\leqslant sup\{A_{1}^{+}(x),A_{2}^{+}(x)\}$$\blacksquare$
 Proposition 2.14.
-The following statements hold for $A,B\in HF(U)$,
-(1) $A\cap B\subset_{p}A$.
-(2) $A\cap B\subset_{p}B$.
-(3) $A\subset_{p}A\cup B$.
-(4) $B\subset_{p}A\cup B$.
-(5) $A\cap B\subset_{p}A\cup B$.
-Proof (1) $min\{A^{+}(x),B^{+}(x)\}\leqslant A^{+}(x)$.
-(2) $min\{A^{+}(x),B^{+}(x)\}\leqslant B^{+}(x)$.
-(3) $A^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$.
-(4) $B^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$.
-(5) $min\{A^{+}(x),B^{+}(x)\}\leqslant max\{A^{+}(x),B^{+}(x)\}$$\blacksquare$
-Proposition 2.12.
-The following statements hold for $A,B\in HF(U)$,
-(1) $A\cap B\subset_{a}A$.
-(2) $A\cap B\subset_{a}B$.
-(3) $A\subset_{a}A\cup B$.
-(4) $B\subset_{a}A\cup B$.
-(5) $A\cap B\subset_{a}A\cup B$.
-Proof (1) $min\{A^{+}(x),B^{+}(x)\}\leqslant A^{+}(x)$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant A^{-}(x)$.
-(2) $min\{A^{+}(x),B^{+}(x)\}\leqslant B^{+}(x)$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant B^{-}(x)$.
-(3) $A^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $A^{-}(x)\leqslant max\{A^{-}(x),B^{-}(x)\}$.
-(4) $B^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $B^{-}(x)\leqslant max\{A^{-}(x),B^{-}(x)\}$.
-(5) $min\{A^{+}(x),B^{+}(x)\}\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant max\{A^{-}(x),B^{-}(x)\}$$\blacksquare$
+The following statements hold for $A_{1},A_{2}\in HF(U)$,
+(1) $A_{1}\cap A_{2}\subset_{a}A_{i}$ for $i=1,2$.
+(2) $A_{i}\subset_{a}A_{1}\cup A_{2}$ for $i=1,2$.
+(3) $A_{1}\cap A_{2}\subset_{a}A_{1}\cup A_{2}$.
+Proof (1) and (2) are obvious.
+(3) $inf\{A_{1}^{+}(x),A_{2}^{+}(x)\}\leqslant sup\{A_{1}^{+}(x),A_{2}^{+}(x)\}$ and $inf\{A_{1}^{-}(x),A_{2}^{-}(x)\}\leqslant sup\{A_{1}^{-}(x),A_{2}^{-}(x)\}$$\blacksquare$
 Proposition 2.15.
@@
 (c1) $A^{+}(x)=B^{+}(x)$.
-$\sum\limits_{h\in(A\cap B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})$.
-$mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=max(\bar{a},\bar{b})$.
+$\sum\limits_{h\in(A\cap B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\leqslant sup(\bar{a},\bar{b})(k_{1}+k_{2})$.
+$mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})\leqslant sup(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=sup(\bar{a},\bar{b})$.
 Then, one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
 (c2) $A^{+}(x)\neq B^{+}(x)$.
-$A^{+}(x)\neq B^{+}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=min(V)$, then $h^{\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$.
-Since $h^{\prime}>max\{h:h\in(A\cap B)(x)\}$ and $|(A\cap B)(x)|+|V|=k_{1}+k_{2}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
-Since $h^{\prime}=min(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
-Then, we can obtain $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=max(\bar{a},\bar{b})$.
+$A^{+}(x)\neq B^{+}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=inf(V)$, then $h^{\prime}>inf(A^{+}(x),B^{+}(x))=sup\{h:h\in(A\cap B)(x)\}$.
+Since $h^{\prime}>sup\{h:h\in(A\cap B)(x)\}$ and $|(A\cap B)(x)|+|V|=k_{1}+k_{2}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
+Since $h^{\prime}=inf(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
+Then, we can obtain $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\leqslant sup(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=sup(\bar{a},\bar{b})$.
 Then, one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
@@
 (c1) $A^{-}(x)=B^{-}(x)$.
-$\sum\limits_{h\in(A\cup B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})$.
-$\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=min(\bar{a},\bar{b})$.
+$\sum\limits_{h\in(A\cup B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\geqslant inf(\bar{a},\bar{b})(k_{1}+k_{2})$.
+$\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})\geqslant inf(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=inf(\bar{a},\bar{b})$.
 Then, one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
 (c2) $A^{-}(x)\neq B^{-}(x)$.
-$A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$. $h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$.
-Since $h^{\prime}=max(V)$ and $|(A\cup B)(x)|+|V|=k_{1}+k_{2}$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
-Since $h^{\prime}<min\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
-Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=min(\bar{a},\bar{b})$.
+$A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=sup(V)$. $h^{\prime}<sup(A^{-}(x),B^{-}(x))=inf\{h:h\in(A\cup B)(x)\}$.
+Since $h^{\prime}=sup(V)$ and $|(A\cup B)(x)|+|V|=k_{1}+k_{2}$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
+Since $h^{\prime}<inf\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
+Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\geqslant inf(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=inf(\bar{a},\bar{b})$.
 Then, one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
@@
 In this case, $|(A\cap B)(x)|<k_{1}+k_{2}$ and $|(A\cup B)(x)|=k_{1}+k_{2}$.
-Let $V=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=min(V)$.
-$h^{\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$.
-Since $|(A\cap B)(x)|+|V|=|(A\cup B)(x)|=k_{1}+k_{2}$ and $h^{\prime}>max\{h:h\in(A\cap B)(x)\}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
-Since $h^{\prime}=min(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})$.
+Let $V=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=inf(V)$.
+$h^{\prime}>inf(A^{+}(x),B^{+}(x))=sup\{h:h\in(A\cap B)(x)\}$.
+Since $|(A\cap B)(x)|+|V|=|(A\cup B)(x)|=k_{1}+k_{2}$ and $h^{\prime}>sup\{h:h\in(A\cap B)(x)\}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
+Since $h^{\prime}=inf(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})$.
 Based on the results above, then $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})=mean[(A\cup B)(x)]$, i.e., $(A\cap B)(x)\subset_{m}(A\cup B)(x)$.
@@
 In this case, $|(A\cap B)(x)|=k_{1}+k_{2}$ and $|(A\cup B)(x)|<k_{1}+k_{2}$.
-Let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$.
-$h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$.
-Since $|(A\cup B)(x)|+|V|=|(A\cap B)(x)|=k_{1}+k_{2}$ and $h^{\prime}<min\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
-Since $h^{\prime}=max(V)$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})$.
+Let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=sup(V)$.
+$h^{\prime}<sup(A^{-}(x),B^{-}(x))=inf\{h:h\in(A\cup B)(x)\}$.
+Since $|(A\cup B)(x)|+|V|=|(A\cap B)(x)|=k_{1}+k_{2}$ and $h^{\prime}<inf\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
+Since $h^{\prime}=sup(V)$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})$.
 Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})=mean[(A\cap B)(x)]$, i.e., $(A\cap B)(x)\subset_{m}(A\cup B)(x)$.
 (c4) $A^{-}(x)\neq B^{-}(x)$ and $A^{+}(x)\neq B^{+}(x)$.
-Based on the condition $A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$, then $h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$, and $|V|+|(A\cup B)(x)|=k_{1}+k_{2}$.
+Based on the condition $A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=sup(V)$, then $h^{\prime}<sup(A^{-}(x),B^{-}(x))=inf\{h:h\in(A\cup B)(x)\}$, and $|V|+|(A\cup B)(x)|=k_{1}+k_{2}$.
 Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
-Based on the condition $A^{+}(x)\neq B^{+}(x)$, let $W=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime\prime}=min(W)$, then $h^{\prime\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$, and $|W|+|(A\cap B)(x)|=k_{1}+k_{2}$.
+Based on the condition $A^{+}(x)\neq B^{+}(x)$, let $W=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime\prime}=inf(W)$, then $h^{\prime\prime}>inf(A^{+}(x),B^{+}(x))=sup\{h:h\in(A\cap B)(x)\}$, and $|W|+|(A\cap B)(x)|=k_{1}+k_{2}$.
 Based on the results above, then $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime\prime}|W|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})<mean[(A\cup B)(x)]$, i.e., $(A\cap B)(x)\subset_{m}(A\cup B)(x)$.
@@
 $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$.
-For random two hesitant fuzzy sets $A$ and $B$, Example 2.14 shows that $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ do not hold simultaneously, $A(y)\subset_{m}(A\cup B)(y)$ and $B(y)\subset_{m}(A\cup B)(y)$ do not hold simultaneously.
+For two random hesitant fuzzy sets $A$ and $B$, Example 2.16 shows that $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ do not hold simultaneously, $A(y)\subset_{m}(A\cup B)(y)$ and $B(y)\subset_{m}(A\cup B)(y)$ do not hold simultaneously.
 $(A\cap B)(x)\subset_{m}B(x)$ and $(A\cap B)(x)\not\subset_{m}A(x)$.
@@
 It means that (1) of Proposition 2.15 cannot be written as that one of $A\cap B\subset_{m}A$ and $A\cap B\subset_{m}B$ holds.
+When we discuss the case ``$\subset_{m}"$ we need to note that
+$\begin{cases}A_{1}\cap A_{2}\subset_{m}A_{i}\ \text{is uncertain for}\ i=1,2,\\ A_{i}\subset_{m}A_{1}\cup A_{2}\ \text{is uncertain for}\ i=1,2.\\ \end{cases}$
+Proposition 2.17.
+The following statements hold for $x\in U$ and $A_{1},A_{2}\in HF(U)$,
+(1) At least one of $A_{i}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{i}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds for $i=1,2$.
+(2) At least one of $(A_{1}\cap A_{2})(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $(A_{1}\cap A_{2})(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.
+Proof (1) We prove the case that one of $A_{1}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{1}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds in follows.
+For each $x\in U$, if $A_{1}^{-}(x)\geqslant A_{2}^{-}(x)$, then $h\in(A_{1}\cup A_{2})(x)$ for all $h\in A_{1}(x)$, i.e., $A_{1}(x)$ is a subsequence of $(A_{1}\cup A_{2})(x)$.
+$A_{1}(x)\preccurlyeq A_{1}(x)$ and $|A_{1}(x)|\leqslant|(A_{1}\cup A_{2})(x)|$, by Remark 2.11 (1) and (2), then one of $A_{1}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{1}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.
+If $A_{1}^{-}(x)<A_{2}^{-}(x)$, let $K=A_{1}(x)\sqcap(A_{1}\cup A_{2})(x)$, then $sup(A_{1}(x)-K)<inf(K)\leqslant sup(K)<sup((A_{1}\cup A_{2})(x)-K)$, by Remark 2.11 (4), then one of $A_{1}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{1}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.
+With the same manner, we can prove the case that one of $A_{2}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{2}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.
+(2) Let $K=(A_{1}\cap A_{2})(x)\sqcap(A_{1}\cup A_{2})(x)$, then $sup((A_{1}\cap A_{2})(x)-K)<inf(K)\leqslant sup(K)<inf((A_{1}\cup A_{2})(x)-K)$, by Remark 2.11 (4), one of $(A_{1}\cap A_{2})(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $(A_{1}\cap A_{2})(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.$\blacksquare$
+If two inclusion relationships $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ in Definition 2.10 are combined in $H_{1}(x)\subset_{sot}H_{2}(x)$ ($\subset_{sot}$ means $\subset_{s}$ or $\subset_{t}$), then Proposition 2.17 can be described as follows:
+Proposition 2.17 The following statements hold for $A_{1},A_{2}\in HF(U)$,
+(1) $A_{i}\subset_{sot}A_{1}\cup A_{2}$ for $i=1,2$.
+(2) $A_{1}\cap A_{2}\subset_{sot}A_{1}\cup A_{2}$.
 Proposition 2.18.
 The following statements hold for $x\in U$ and $A,B\in HF(U)$,
-(1) At least one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
-(2) At least one of $B(x)\subset_{s}(A\cup B)(x)$ and $B(x)\subset_{t}(A\cup B)(x)$ holds.
-(3) At least one of $(A\cap B)(x)\subset_{s}(A\cup B)(x)$ and $(A\cap B)(x)\subset_{t}(A\cup B)(x)$ holds.
-Proof (1) For each $x\in U$, if $A^{-}(x)\geqslant B^{-}(x)$, then $h\in(A\cup B)(x)$ for all $h\in A(x)$, i.e., $A(x)$ is a subsequence of $(A\cup B)(x)$.
-$A(x)\preccurlyeq A(x)$ and $|A(x)|\leqslant|(A\cup B)(x)|$, by Remark 2.8 (1) and (2), then one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
-If $A^{-}(x)<B^{-}(x)$, let $K=A(x)\sqcap(A\cup B)(x)$, then $max(A(x)-K)<min(K)\leqslant max(K)<max((A\cup B)(x)-K)$, by Remark 2.8 (4), then one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
-(2) It can be proved by the similar method of (1).
-(3) Let $K=(A\cap B)(x)\sqcap(A\cup B)(x)$, then $max((A\cap B)(x)-K)<min(K)\leqslant max(K)<min((A\cup B)(x)-K)$, by Remark 2.8 (4), one of $(A\cap B)(x)\subset_{s}(A\cup B)(x)$ and $(A\cap B)(x)\subset_{t}(A\cup B)(x)$ holds.$\blacksquare$
-If two inclusion relationships $H_{1}(x)\subset_{s}H_{2}(x)$ and $H_{1}(x)\subset_{t}H_{2}(x)$ in Definition 2.7 are combined in $H_{1}(x)\subset_{sot}H_{2}(x)$ ($\subset_{sot}$ means $\subset_{s}$ or $\subset_{t}$), then Proposition 2.15 can be described as follows:
-Proposition ${}^{\prime}$ The following statements hold for $A,B\in HF(U)$,
-(1) $A\subset_{sot}A\cup B$.
-(2) $B\subset_{sot}A\cup B$.
-(3) $A\cap B\subset_{sot}A\cup B$.
-Proposition 2.16.
-The following statements hold for $x\in U$ and $A,B,C\in HF(U)$,
 (1) If $A\subset_{p}B$, then at least one of $A(x)\subset_{s}(A\cap B)(x)$ and $A(x)\subset_{t}(A\cap B)(x)$ holds.
@@
 (5) If $A\subset_{n}B$, then at least one of $A(x)\subset_{s}(A\cap B)(x)$ and $A(x)\subset_{t}(A\cap B)(x)$ holds.
-(6) For a random $C\in HF(U)$, if $A\subset_{s}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
-(7) For a random $C\in HF(U)$, if $A\subset_{t}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
-(8) For a random $C\in HF(U)$, if $A\subset_{n}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
 Proof (1) $A\subset_{p}B$ implies that $A^{+}(x)\leqslant B^{+}(x)$ for all $x\in U$.
@@
 (4) $A\subset_{t}B$ implies $A\subset_{p}B$, by the result of (1), then (4) holds.
-(5) $A\subset_{n}B$ implies $A\subset_{p}B$, by the result of (1), then (5) holds.
+(5) $A\subset_{n}B$ implies $A\subset_{p}B$, by the result of (1), then (5) holds.$\blacksquare$
+Proposition 2.18 The following statements hold for $A,B,C\in HF(U)$,
+(1) If $A\subset_{p}B$, then $A\subset_{sot}A\cap B$.
+(2) If $A\subset_{a}B$, then $A\subset_{sot}A\cap B$.
+(3) If $A\subset_{sot}B$, then $A\subset_{sot}A\cap B$.
+(4) If $A\subset_{n}B$, then $A\subset_{sot}A\cap B$.
+Example 2.19.
+Let $U=\{x,y\}$, $A=\frac{\{0.1,0.2,0.5,0.6,0.9\}}{x}+\frac{\{0.1,0.7\}}{y}$, $B=\frac{\{0.05,0.3,0.4,0.7,0.8\}}{x}+\frac{\{0.8,0.9,0.9\}}{y}$.
+$(A\cap B)(x)=\{0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8\}$, then $(A\cap B)(x)\not\subset_{s}A(x)$, $(A\cap B)(x)\not\subset_{s}B(x)$, $(A\cap B)(x)\not\subset_{t}A(x)$, and $(A\cap B)(x)\not\subset_{t}B(x)$.
+$(A\cup B)(y)=\{0.8,0.9,0.9\}$, then $A(y)\subset_{t}(A\cup B)(y)$ and $B(y)\subset_{s}(A\cup B)(y)$.
+Proposition 2.20.
+The following statements hold for $A,B\in HF(U)$,
+(1) If $A\subset_{n}B$, $A\cap B\subset_{n}B$.
+(2) If $A\subset_{n}B$, $A\subset_{n}A\cup B$.
+(3) If $A\subset_{n}B$, $A\cap B\subset_{n}A\cup B$.
+Proof Since $A\subset_{n}B$, then $A^{+}(x)\leqslant B^{-}(x)$ for all $x\in U$, then (1), (2) and (3) are obvious.$\blacksquare$
+Example 2.21.
+Let $U=\{x\}$, $A=\frac{\{0.1,0.3,0.5\}}{x}$, $B=\frac{\{0.2,0.4,0.6\}}{x}$.
+$(A\cap B)(x)=\{0.1,0.2,0.3,0.4,0.5\}$, $(A\cup B)(x)=\{0.2,0.3,0.4,0.5,0.6\}$.
+$A\cap B\not\subset_{n}A$, $A\cap B\not\subset_{n}B$, $A\not\subset_{n}A\cup B$, $B\not\subset_{n}A\cup B$, $A\cap B\not\subset_{n}A\cup B$.
+2.4 Some propositions about a pair of hesitant fuzzy sets (one being a subset of the other) and a random additional hesitant fuzzy set
+For three classical sets $\bar{A}$, $\bar{B}$ and $\bar{C}$, if $\bar{A}\sqsubset\bar{B}$, then $\bar{A}\sqsubset\bar{B}\sqcup\bar{C}$, $\bar{A}\sqcap\bar{C}\sqsubset\bar{B}\sqcap\bar{C}$ and $\bar{A}\sqcup\bar{C}\sqsubset\bar{B}\sqcup\bar{C}$.
+These rules in the cases of three hesitant fuzzy sets $A$, $B$ and $C$ are investigated, shown as Proposition 2.22 and 2.24.
+Proposition 2.22.
+The following statements hold for $x\in U$ and $A,B,C\in HF(U)$,
+(1) If $A\subset_{p}B$, then $A\subset_{p}B\cup C$.
+(2) If $A\subset_{a}B$, then $A\subset_{a}B\cup C$.
+(3) If $A\subset_{s}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
+(4) If $A\subset_{t}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
+(5) If $A\subset_{n}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
+Proof (1) and (2) are obvious.
 (3) We prove this proposition after dividing the relationships of $h\in C(x)$ and the interval $[B^{-}(x),B^{+}(x)]$ into two cases (i) and (ii) for each $x\in U$.
@@
 For $o_{j}\in\{o_{b+1},o_{b+2},\cdots,o_{a}\}\sqsubset A(x)$ and all $h\in(B\cup C)(x)$, we have $h\geqslant B^{-}(x)=v_{b}\geqslant o_{b}\geqslant o_{b+1}\geqslant o_{j}$, where $j=b+1,b+2,\cdots,a$.
-Based on the illustrations of sub-case $C^{-}(x)\leqslant B^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
+Based on the illustrations of sub-case $C^{-}(x)\leqslant B^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=inf\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.11 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
 When $B^{-}(x)<C^{-}(x)$.
@@
 For $o_{j}\in\{o_{m},o_{m+1},\cdots,o_{a}\}\sqsubset A(x)$ and all $h\in(B\cup C)(x)$, we have $h\geqslant C^{-}(x)=w_{c}>v_{m}\geqslant o_{m}\geqslant o_{j}$, where $j=m,m+1,\cdots,a$.
-Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
+Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=inf\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.11 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
 (4) We prove this proposition after dividing the relationships of $h\in C(x)$ and the interval $[B^{-}(x),B^{+}(x)]$ into two cases (i) and (ii) for each $x\in U$.
@@
 Since $o_{m}\leqslant v_{m}<w_{c}=C^{-}(x)$, for $o_{j}\in\{o_{m},o_{m+1},\cdots,o_{a}\}\sqsubset A(x)$ and all $h\in(B\cup C)(x)$, it implies $h>o_{j}$, where $j=m,m+1,\cdots,a$.
-Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$ is obtained, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
-(8) By Proposition 2.9 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for all $x\in U$. By the results of (6) and (7), (8) holds.$\blacksquare$
-Proposition ${}^{\prime}$ The following statements hold for $A,B,C\in HF(U)$,
-(1) If $A\subset_{p}B$, then $A\subset_{sot}A\cap B$.
-(2) If $A\subset_{a}B$, then $A\subset_{sot}A\cap B$.
-(3) If $A\subset_{sot}B$, then $A\subset_{sot}A\cap B$.
-(4) If $A\subset_{n}B$, then $A\subset_{sot}A\cap B$.
-(5) If $A\subset_{sot}B$, then $A\subset_{sot}B\cup C$ for a random $C\in HF(U)$.
-(6) If $A\subset_{n}B$, then $A\subset_{sot}B\cup C$ for a random $C\in HF(U)$.
-Example 2.17.
-Let $U=\{x,y\}$, $A=\frac{\{0.1,0.2,0.5,0.6,0.9\}}{x}+\frac{\{0.1,0.7\}}{y}$, $B=\frac{\{0.05,0.3,0.4,0.7,0.8\}}{x}+\frac{\{0.8,0.9,0.9\}}{y}$.
-$(A\cap B)(x)=\{0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8\}$, then $(A\cap B)(x)\not\subset_{s}A(x)$, $(A\cap B)(x)\not\subset_{s}B(x)$, $(A\cap B)(x)\not\subset_{t}A(x)$, and $(A\cap B)(x)\not\subset_{t}B(x)$.
-$(A\cup B)(y)=\{0.8,0.9,0.9\}$, then $A(y)\subset_{t}(A\cup B)(y)$ and $B(y)\subset_{s}(A\cup B)(y)$.
+Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$ is obtained, where $q=inf\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.11 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
+(5) By Proposition 2.12 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for all $x\in U$. By the results of (3) and (4) of Proposition 2.22, (5) holds.$\blacksquare$
+(3) and (4) of Proposition 2.22 can be combined as (1) of Proposition 2.22. (5) of Proposition 2.22 can be written into (2) of Proposition 2.22.
+Proposition 2.22 The following statements hold for $A,B,C\in HF(U)$,
+(1) If $A\subset_{sot}B$, then $A\subset_{sot}B\cup C$.
+(2) If $A\subset_{n}B$, then $A\subset_{sot}B\cup C$.
 Proposition 2.23.
-The following statements hold for $A,B\in HF(U)$,
-(1) If $A\subset_{n}B$, $A\cap B\subset_{n}B$.
-(2) If $A\subset_{n}B$, $A\subset_{n}A\cup B$.
-(3) If $A\subset_{n}B$, $A\cap B\subset_{n}A\cup B$.
-Proof Since $A\subset_{n}B$, then $A^{+}(x)\leqslant B^{-}(x)$ for all $x\in U$, then (1), (2) and (3) are obvious.$\blacksquare$
-Example 2.19.
-Let $U=\{x\}$, $A=\frac{\{0.1,0.3,0.5\}}{x}$, $B=\frac{\{0.2,0.4,0.6\}}{x}$.
-$(A\cap B)(x)=\{0.1,0.2,0.3,0.4,0.5\}$, $(A\cup B)(x)=\{0.2,0.3,0.4,0.5,0.6\}$.
-$A\cap B\not\subset_{n}A$, $A\cap B\not\subset_{n}B$, $A\not\subset_{n}A\cup B$, $B\not\subset_{n}A\cup B$, $A\cap B\not\subset_{n}A\cup B$.
-For three classical sets $\bar{A}$, $\bar{B}$ and $\bar{C}$, if $\bar{A}\sqsubset\bar{B}$, then $\bar{A}\sqcap\bar{C}\sqsubset\bar{B}\sqcap\bar{C}$ and $\bar{A}\sqcup\bar{C}\sqsubset\bar{B}\sqcup\bar{C}$.
-These rules in the cases of three hesitant fuzzy sets $A$, $B$ and $C$ are investigated, shown as Proposition 2.20.
+The following statements hold for $A,B,C\in HF(U)$,
+(1) If $A\subset_{p}B$, $A\cap C\subset_{p}B\cap C$.
+(2) If $A\subset_{a}B$, $A\cap C\subset_{a}B\cap C$.
+(3) If $A\subset_{s}B$, $A\cap C\subset_{a}B\cap C$.
+(4) If $A\subset_{t}B$, $A\cap C\subset_{p}B\cap C$.
+(5) If $A\subset_{n}B$, $A\cap C\subset_{a}B\cap C$.
+Proof (1)-(5) are obvious.
 Proposition 2.24.
 The following statements hold for $x\in U$ and $A,B,C\in HF(U)$,
-(1) If $A\subset_{p}B$, $A\cap C\subset_{p}B\cap C$.
 (1) If $A\subset_{p}B$, $A\cup C\subset_{p}B\cup C$.
-(3) If $A\subset_{a}B$, $A\cap C\subset_{a}B\cap C$.
 (2) If $A\subset_{a}B$, $A\cup C\subset_{a}B\cup C$.
-(5) If $A\subset_{s}B$, $A\cap C\subset_{a}B\cap C$.
 (3) If $A\subset_{s}B$, $A\cup C\subset_{a}B\cup C$.
-(7) If $A\subset_{t}B$, $A\cap C\subset_{p}B\cap C$.
 (4) If $A\subset_{t}B$, $A\cup C\subset_{p}B\cup C$.
-(9) If $A\subset_{n}B$, $A\cap C\subset_{a}B\cap C$.
 (5) If $A\subset_{n}B$, $A\cup C\subset_{a}B\cup C$.
@@
 $v_{b}=B^{-}(x)\leqslant(B\cup C)^{-}(x)$, then $(B\cup C)(x)$ is a subsequence of $W^{2}$.
-Let $q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$.
-$w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
+Let $q=inf\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$.
+$w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=inf\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
 By Remark 2.11 (3), one of $(A\cup C)(x)\subset_{s}(B\cup C)(x)$ and $(A\cup C)(x)\subset_{t}(B\cup C)(x)$ holds.
@@
 $o_{a}=A^{-}(x)\leqslant(A\cup C)^{-}(x)$, then $(A\cup C)(x)$ is a subsequence of $W^{1}$.
-Let $q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$.
-$w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
+Let $q=inf\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$.
+$w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=inf\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
 By Remark 2.11 (3), one of $(A\cup C)(x)\subset_{s}(B\cup C)(x)$ and $(A\cup C)(x)\subset_{t}(B\cup C)(x)$ holds.
-(13) By Proposition 2.9 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for $x\in U$. By the results of (11) and (12), then (13) holds.$\blacksquare$
-(11) and (12) of Proposition 2.20 can be combined as (1) of Proposition ${}^{\prime}$. (13) of Proposition 2.20 can be written into (2) of Proposition ${}^{\prime}$.
-Proposition ${}^{\prime}$ The following statements hold for $A,B,C\in HF(U)$,
+(8) By Proposition 2.12 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for $x\in U$. By the results of (6) and (7) of Proposition 2.24, then (8) holds.$\blacksquare$
+(6) and (7) of Proposition 2.24 can be combined as (1) of Proposition 2.24. (8) of Proposition 2.24 can be written into (2) of Proposition 2.24.
+Proposition 2.24 The following statements hold for $A,B,C\in HF(U)$,
 (1) If $A\subset_{sot}B$, then $A\cup C\subset_{sot}B\cup C$.
@@
 (4) $A\subset_{n}B$ and $A\subset_{n}C$ if and only if $A\subset_{n}B\cap C$.
-Proof (1) For all $x\in U$, $A^{+}(x)\leqslant B^{+}(x)$ and $A^{+}(x)\leqslant C^{+}(x)$ if and only if $A^{+}(x)\leqslant\min\{B^{+}(x),C^{+}(x)\}$.
-(2) For all $x\in U$, $A^{-}(x)\leqslant B^{-}(x)$ and $A^{-}(x)\leqslant C^{-}(x)$ if and only if $A^{-}(x)\leqslant\min\{B^{-}(x),C^{-}(x)\}$. To combine the result of (1), then (2) holds.
+Proof (1) For all $x\in U$, $A^{+}(x)\leqslant B^{+}(x)$ and $A^{+}(x)\leqslant C^{+}(x)$ if and only if $A^{+}(x)\leqslant inf\{B^{+}(x),C^{+}(x)\}$.
+(2) For all $x\in U$, $A^{-}(x)\leqslant B^{-}(x)$ and $A^{-}(x)\leqslant C^{-}(x)$ if and only if $A^{-}(x)\leqslant inf\{B^{-}(x),C^{-}(x)\}$. To combine the result of (1), then (2) holds.
 (3) For a random $x\in U$, the relationships of $h\in C(x)$ and interval $[B^{-}(x),B^{+}(x)]$ are divided into two cases (i) and (ii) as follows,
@@
 To sum up, $A(x)\subset_{t}(B\cap C)(x)$ for random $x\in U$, then $A\subset_{t}B\cap C$.
-(4) For all $x\in U$, $A^{+}(x)\leqslant B^{-}(x)$ and $A^{+}(x)\leqslant C^{-}(x)$ if and only if $A^{+}(x)\leqslant\min\{B^{-}(x),C^{-}(x)\}$$\blacksquare$
+(4) For all $x\in U$, $A^{+}(x)\leqslant B^{-}(x)$ and $A^{+}(x)\leqslant C^{-}(x)$ if and only if $A^{+}(x)\leqslant inf\{B^{-}(x),C^{-}(x)\}$$\blacksquare$
 Suppose $A(x)=\{0.3,0.5,0.7\}$, $B(x)=\{0.8,0.9\}$, $C(x)=\{0.6,0.8,0.9\}$. $A(x)\subset_{t}(B\cap C)(x)$, but $A(x)\not\subset_{t}B(x)$. The converse of Proposition 2.26 (3) does not hold.
-Definition 2.23.
-TV For $x\in U$ and $H\in HF(U)$, the complement of $H$ is denoted as $H^{c}$, where
-$H^{c}(x)=\sqcup_{\gamma\in H(x)}\{1-\gamma\}$.
-Theorem 2.24.
-TV ; XX The following statements hold for $A,B,C\in HF(U)$,
-(1) $(A^{c})^{c}=A$.
-(2) $(A\cap B)^{c}=A^{c}\cup B^{c}$.
-(3) $(A\cup B)^{c}=A^{c}\cap B^{c}$.
-(4) $A\cap B=B\cap A$, $A\cup B=B\cup A$.
-(5) $(A\cap B)\cap C=A\cap(B\cap C)$, $(A\cup B)\cup C=A\cup(B\cup C)$.
+2.5 The proposition of the complement of hesitant fuzzy set and the transitivity of the inclusion relationships in hesitant fuzzy sets
 Proposition 2.27.
@@
 Example 2.28.
-Let $U=\{x,y\}$, $A=\frac{0.4,0.4}{x}+\frac{0.2,0.25}{y}$, $B=\frac{0.1,0.1,0.41}{x}+\frac{0.1,0.2,0.3}{y}$.
+Let $U=\{x,y\}$, $A=\frac{\{0.4,0.4\}}{x}+\frac{\{0.2,0.25\}}{y}$, $B=\frac{\{0.1,0.1,0.41\}}{x}+\frac{\{0.1,0.2,0.3\}}{y}$.
 (1) $A(x)\subset_{p}B(x)$. $A^{c}(x)=\{0.6,0.6\}$, $B^{c}(x)=\{0.9,0.9,0.59\}$. $B^{c}(x)\not\subset_{p}A^{c}(x)$, $B^{c}(x)\not\subset_{m}A^{c}(x)$.
@@
 $A\subset_{p}B\nRightarrow B^{c}\subset_{p}A^{c}.$
-2.3 Equivalence relationships of hesitant fuzzy sets
+Proposition 2.29.
+The following statements hold for $A,B,C\in HF(U)$,
+(1) If $A\subset_{p}B$ and $B\subset_{p}C$, then $A\subset_{p}C$.
+(2) If $A\subset_{a}B$ and $B\subset_{a}C$, then $A\subset_{a}C$.
+(3) If $A\subset_{m}B$ and $B\subset_{m}C$, then $A\subset_{m}C$.
+(4) If $A\subset_{s}B$ and $B\subset_{s}C$, then $A\subset_{s}C$.
+(5) If $A\subset_{t}B$ and $B\subset_{t}C$, then $A\subset_{t}C$.
+(6) If $A\subset_{n}B$ and $B\subset_{n}C$, then $A\subset_{n}C$.
+Proof (1) and (2) are obvious.
+(3) $A\subset_{m}B$ and $B\subset_{m}C$ imply that $mean[A(x)]\leqslant mean[B(x)]\leqslant mean[C(x)]$ for all $x\in U$, where $mean[\cdot]$ is the mean value operator. Then $A\subset_{m}C$.
+(4) $A\subset_{s}B$ and $B\subset_{s}C$ imply that $|A(x)|\geqslant|B(x)|\geqslant|C(x)|$ for all $x\in U$.
+Let $C(x)=\{h^{C}_{1},h^{C}_{2},\cdots,h^{C}_{|C(x)|}\}$, $B(x)=\{h^{B}_{1},h^{B}_{2},\cdots,h^{B}_{|B(x)|}\}$ and $A(x)=\{h^{A}_{1},h^{A}_{2},\cdots,h^{A}_{|A(x)|}\}$ be three descending sequences, since $A(x)\subset_{s}B(x)$ and $B(x)\subset_{s}C(x)$, then $h^{C}_{1}\geqslant h^{B}_{1}\geqslant h^{A}_{1}$, $h^{C}_{2}\geqslant h^{B}_{2}\geqslant h^{A}_{2}$$\cdots$, and $h^{C}_{|C(x)|}\geqslant h^{B}_{|C(x)|}\geqslant h^{A}_{|C(x)|}$.
+Hence, $A\subset_{s}C$.
+(5) $A\subset_{t}B$ and $B\subset_{t}C$ imply that $|A(x)|<|B(x)|<|C(x)|$ for all $x\in U$.
+Let $C(x)=\{h^{C}_{1},h^{C}_{2},\cdots,h^{C}_{|C(x)|}\}$, $B(x)=\{h^{B}_{1},h^{B}_{2},\cdots,h^{B}_{|B(x)|}\}$ and $A(x)=\{h^{A}_{1},h^{A}_{2},\cdots,h^{A}_{|A(x)|}\}$ be three descending sequences, since $A(x)\subset_{t}B(x)$ and $B(x)\subset_{t}C(x)$, then $h^{C}_{1}\geqslant h^{B}_{1}\geqslant h^{A}_{1}$, $h^{C}_{2}\geqslant h^{B}_{2}\geqslant h^{A}_{2}$$\cdots$, and $h^{C}_{|A(x)|}\geqslant h^{B}_{|A(x)|}\geqslant h^{A}_{|A(x)|}$.
+Hence, $A\subset_{t}C$.
+(6) $A\subset_{n}B$ and $B\subset_{n}C$ imply that $A^{+}(x)\leqslant B^{-}(x)\leqslant B^{+}(x)\leqslant C^{-}(x)$ for all $x\in U$, then $A\subset_{n}C$$\blacksquare$
+2.6 The situation of equality in hesitant fuzzy sets
 Theorem 2.30.
@@
 For the proofs of (4)-(7), we give a simple example as follows.
-By the results of fuzzy sets ZADE , $(A^{\prime}(x)\cup B^{\prime}(x))\cap A^{\prime}(x)=A^{\prime}(x)$ and $(A^{\prime\prime}(x)\cup B^{\prime\prime}(x))\cap A^{\prime\prime}(x)=A^{\prime\prime}(x)$, i.e., $(A^{+}(x)\cup B^{+}(x))\cap A^{+}(x)=A^{+}(x)$ and $(A^{-}(x)\cup B^{-}(x))\cap A^{-}(x)=A^{-}(x)$, then $(A\cup B)\cap A=_{a}A$.
-(4)-(7) are easy to be proved by the results of fuzzy sets in ZADE $\blacksquare$
+By the results of fuzzy sets [6], $(A^{\prime}(x)\cup B^{\prime}(x))\cap A^{\prime}(x)=A^{\prime}(x)$ and $(A^{\prime\prime}(x)\cup B^{\prime\prime}(x))\cap A^{\prime\prime}(x)=A^{\prime\prime}(x)$, i.e., $(A^{+}(x)\cup B^{+}(x))\cap A^{+}(x)=A^{+}(x)$ and $(A^{-}(x)\cup B^{-}(x))\cap A^{-}(x)=A^{-}(x)$, then $(A\cup B)\cap A=_{a}A$.
+(4)-(7) are easy to be proved by the results of fuzzy sets in [6]$\blacksquare$
 Example 2.32.
-Let $U=\{x\}$, $A=\frac{0.1,0.2,0.3}{x}$, $B=\frac{0.3,0.4,0.5}{x}$, $C=\frac{0.3,0.45,0.5}{x}$.
+Let $U=\{x\}$, $A=\frac{\{0.1,0.2,0.3\}}{x}$, $B=\frac{\{0.3,0.4,0.5\}}{x}$, $C=\frac{\{0.3,0.45,0.5\}}{x}$.
 $(A\cap B)(x)=\{0.1,0.2,0.3,0.3\}$, $((A\cap B)\cup A)(x)=\{0.1,0,1,0.2,0.2,0.3,0.3,0.3\}\neq_{m}A(x)$.
@@
 $\begin{cases}(A\cap B)\cup A\not=A,\\ (A\cup B)\cap A\not=A,\\ (A\cup B)\cap C\not=(A\cap C)\cup(B\cap C),\\ (A\cap B)\cup C\not=(A\cup C)\cap(B\cup C),\\ (A\cap B)\cup A\neq_{n}A,\\ (A\cup B)\cap A\neq_{n}A,\\ (A\cup B)\cap C\neq_{n}(A\cap C)\cup(B\cap C),\\ (A\cap B)\cup C\neq_{n}(A\cup C)\cap(B\cup C).\\ \end{cases}$
-For the $\subset_{p}$ and $=_{p}$ relationships, the upper bound with respect to an object $x\in U$ and a hesitant fuzzy set over $U$ is the unique thing that needs to be investigated.
-Hence, many inclusion relationships and equivalence relationships of fuzzy sets also hold on hesitant fuzzy sets as the forms of $\subset_{p}$ and $=_{p}$ relationships.
-However, $A\subset_{p}B\nRightarrow B^{c}\subset_{p}A^{c}$, then the $\subset_{p}$ and $=_{p}$ relationships are not consistent with the inclusion relationships and equivalence relationships of fuzzy sets.
-Section 2.2 and 2.3 reveal that some common inclusion relationships and equivalence relationships do not hold on hesitant fuzzy sets but hold on classical sets.
+When considering the $\subset_{p}$ and $=_{p}$ relationships, the primary focus lies on determining the upper bound for an object $x\in U$ and a hesitant fuzzy set defined over $U$.
+Consequently, several inclusion and equal situations that are applicable to fuzzy sets can also be extended to hesitant fuzzy sets in the form of $\subset_{p}$ and $=_{p}$ relationships.
+Nevertheless, the absence of the implication $A\subset_{p}B\nRightarrow B^{c}\subset_{p}A^{c}$ indicates that the $\subset_{p}$ and $=_{p}$ relationships do not align with the inclusion relationships and equal situations observed in fuzzy sets.
+Sections 2.2 and 2.3 reveal that certain commonly observed inclusion and equal situations do not apply to hesitant fuzzy sets but are valid for classical sets.
 3 Foundations of families of hesitant fuzzy sets
@@
 If $H_{inf}^{+}(x)\leqslant H_{sup}^{-}(x)$, by Proposition 2.12 (9), ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}\})(x)$ is obvious.
-If $H_{sup}^{-}(x)<H_{inf}^{+}(x)$, let $V=(\bigcap\{H:H\in\mathscr{H}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}\})(x)$, $min((\bigcup\{H:H\in\mathscr{H}\})(x)-V)>H_{inf}^{+}(x)>H_{sup}^{-}(x)>max((\bigcap\{H:H\in\mathscr{H}\})(x)-V)$, by Remark 2.8 (4), ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}\})(x)$$\blacksquare$
+If $H_{sup}^{-}(x)<H_{inf}^{+}(x)$, let $V=(\bigcap\{H:H\in\mathscr{H}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}\})(x)$, $inf((\bigcup\{H:H\in\mathscr{H}\})(x)-V)>H_{inf}^{+}(x)>H_{sup}^{-}(x)>sup((\bigcap\{H:H\in\mathscr{H}\})(x)-V)$, by Remark 2.11 (4), ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}\})(x)$$\blacksquare$
 Theorem 3.34.
@@
 If $h\in(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)$ and $sup_{2}^{-}\leqslant h$, then $h\in(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.
-It means that $max((\bigcup\{H:H\in\mathscr{H}_{1}\})(x)-V)<sup_{2}^{-}$.
-On the other hand, $sup_{2}^{-}<min((\bigcup\{H:H\in\mathscr{H}_{2}\})(x)-V)$, by Remark 2.8 (4), then ($(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.
+It means that $sup((\bigcup\{H:H\in\mathscr{H}_{1}\})(x)-V)<sup_{2}^{-}$.
+On the other hand, $sup_{2}^{-}<inf((\bigcup\{H:H\in\mathscr{H}_{2}\})(x)-V)$, by Remark 2.11 (4), then ($(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.
 (8) For each $x\in U$, let $inf_{1}^{+}=inf\{H^{+}(x):H\in\mathscr{H}_{1}\}$, $sup_{2}^{-}=sup\{H^{-}(x):H\in\mathscr{H}_{2}\}$.
@@
 Base on the condition $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, if $h\in(\bigcap\{H:H\in\mathscr{H}_{1}\})(x)$ and $sup_{2}^{-}\leqslant h$, then $h\in(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.
-It means that $max((\bigcap\{H:H\in\mathscr{H}_{1}\})(x)-V)<sup_{2}^{-}$.
-On the other hand, $sup_{2}^{-}<min((\bigcup\{H:H\in\mathscr{H}_{2}\})(x)-V)$, by Remark 2.8 (4), then ($(\bigcap\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$$\blacksquare$
-Theorem 3.33.
-Let $\mathscr{H}_{1},\mathscr{H}_{2}\sqsubset HF(U)$ be two families of hesitant fuzzy sets over $U$. The following statements hold,
-(1) There is a $H_{\gamma^{*}}\in\mathscr{H}_{2}$ satisfied $H_{\alpha}\subset_{p}H_{\gamma^{*}}$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{p}\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-(2) There is a $H_{\alpha^{*}}\in\mathscr{H}_{1}$ satisfied $H_{\alpha^{*}}\subset_{p}H_{\gamma}$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{p}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-(3) If $H_{\alpha}\subset_{p}H_{\gamma}$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{p}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-(4) There is a $H_{\gamma^{*}}\in\mathscr{H}_{2}$ satisfied $H_{\alpha}\subset_{a}H_{\gamma^{*}}$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{a}\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-(5) There is a $H_{\alpha^{*}}\in\mathscr{H}_{1}$ satisfied $H_{\alpha^{*}}\subset_{a}H_{\gamma}$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{a}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-(6) If $H_{\alpha}\subset_{a}H_{\gamma}$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{a}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-(7) There is a $H_{\gamma^{*}}\in\mathscr{H}_{2}$ satisfied $H_{\alpha}\subset_{n}H_{\gamma^{*}}$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{n}\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-(8) There is a $H_{\alpha^{*}}\in\mathscr{H}_{1}$ satisfied $H_{\alpha^{*}}\subset_{n}H_{\gamma}$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{n}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-(9) If $H_{\alpha}\subset_{n}H_{\gamma}$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, then $\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\}\subset_{n}\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\}$.
-Proof (1) For each $x\in U$, there is $H_{\gamma^{*}}\in\mathscr{H}_{2}$ such that $H_{\alpha}^{+}(x)\leqslant H_{\gamma^{*}}^{+}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma^{*}}^{+}(x)\leqslant sup\{H_{\gamma}^{+}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-Then ($(\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{p}(\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
-(2) For each $x\in U$, $H_{\alpha^{*}}^{+}(x)\leqslant H_{\gamma}^{+}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $inf\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\alpha^{*}}^{+}(x)\leqslant inf\{H_{\gamma}^{+}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-Then ($(\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{p}(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
-(3) For each $x\in U$, $H_{\alpha}^{+}(x)\leqslant H_{\gamma}^{+}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, thus $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma}^{+}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant inf\{H_{\gamma}^{+}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-Then ($(\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{p}(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
-(4) For each $x\in U$, there is $H_{\gamma^{*}}\in\mathscr{H}_{2}$ such that $H_{\alpha}^{-}(x)\leqslant H_{\gamma^{*}}^{-}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $sup\{H_{\alpha}^{-}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma^{*}}^{-}(x)\leqslant sup\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-To combine the result of (1), then (4) holds.
-(5) For each $x\in U$, $H_{\alpha^{*}}^{-}(x)\leqslant H_{\gamma}^{-}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $inf\{H_{\alpha}^{-}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\alpha^{*}}^{-}(x)\leqslant inf\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-To combine the result of (2), then (5) holds.
-(6) For each $x\in U$, $H_{\alpha}^{-}(x)\leqslant H_{\gamma}^{-}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, thus $sup\{H_{\alpha}^{-}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma}^{-}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $sup\{H_{\alpha}^{-}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant inf\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-To combine the result of (3), then (6) holds.
-(7) For each $x\in U$, there is $H_{\gamma^{*}}\in\mathscr{H}_{2}$ such that $H_{\alpha}^{+}(x)\leqslant H_{\gamma^{*}}^{-}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$, then $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma^{*}}^{-}(x)\leqslant sup\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-Then ($(\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{n}(\bigcup\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
-(8) For each $x\in U$, $H_{\alpha^{*}}^{+}(x)\leqslant H_{\gamma}^{-}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $inf\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\alpha^{*}}^{+}(x)\leqslant inf\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-Then ($(\bigcap\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{n}(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$.
-(9) For each $x\in U$, $H_{\alpha}^{+}(x)\leqslant H_{\gamma}^{-}(x)$ for all $H_{\alpha}\in\mathscr{H}_{1}$ and all $H_{\gamma}\in\mathscr{H}_{2}$, thus $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant H_{\gamma}^{-}(x)$ for all $H_{\gamma}\in\mathscr{H}_{2}$, then $sup\{H_{\alpha}^{+}(x):H_{\alpha}\in\mathscr{H}_{1}\}\leqslant inf\{H_{\gamma}^{-}(x):H_{\gamma}\in\mathscr{H}_{2}\}$.
-Then ($(\bigcup\{H_{\alpha}:H_{\alpha}\in\mathscr{H}_{1}\})(x)\subset_{n}(\bigcap\{H_{\gamma}:H_{\gamma}\in\mathscr{H}_{2}\})(x)$$\blacksquare$
-4 Parameter reductions of hesitant fuzzy information systems
-In this section, we investigate some propositions with regard to parameter reductions in hesitant fuzzy information systems, and then present an algorithm for parameter reductions of hesitant fuzzy information systems.
-4.1 Parameter reductions based on $\beta$-partition in hesitant fuzzy information systems
-In the classical information systems, if an object has the characteristic with respect to attribute $e$, then $F(e)(x)$ is denoted as 1, otherwise, $F(e)(x)$ is denoted as 0.
-Inspired by the idea of classical information systems, if $\beta(e)\subset_{p}F(e)(x)$ then $x$ is considered to have the characteristic with respect to attribute $e$ based on $p\beta$-partition in hesitant fuzzy information systems.
-$\beta=\sum\limits_{e_{i}\in E}\frac{\beta(e_{i})}{e_{i}}$ can be considered as an indicator to investigate whether object $x$ has characteristic $e_{i}\in E$, where $\beta(e_{i})$ is a hesitant fuzzy membership degree of $e_{i}$ in indicator $\beta$.
-As shown in Example 4.44, $\beta(e_{1})\not\subset_{p}F(e_{1})(x)$ for $x\in\{x_{1},x_{2},x_{3}\}$, $\beta(e_{1})\subset_{p}F(e_{1})(y)$ for $y\in\{x_{4},x_{5},x_{6}\}$.
-A family of $U$ can be denoted as $N_{e_{1}}^{p\beta}=\{\{x:\beta(e_{1})\not\subset_{p}F(e_{1})(x),x\in U\},\{y:\beta(e_{1})\subset_{p}F(e_{1})(y),y\in U\}\}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$.
-$N_{e_{1}}^{p\beta}(x_{1})=N_{e_{1}}^{p\beta}(x_{2})=N_{e_{1}}^{p\beta}(x_{3})=\{x_{1},x_{2},x_{3}\}$, $N_{e_{1}}^{p\beta}(x_{4})=N_{e_{1}}^{p\beta}(x_{5})=N_{e_{1}}^{p\beta}(x_{6})=\{x_{4},x_{5},x_{6}\}$.
-$N_{e_{1}}^{p\beta}$ is a $p\beta$-partition of $U$ with respect to $e_{1}$.
-Definition 4.34.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$.
-(1) Let $N_{e}^{p\beta}=\{\{x:\beta(e)\subset_{p}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{p}F(e)(y),y\in U\}\}$ be a $p\beta$-partition of $U$ with respect to $e$, $N_{e}^{p\beta}(x)$ be an element of $N_{e}^{p\beta}$ which contains $x$, $N_{A}^{p\beta}(x)=\sqcap\{N_{e}^{p\beta}(x):e\in A\}$, and $Cov^{p\beta}(A)=\{N_{A}^{p\beta}(x):x\in U\}$.
-(2) Let $N_{e}^{a\beta}=\{\{x:\beta(e)\subset_{a}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{a}F(e)(y),y\in U\}\}$ be a $a\beta$-partition of $U$ with respect to $e$, $N_{e}^{a\beta}(x)$ be an element of $N_{e}^{a\beta}$ which contains $x$, $N_{A}^{a\beta}(x)=\sqcap\{N_{e}^{a\beta}(x):e\in A\}$, and $Cov^{a\beta}(A)=\{N_{A}^{a\beta}(x):x\in U\}$.
-(3) Let $N_{e}^{m\beta}=\{\{x:\beta(e)\subset_{m}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{m}F(e)(y),y\in U\}\}$ be a $m\beta$-partition of $U$ with respect to $e$, $N_{e}^{m\beta}(x)$ be an element of $N_{e}^{m\beta}$ which contains $x$, $N_{A}^{m\beta}(x)=\sqcap\{N_{e}^{m\beta}(x):e\in A\}$, and $Cov^{m\beta}(A)=\{N_{A}^{m\beta}(x):x\in U\}$.
-(4) Let $N_{e}^{s\beta}=\{\{x:\beta(e)\subset_{s}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{s}F(e)(y),y\in U\}\}$ be a $s\beta$-partition of $U$ with respect to $e$, $N_{e}^{s\beta}(x)$ be an element of $N_{e}^{s\beta}$ which contains $x$, $N_{A}^{s\beta}(x)=\sqcap\{N_{e}^{s\beta}(x):e\in A\}$, and $Cov^{s\beta}(A)=\{N_{A}^{s\beta}(x):x\in U\}$.
-(5) Let $N_{e}^{t\beta}=\{\{x:\beta(e)\subset_{t}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{t}F(e)(y),y\in U\}\}$ be a $t\beta$-partition of $U$ with respect to $e$, $N_{e}^{t\beta}(x)$ be an element of $N_{e}^{t\beta}$ which contains $x$, $N_{A}^{t\beta}(x)=\sqcap\{N_{e}^{t\beta}(x):e\in A\}$, and $Cov^{t\beta}(A)=\{N_{A}^{t\beta}(x):x\in U\}$.
-(6) Let $N_{e}^{n\beta}=\{\{x:\beta(e)\subset_{n}F(e)(x),x\in U\},\{y:\beta(e)\not\subset_{n}F(e)(y),y\in U\}\}$ be a $n\beta$-partition of $U$ with respect to $e$, $N_{e}^{n\beta}(x)$ be an element of $N_{e}^{n\beta}$ which contains $x$, $N_{A}^{n\beta}(x)=\sqcap\{N_{e}^{n\beta}(x):e\in A\}$, and $Cov^{n\beta}(A)=\{N_{A}^{n\beta}(x):x\in U\}$.
-Proposition 4.35.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $e\in E$. The following statements hold,
-(1) Suppose $N_{e}^{p\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
-(2) Suppose $N_{e}^{a\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
-(3) Suppose $N_{e}^{m\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
-(4) Suppose $N_{e}^{s\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
-(5) Suppose $N_{e}^{t\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
-(6) Suppose $N_{e}^{n\beta}=\{K_{1},K_{2}\}$, then $K_{1}\sqcap K_{2}=\emptyset$ and $K_{1}\sqcup K_{2}=U$.
-Proof (1)-(6) are obvious.
-Proposition 4.36.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $e\in E$. The following statements hold,
-(1) If $y\in N_{e}^{p\beta}(x)$, then $N_{e}^{p\beta}(y)=N_{e}^{p\beta}(x)$.
-(2) If $y\in N_{e}^{a\beta}(x)$, then $N_{e}^{a\beta}(y)=N_{e}^{a\beta}(x)$.
-(3) If $y\in N_{e}^{m\beta}(x)$, then $N_{e}^{m\beta}(y)=N_{e}^{m\beta}(x)$.
-(4) If $y\in N_{e}^{s\beta}(x)$, then $N_{e}^{s\beta}(y)=N_{e}^{s\beta}(x)$.
-(5) If $y\in N_{e}^{t\beta}(x)$, then $N_{e}^{t\beta}(y)=N_{e}^{t\beta}(x)$.
-(6) If $y\in N_{e}^{n\beta}(x)$, then $N_{e}^{n\beta}(y)=N_{e}^{n\beta}(x)$.
-Proof (1)-(6) are obvious.
-Definition 4.37.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $a\in A\sqsubset E$.
-(1) If $Cov^{p\beta}(A)=Cov^{p\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $p\beta$-partition, denoted as $a\in R^{p\beta}(A)$; if $Cov^{p\beta}(A)\neq Cov^{p\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $p\beta$-partition, denoted as $a\in Core^{p\beta}(A)$.
-(2) If $Cov^{a\beta}(A)=Cov^{a\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $a\beta$-partition, denoted as $a\in R^{a\beta}(A)$; if $Cov^{a\beta}(A)\neq Cov^{a\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $a\beta$-partition, denoted as $a\in Core^{a\beta}(A)$.
-(3) If $Cov^{m\beta}(A)=Cov^{m\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $m\beta$-partition, denoted as $a\in R^{m\beta}(A)$; if $Cov^{m\beta}(A)\neq Cov^{m\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $m\beta$-partition, denoted as $a\in Core^{m\beta}(A)$.
-(4) If $Cov^{s\beta}(A)=Cov^{s\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $s\beta$-partition, denoted as $a\in R^{s\beta}(A)$; if $Cov^{s\beta}(A)\neq Cov^{s\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $s\beta$-partition, denoted as $a\in Core^{s\beta}(A)$.
-(5) If $Cov^{t\beta}(A)=Cov^{t\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $t\beta$-partition, denoted as $a\in R^{t\beta}(A)$; if $Cov^{t\beta}(A)\neq Cov^{t\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $t\beta$-partition, denoted as $a\in Core^{t\beta}(A)$.
-(6) If $Cov^{n\beta}(A)=Cov^{n\beta}(A-\{a\})$, then $a$ is called a reducible parameter of hesitant fuzzy information system $(U,F,A)$ based on $n\beta$-partition, denoted as $a\in R^{n\beta}(A)$; if $Cov^{n\beta}(A)\neq Cov^{n\beta}(A-\{a\})$, then $a$ is called an irreducible parameter of $(U,F,A)$ based on $n\beta$-partition, denoted as $a\in Core^{n\beta}(A)$.
-Theorem 4.38.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$. The following statements hold,
-(1) $\sqcup\{C:C\in Cov^{p\beta}(A)\}=U$.
-(2) $\sqcup\{C:C\in Cov^{a\beta}(A)\}=U$.
-(3) $\sqcup\{C:C\in Cov^{m\beta}(A)\}=U$.
-(4) $\sqcup\{C:C\in Cov^{s\beta}(A)\}=U$.
-(5) $\sqcup\{C:C\in Cov^{t\beta}(A)\}=U$.
-(6) $\sqcup\{C:C\in Cov^{n\beta}(A)\}=U$.
-Proof (1) For all $x\in U$, $x\in N_{e}^{p\beta}(x)$, then $x\in\sqcap\{N_{e}^{p\beta}(x):e\in A\}=N_{A}^{p\beta}(x)\in Cov^{p\beta}(A)$. Then, $U\sqsubset\sqcup\{C:C\in Cov^{p\beta}(A)\}$.
-On the other hand, $C\sqsubset U$ for all $C\in Cov^{p\beta}(A)$, then $\sqcup\{C:C\in Cov^{p\beta}(A)\}\sqsubset U$.
-Hence, $\sqcup\{C:C\in Cov^{p\beta}(A)\}=U$.
-The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
-Theorem 4.39.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$. The following statements hold,
-(1) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{p\beta}(A)$.
-(2) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{a\beta}(A)$.
-(3) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{m\beta}(A)$.
-(4) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{s\beta}(A)$.
-(5) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{t\beta}(A)$.
-(6) If $C_{i}\neq C_{j}$, then $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{n\beta}(A)$.
-Proof (1) If there are $C_{i},C_{j}\in Cov^{p\beta}(A)$ such that $C_{i}\sqcap C_{j}\neq\emptyset$, then there is $x\in C_{i}\sqcap C_{j}$.
-Since $C_{i}\neq C_{j}$, then one of $C_{i}-C_{j}\neq\emptyset$ and $C_{j}-C_{i}\neq\emptyset$ holds.
-These two cases can derive the following two cases, case 1, $y\in C_{i}$ and $y\not\in C_{j}$; and case 2, $y\not\in C_{i}$ and $y\in C_{j}$.
-The proof of case 1 is shown as follows,
-(i) Since $x,y\in C_{i}\in Cov^{p\beta}(A)$, then $\beta(e)\subset_{p}F(e)(x)$ if and only if $\beta(e)\subset_{p}F(e)(y)$ for $e\in\overline{A}\sqsubset A$; and $\beta(e^{\prime})\not\subset_{p}F(e^{\prime})(x)$ if and only if $\beta(e^{\prime})\not\subset_{p}F(e^{\prime})(y)$ for others $e^{\prime}\in A-\overline{A}$.
-(ii) Since $x\in C_{j}$, $y\not\in C_{j}$ and $C_{j}\in Cov^{p\beta}(A)$, then there is at least one $e^{*}\in A$ such that $\beta(e^{*})\subset_{p}F(e^{*})(x)$ and $\beta(e^{*})\not\subset_{p}F(e^{*})(y)$, or $\beta(e^{*})\not\subset_{p}F(e^{*})(x)$ and $\beta(e^{*})\subset_{p}F(e^{*})(y)$.
-There is a contradiction in (i) and (ii). Hence, $C_{i}\sqcap C_{j}=\emptyset$.
-The proof of case 2 ($y\not\in C_{i}$ and $y\in C_{j}$) is similar to the proof of case 1 ($y\in C_{i}$ and $y\not\in C_{j}$), and case 2 also can obtain $C_{i}\sqcap C_{j}=\emptyset$.
-Then, $C_{i}\sqcap C_{j}=\emptyset$ for all $C_{i},C_{j}\in Cov^{p\beta}(A)$.
-The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
-Theorem 4.40.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$, $x,y\in U$. The following statements hold,
-(1) If $y\in N_{A}^{p\beta}(x)$, then $N_{A}^{p\beta}(x)=N_{A}^{p\beta}(y)$.
-(2) If $y\in N_{A}^{a\beta}(x)$, then $N_{A}^{a\beta}(x)=N_{A}^{a\beta}(y)$.
-(3) If $y\in N_{A}^{m\beta}(x)$, then $N_{A}^{m\beta}(x)=N_{A}^{m\beta}(y)$.
-(4) If $y\in N_{A}^{s\beta}(x)$, then $N_{A}^{s\beta}(x)=N_{A}^{s\beta}(y)$.
-(5) If $y\in N_{A}^{t\beta}(x)$, then $N_{A}^{t\beta}(x)=N_{A}^{t\beta}(y)$.
-(6) If $y\in N_{A}^{n\beta}(x)$, then $N_{A}^{n\beta}(x)=N_{A}^{n\beta}(y)$.
-Proof (1) Suppose $N_{A}^{p\beta}(x)\neq N_{A}^{p\beta}(y)$, since $N_{A}^{p\beta}(x),N_{A}^{p\beta}(y)\in Cov^{p\beta}(A)$, let $N_{A}^{p\beta}(x)=C_{i}$ and $N_{A}^{p\beta}(y)=C_{j}$.
-By Theorem 4.39, $C_{i}\sqcap C_{j}=\emptyset$.
-On the other hand, $y\in N_{A}^{p\beta}(x)$ and $y\in N_{A}^{p\beta}(y)$, then $\emptyset\neq\{y\}\sqsubset C_{i}\sqcap C_{j}$. There is a contradiction.
-Hence, $N_{A}^{p\beta}(x)=N_{A}^{p\beta}(y)$.
-The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
-Theorem 4.41.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $a\in A\sqsubset E$. The following statements hold,
-(1) $a\in R^{p\beta}(A)$ if and only if $N_{a}^{p\beta}(x)=\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}$ for all $x\in U$.
-(2) $a\in R^{a\beta}(A)$ if and only if $N_{a}^{a\beta}(x)=\sqcup\{N_{A-\{a\}}^{a\beta}(y):y\in N_{a}^{a\beta}(x)\}$ for all $x\in U$.
-(3) $a\in R^{m\beta}(A)$ if and only if $N_{a}^{m\beta}(x)=\sqcup\{N_{A-\{a\}}^{m\beta}(y):y\in N_{a}^{m\beta}(x)\}$ for all $x\in U$.
-(4) $a\in R^{s\beta}(A)$ if and only if $N_{a}^{s\beta}(x)=\sqcup\{N_{A-\{a\}}^{s\beta}(y):y\in N_{a}^{s\beta}(x)\}$ for all $x\in U$.
-(5) $a\in R^{t\beta}(A)$ if and only if $N_{a}^{t\beta}(x)=\sqcup\{N_{A-\{a\}}^{t\beta}(y):y\in N_{a}^{t\beta}(x)\}$ for all $x\in U$.
-(6) $a\in R^{n\beta}(A)$ if and only if $N_{a}^{n\beta}(x)=\sqcup\{N_{A-\{a\}}^{n\beta}(y):y\in N_{a}^{n\beta}(x)\}$ for all $x\in U$.
-Proof (1) The hesitant fuzzy information system $(U,F,A)$ has one more parameter $a$ than $(U,F,A-\{a\})$.
-Then $N_{A}^{p\beta}(x)=\sqcap\{N_{e}^{p\beta}(x):e\in A\}=N_{A-\{a\}}^{p\beta}(x)\sqcap N_{a}^{p\beta}(x)=\sqcap\{N_{e}^{p\beta}(x):e\in A-\{a\}\}\sqcap N_{a}^{p\beta}(x)$.
-The proof of necessity.
-$a\in R^{p\beta}(A)$, then $N_{A}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)$ for all $x\in U$.
-$N_{A}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)\sqcap N_{a}^{p\beta}(x)$, then $N_{A-\{a\}}^{p\beta}(x)\sqsubset N_{a}^{p\beta}(x)$ for all $x\in U$.
-By Proposition 4.36, if $y\in N_{a}^{p\beta}(x)$, then $N_{a}^{p\beta}(x)=N_{a}^{p\beta}(y)$.
-If $y\in N_{a}^{p\beta}(x)$, then $N_{A-\{a\}}^{p\beta}(y)\sqsubset N_{a}^{p\beta}(y)=N_{a}^{p\beta}(x)$, then $\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}\sqsubset N_{a}^{p\beta}(x)$.
-Since $y\in N_{A-\{a\}}^{p\beta}(y)$, then $N_{a}^{p\beta}(x)\sqsubset\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}$ holds.
-Then $N_{a}^{p\beta}(x)=\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}$.
-The proof of sufficiency.
-If $N_{a}^{p\beta}(x)=\sqcup\{N_{A-\{a\}}^{p\beta}(y):y\in N_{a}^{p\beta}(x)\}$, since $x\in N_{a}^{p\beta}(x)$, then $N_{A-\{a\}}^{p\beta}(x)\sqsubset N_{a}^{p\beta}(x)$.
-Then $N_{A}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)\sqcap N_{a}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)$, i.e., $a\in R^{p\beta}(A)$.
-The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
-Theorem 4.42.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $a,b\in A\sqsubset E$. The following statements hold,
-(1) If $N_{a}^{p\beta}=N_{b}^{p\beta}$, then $a\in R^{p\beta}(A)$ ($b\in R^{p\beta}(A)$).
-(2) If $N_{a}^{a\beta}=N_{b}^{a\beta}$, then $a\in R^{a\beta}(A)$ ($b\in R^{a\beta}(A)$).
-(3) If $N_{a}^{m\beta}=N_{b}^{m\beta}$, then $a\in R^{m\beta}(A)$ ($b\in R^{m\beta}(A)$).
-(4) If $N_{a}^{s\beta}=N_{b}^{s\beta}$, then $a\in R^{s\beta}(A)$ ($b\in R^{s\beta}(A)$).
-(5) If $N_{a}^{t\beta}=N_{b}^{t\beta}$, then $a\in R^{t\beta}(A)$ ($b\in R^{t\beta}(A)$).
-(6) If $N_{a}^{n\beta}=N_{b}^{n\beta}$, then $a\in R^{n\beta}(A)$ ($b\in R^{n\beta}(A)$).
-Proof (1) $N_{a}^{p\beta}=N_{b}^{p\beta}$, i.e., $N_{a}^{p\beta}(x)=N_{b}^{p\beta}(x)$ for all $x\in U$, then $N_{A}^{p\beta}(x)=N_{A-\{a,b\}}^{p\beta}(x)\sqcap N_{b}^{p\beta}(x)\sqcap N_{a}^{p\beta}(x)=N_{A-\{a,b\}}^{p\beta}(x)\sqcap N_{b}^{p\beta}(x)\sqcap N_{b}^{p\beta}(x)=N_{A-\{a,b\}}^{p\beta}(x)\sqcap N_{b}^{p\beta}(x)=N_{A-\{a\}}^{p\beta}(x)$, i.e., $a\in R^{p\beta}(A)$.
-$b\in R^{p\beta}(A)$ can be proved by the same manner of the case $a\in R^{p\beta}(A)$.
-The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
-$a\in R^{p\beta}(A)$ or $b\in R^{p\beta}(A)$ holds, however, $\{a,b\}\sqsubset R^{p\beta}(A)$ is not determinate. The situations of (2)-(6) are the same as the situations of (1).
-Theorem 4.43.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $a,b\in A\sqsubset E$. The following statements hold,
-(1) If $b\in R^{p\beta}(A-\{a\})$, then $b\in R^{p\beta}(A)$.
-(2) If $b\in R^{a\beta}(A-\{a\})$, then $b\in R^{a\beta}(A)$.
-(3) If $b\in R^{m\beta}(A-\{a\})$, then $b\in R^{m\beta}(A)$.
-(4) If $b\in R^{s\beta}(A-\{a\})$, then $b\in R^{s\beta}(A)$.
-(5) If $b\in R^{t\beta}(A-\{a\})$, then $b\in R^{t\beta}(A)$.
-(6) If $b\in R^{n\beta}(A-\{a\})$, then $b\in R^{n\beta}(A)$.
-Proof (1) Since $b\in R^{p\beta}(A-\{a\})$, by Theorem 4.41, $N_{b}^{p\beta}(x)=\sqcup\{N_{A-\{a,b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}$.
-Since $N_{A-\{b\}}^{p\beta}(y)=N_{A-\{a,b\}}^{p\beta}(y)\sqcap N_{a}^{p\beta}(y)\sqsubset N_{A-\{a,b\}}^{p\beta}(y)$, then $N_{b}^{p\beta}(x)=\sqcup\{N_{A-\{a,b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}\sqsupset\sqcup\{N_{A-\{b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}$.
-Since $y\in N_{A-\{b\}}^{p\beta}(y)$, then $N_{b}^{p\beta}(x)\sqsubset\sqcup\{N_{A-\{b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}$ holds.
-Based on the results above, then $N_{b}^{p\beta}(x)=\sqcup\{N_{A-\{b\}}^{p\beta}(y):y\in N_{b}^{p\beta}(x)\}$, by Theorem 4.41, $b\in R^{p\beta}(A)$.
-The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
-Table 1. A triple $(U,F,E)$.
-$U$ $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$
-$x_{1}$ 0.5, 0.3, 0.2 0.7, 0.6, 0.5 0.8, 0.7 0.7, 0.2 0.6, 0.5, 0.3, 0.3 0.6, 0.3, 0.3 0.8, 0.8, 0.6, 0.5
-$x_{2}$ 0.4, 0.4, 0.3 0.7, 0.7 0.7, 0.6, 0.5 0.7, 0.5, 0.5 0.7, 0.5, 0.3, 0.3 0.6, 0.3, 0.3 0.6, 0.4, 0.3
-$x_{3}$ 0.3, 0.2, 0.1 0.7, 0.6 0.6, 0.5 0.7, 0.2 0.6, 0.5, 0.3, 0.1 0.6, 0.3 0.3, 0.3
-$x_{4}$ 0.8, 0.7, 0.7 0.4, 0.3 0.3, 0.3, 0.2 0.9, 0.7 0.5, 0.5, 0.2 0.9, 0.8, 0.7, 0.6 0.9, 0.7
-$x_{5}$ 0.8, 0.7, 0.6 0.3, 0.3, 0.2 0.3, 0.3, 0.1 0.8, 0.8, 0.7 0.6, 0.2, 0.2 0.7, 0.6, 0.5, 0.3 0.5, 0.3
-$x_{6}$ 0.7, 0.7, 0.6 0.8, 0.6, 0.5 0.2, 0.1, 0.1 0.8, 0.7, 0.7 0.6, 0.2 0.6, 0.5, 0.3, 0.1 0.5, 0.3, 0.1, 0.1
-Example 4.44.
-$U=\{x_{1},x_{2},\cdots,x_{6}\}$, $E=\{e_{1},e_{2},\cdots,e_{7}\}$, a triple $(U,F,E)$ is shown in Table 1.
-Let $\beta=\frac{0.6,0.5,0.3}{e_{1}}+\frac{0.5,0.5,0.3}{e_{2}}+\frac{0.4,0.3,0.3}{e_{3}}+\frac{0.7,0.6,0.5}{e_{4}}+\frac{0.6,0.3,0.2}{e_{5}}+\frac{0.6,0.3,0.1}{e_{6}}+\frac{0.6,0.3,0.2}{e_{7}}$.
-(1) For the Theorem 4.43 (1), $a\in R^{p\beta}(A)$ and $b\in R^{p\beta}(A)$, however $b\not\in R^{p\beta}(A-\{a\})$.
-$N_{e_{1}}^{p\beta}=N_{e_{3}}^{p\beta}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$, $N_{e_{2}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{6}\},\{x_{4},x_{5}\}\}$.
-Let $a=e_{1}$, $b=e_{3}$ and $A=\{e_{1},e_{2},e_{3}\}$.
-$Cov^{p\beta}(A)=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5}\},\{x_{6}\}\}=Cov^{p\beta}(A-\{a\})=Cov^{p\beta}(A-\{b\})$, i.e., $a\in R^{p\beta}(A)$ and $b\in R^{p\beta}(A)$.
-$Cov^{p\beta}(A-\{a,b\})=\{\{x_{1},x_{2},x_{3},x_{6}\},\{x_{4},x_{5}\}\}$. $Cov^{p\beta}(A-\{a,b\})\neq Cov^{p\beta}(A-\{a\})$, i.e., $b\not\in R^{p\beta}(A-\{a\})$.
-(2) The cases of Theorem 4.43 (2), (3), (4) and (6) are similar to the case of Theorem 4.43 (1).
-$a\in R^{p\beta}(A)$ and $b\in R^{p\beta}(A)$, $b\not\in R^{p\beta}(A-\{a\})$.
-With the similar ways above, $Cov^{a\beta}(A)=Cov^{a\beta}(A-\{a\})=Cov^{a\beta}(A-\{b\})$, $Cov^{a\beta}(A-\{a,b\})\neq Cov^{a\beta}(A-\{a\})$, i.e., $a\in R^{a\beta}(A)$ and $b\in R^{a\beta}(A)$, $b\not\in R^{a\beta}(A-\{a\})$.
-Furthermore, $a\in R^{m\beta}(A)$ and $b\in R^{m\beta}(A)$, $b\not\in R^{m\beta}(A-\{a\})$; $a\in R^{s\beta}(A)$ and $b\in R^{s\beta}(A)$, $b\not\in R^{s\beta}(A-\{a\})$; and $a\in R^{n\beta}(A)$ and $b\in R^{n\beta}(A)$, $b\not\in R^{n\beta}(A-\{a\})$.
-(3) For the Theorem 4.43 (5), $a\in R^{t\beta}(A)$ and $b\in R^{t\beta}(A)$, however $b\not\in R^{t\beta}(A-\{a\})$.
-$N_{e_{5}}^{t\beta}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}=N_{e_{6}}^{t\beta}$, $N_{e_{7}}^{t\beta}=\{\{x_{1}\},\{x_{2},x_{3},x_{4},x_{5},x_{6}\}\}$.
-Let $a=e_{5}$, $b=e_{6}$, $A=\{e_{5},e_{6},e_{7}\}$.
-$Cov^{t\beta}(A)=Cov^{t\beta}(A-\{a\})=Cov^{t\beta}(A-\{b\})=\{\{x_{1}\},\{x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$.
-$Cov^{t\beta}(A-\{a,b\})=\{\{x_{1}\},\{x_{2},x_{3},x_{4},x_{5},x_{6}\}\neq Cov^{t\beta}(A-\{a\})$.
-$a\in R^{t\beta}(A)$ and $b\in R^{t\beta}(A)$, $b\not\in R^{t\beta}(A-\{a\})$.
-Let $\mathscr{B}^{p}=\{B^{p}:Cov^{p\beta}(A)=Cov^{p\beta}(A-B^{p})\}$ and $\mathcal{I}^{p}=\{I^{p}:I^{p}=A-B^{p},B^{p}\in\mathscr{B}^{p}\}$ be two families over $A$ in $S_{p}=(U,F,A)_{\beta}$.
-The similar ways to define ($\mathscr{B}^{a},\mathcal{I}^{a}$), ($\mathscr{B}^{m},\mathcal{I}^{m}$), ($\mathscr{B}^{s},\mathcal{I}^{s}$), ($\mathscr{B}^{t},\mathcal{I}^{t}$) and ($\mathscr{B}^{n},\mathcal{I}^{n}$) in $S_{a}=(U,F,A)_{\beta}$, $S_{m}=(U,F,A)_{\beta}$, $S_{s}=(U,F,A)_{\beta}$, $S_{t}=(U,F,A)_{\beta}$ and $S_{n}=(U,F,A)_{\beta}$, respectively.
-Theorem 4.45.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$. The following statements hold,
-(1) $Core^{p\beta}(A)=\sqcap_{I^{p}\in\mathcal{I}^{p}}I^{p}$.
-(2) $Core^{a\beta}(A)=\sqcap_{I^{a}\in\mathcal{I}^{a}}I^{a}$.
-(3) $Core^{m\beta}(A)=\sqcap_{I^{m}\in\mathcal{I}^{m}}I^{m}$.
-(4) $Core^{s\beta}(A)=\sqcap_{I^{s}\in\mathcal{I}^{s}}I^{s}$.
-(5) $Core^{t\beta}(A)=\sqcap_{I^{t}\in\mathcal{I}^{t}}I^{t}$.
-(6) $Core^{n\beta}(A)=\sqcap_{I^{n}\in\mathcal{I}^{n}}I^{n}$.
-Proof (1) For all $B^{p}\in\mathscr{B}^{p}$, $Core^{p\beta}(A)\sqsubset A-B^{p}$, then $Core^{p\beta}(A)\sqsubset\sqcap_{B^{p}\in\mathscr{B}^{p}}(A-B^{p})=\sqcap_{I^{p}\in\mathcal{I}^{p}}I^{p}$.
-For all $e\not\in Core^{p\beta}(A)$, $\{e\}\in\mathscr{B}^{p}$, then $(Core^{p\beta}(A))^{c}\sqsubset\sqcup_{B^{p}\in\mathscr{B}^{p}}B^{p}$.
-Then $Core^{p\beta}(A)\sqsupset(\sqcup_{B^{p}\in\mathscr{B}^{p}}B^{p})^{c}=\sqcap_{B^{p}\in\mathscr{B}^{p}}(B^{p})^{c}=\sqcap_{B^{p}\in\mathscr{B}^{p}}(A-B^{p})=\sqcap_{I^{p}\in\mathcal{I}^{p}}I^{p}$.
-Hence, $Core^{p\beta}(A)=\sqcap_{I^{p}\in\mathcal{I}^{p}}I^{p}$.
-The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
-Theorem 4.46.
-Let $\beta\in HF(E)$, $F:E\rightarrow HF(U)$, $A\sqsubset E$. The following statements hold,
-(1) $e\in Core^{p\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{p_{1}}=(U,F,A)_{\beta}$ and $S_{p_{2}}=(U,F,A-\{e\})_{\beta}$.
-(2) $e\in Core^{a\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{a_{1}}=(U,F,A)_{\beta}$ and $S_{a_{2}}=(U,F,A-\{e\})_{\beta}$.
-(3) $e\in Core^{m\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{m_{1}}=(U,F,A)_{\beta}$ and $S_{m_{2}}=(U,F,A-\{e\})_{\beta}$.
-(4) $e\in Core^{s\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{s_{1}}=(U,F,A)_{\beta}$ and $S_{s_{2}}=(U,F,A-\{e\})_{\beta}$.
-(5) $e\in Core^{t\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{t_{1}}=(U,F,A)_{\beta}$ and $S_{t_{2}}=(U,F,A-\{e\})_{\beta}$.
-(6) $e\in Core^{n\beta}(A)$ if and only if there is at least a pair of $x,y\in U$ whose relations are different in $S_{n_{1}}=(U,F,A)_{\beta}$ and $S_{n_{2}}=(U,F,A-\{e\})_{\beta}$.
-Proof (1) To denote $A^{\prime}=A-\{e\}$.
-The proof of necessity.
-$e\in Core^{p\beta}(A)$, then $Cov^{p\beta}(A)\neq Cov^{p\beta}(A-\{e\})=Cov^{p\beta}(A^{\prime})$.
-There is $x_{0}\in U$ satisfying $N_{A}^{p\beta}(x_{0})\neq N_{A^{\prime}}^{p\beta}(x_{0})$.
-Then there is $y_{0}$ satisfying $y_{0}\in N_{A}^{p\beta}(x_{0})$ and $y_{0}\not\in N_{A^{\prime}}^{p\beta}(x_{0})$, or another case $y_{0}\not\in N_{A}^{p\beta}(x_{0})$ and $y_{0}\in N_{A^{\prime}}^{p\beta}(x_{0})$.
-There is a pair of $x_{0},y_{0}\in U$ that their relations are different in $S_{p_{1}}=(U,F,A)_{\beta}$ and $S_{p_{2}}=(U,F,A-\{e\})_{\beta}$.
-The proof of sufficiency.
-For any $x,y\in U$, if the relations of $x$ and $y$ are different in $S_{p_{1}}=(U,F,A)_{\beta}$ and $S_{p_{2}}=(U,F,A-\{e\})_{\beta}$, i.e., $y\in N_{A}^{p\beta}(x)$ and $y\not\in N_{A^{\prime}}^{p\beta}(x)$, or another case $y\not\in N_{A}^{p\beta}(x)$ and $y\in N_{A^{\prime}}^{p\beta}(x)$.
-In either case, we can deduce $N_{A}^{p\beta}(x)\neq N_{A^{\prime}}^{p\beta}(x)$, i.e., $Cov^{p\beta}(A)\neq Cov^{p\beta}(A-\{e\})$, then $e\in Core^{p\beta}(A)$.
-The proofs of (2)-(6) are similar to the proof of (1)$\blacksquare$
-4.2 Example and algorithm of parameter reduction in hesitant fuzzy information systems
-In this subsection, we provide an example and an algorithm of parameter reduction based on the $p\beta$-partition in hesitant fuzzy information systems.
-The parameter reductions based on other kinds of $\beta$-partitions can be calculated with the similar method of $p\beta$-partition.
-Example 4.47.
-$(U,F,E)$ and $\beta$ are shown in the Example 4.44.
-Let $A=E$, we can thus calculate the families $\mathscr{B}^{p}$ and $\mathcal{I}^{p}$ of hesitant fuzzy information system $(U,F,A)$ based on $p\beta$-partition.
-We show the processes of calculating the $\mathscr{B}^{p}$ and $\mathcal{I}^{p}$ in the following steps:
-Step 1. Calculate $N_{e_{i}}^{p\beta}$ for every $e_{i}\in A$.
-$N_{e_{1}}^{p\beta}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$,
-$N_{e_{2}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{6}\},\{x_{4},x_{5}\}\}$,
-$N_{e_{3}}^{p\beta}=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5},x_{6}\}\}$,
-$N_{e_{4}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\},\emptyset\}$,
-$N_{e_{5}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{5},x_{6}\},\{x_{4}\}\}$,
-$N_{e_{6}}^{p\beta}=\{\{x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}\},\emptyset\}$,
-$N_{e_{7}}^{p\beta}=\{\{x_{1},x_{2},x_{4}\},\{x_{3},x_{5},x_{6}\}\}$.
-Step 2. Obtain an initial $\mathscr{B}^{p}$, if $e_{i}\in R^{p\beta}(A)$, then $\{e_{i}\}\in\mathscr{B}^{p}$. $|B^{p}|=1$ for every $B^{p}$ in the initial $\mathscr{B}^{p}$.
-In this example, the initial $\mathscr{B}^{p}=\{\{e_{1}\},\{e_{3}\},\{e_{4}\},\newline \{e_{5}\},\{e_{6}\}\}$.
-Step 3.
-Select each $B^{p}_{i}\in\mathscr{B}^{p}$ one by one, if $e_{i}\in R^{p\beta}(A-B^{p}_{i})$ for an $e_{i}\in A-B^{p}_{i}$, then remove $B^{p}_{i}$ from $\mathscr{B}^{p}$ and add $B^{p}_{i}\sqcup\{e_{i}\}$ to $\mathscr{B}^{p}$.
-Here, $e_{i}$ may not be the only one that satisfies $e\in R^{p\beta}(A-B^{p}_{i})$ in some cases.
-Thus, if $e_{j}\in R^{p\beta}(A-B^{p}_{i})$ for $e_{j}\in A-B^{p}_{i}$, then $B^{p}_{i}\sqcup\{e_{j}\}$ is also added to $\mathscr{B}^{p}$.
-For $B^{p}_{1}=\{e_{1}\}$, $e_{4}\in R^{p\beta}(A-B^{p}_{1})$, $e_{5}\in R^{p\beta}(A-B^{p}_{1})$ and $e_{6}\in R^{p\beta}(A-B^{p}_{1})$, then $\{e_{1}\}$ is removed from $\mathscr{B}^{p}$, and $B^{p}_{1}\sqcup\{e_{4}\}=\{e_{1},e_{4}\}$, $B^{p}_{1}\sqcup\{e_{5}\}=\{e_{1},e_{5}\}$ and $B^{p}_{1}\sqcup\{e_{6}\}=\{e_{1},e_{6}\}$ are added to $\mathscr{B}^{p}$.
-Based on the results above, then $\mathscr{B}^{p}=\{\{e_{1},e_{4}\},\{e_{1},e_{5}\},\{e_{1},e_{6}\},\{e_{3}\},\{e_{4}\},\{e_{6}\}\}$.
-To deal with other $B^{p}\in\mathscr{B}^{p}$, then we obtain $\mathscr{B}^{p}=\{\{e_{1},e_{4}\},\{e_{1},e_{5}\},\{e_{1},e_{6}\},\{e_{3},e_{4}\},\{e_{3},e_{5}\},\{e_{3},e_{6}\},\newline \{e_{4},e_{5}\},\{e_{4},e_{6}\},\{e_{5},e_{6}\}\}$.
-Step 4.
-Repeat the similar operations of Step 3 until $e\not\in R^{p\beta}(A-B^{p})$ for every $B^{p}\in\mathscr{B}^{p}$ and every $e\in A-B^{p}$.
-Then, the final $\mathscr{B}^{p}$ is obtained as follows:
-$\mathscr{B}^{p}=\{\{e_{1},e_{4},e_{5},e_{6}\},\{e_{3},e_{4},e_{5},e_{6}\}\}$.
-Step 5. Calculate $\mathcal{I}^{p}$, $\mathcal{I}^{p}=\{A-B^{p}:B^{p}\in\mathscr{B}^{p}\}$.
-$\mathcal{I}^{p}=\{\{e_{2},e_{3},e_{7}\},\{e_{1},e_{2},e_{7}\}\}$.
-$Cov^{p\beta}(A)=Cov^{p\beta}(I^{p})$ for all $I^{p}\in\mathcal{I}^{p}$.
-A detailed algorithm of parameter reduction of hesitant fuzzy information systems based on the $p\beta$-partition is shown as follows,
-Algorithm 1: Parameter reduction of hesitant fuzzy information systems based on $p\beta$-partition.
-Input: Hesitant fuzzy information system $S_{p}=(U,F,A)_{\beta}$.
-Output: $\mathscr{B}^{p}$ and $\mathcal{I}^{p}$.
-1. Calculate $N_{e_{i}}^{p\beta}$ for every $e_{i}\in A$.
-2. Obtain an initial $\mathscr{B}^{p}$, if $e_{i}\in R^{p\beta}(A)$, then $\{e_{i}\}\in\mathscr{B}^{p}$.
-3. While There is $B^{p}\in\mathscr{B}^{p}$ and $e\in A-B^{p}$ such that $e\in R^{p\beta}(A-B^{p})$ do
-4. repeat
-5. $\mathscr{B}^{p}\leftarrow\mathscr{B}^{p}-\{B^{p}\}$.
-6. While There is $e\in A-B^{p}$ that satisfied $e\in R^{p\beta}(A-B^{p})$ do
-7. repeat
-8. $\mathscr{B}^{p}\leftarrow\mathscr{B}^{p}\sqcup\{B^{p}\sqcup\{e\}\}$.
-9. until all the $B^{p}\sqcup\{e\}$ are added to $\mathscr{B}^{p}$ for some $e\in A-B^{p}$ that satisfy $e\in R^{p\beta}(A-B^{p})$.
-10. end while
-11. until $e\not\in R^{p\beta}(A-B^{p})$ for every $B^{p}\in\mathscr{B}^{p}$ and every $e\in A-B^{p}$.
-12. end while
-13. $\mathcal{I}^{p}\leftarrow\{A-B^{p}:B^{p}\in\mathscr{B}^{p}\}$.
-5 Application example of foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems
-Some complex systems, such as large wind turbines and shield tunneling machines, have a few failed samples and lots of factors that may affect system failure.
-These kinds of systems make it difficult to make system diagnosis or decision of system replacement.
-We propose a multi-strength intelligent classifier to give an alternative method to deal with the cases stated above.
-We assumed that $|F(e)(x_{i})|=|F(e)(x_{j})|$ for all $e\in E$ and $x_{i},x_{j}\in U$, then the relationship “$\subset_{t}$” will not be considered in Section 5.
-5.1 Knowledge bases and similar degree functions
-Definition 5.48.
-PALA ; PASA A pair $(U,\mathcal{R})$ is called a knowledge base, where $\mathcal{R}\neq\emptyset$ and $\mathcal{R}$ is a family of equivalence relations on $U$.
-As shown in Example 4.44, let $A=\{e_{1},e_{2},e_{3},e_{4}\}$, $Cov^{p\beta}(A)=\{\{x_{1},x_{2},x_{3}\},\{x_{4},x_{5}\},\{x_{6}\}\}$ can be considered as a knowledge, denoted as $(U,R_{(A,p\beta)})$.
-Furthermore, $R_{(A,p\beta)}(x_{1},x_{2})=1$ or $x_{2}\in[x_{1}]_{R_{(A,p\beta)}}$; $R_{(A,p\beta)}(x_{1},x_{4})=0$ or $x_{4}\not\in[x_{1}]_{R_{(A,p\beta)}}$.
-$Cov^{a\beta}(A)$, $Cov^{m\beta}(A)$, $Cov^{s\beta}(A)$ and $Cov^{n\beta}(A)$ are knowledge generated by $R_{(A,a\beta)}$, $R_{(A,m\beta)}$, $R_{(A,s\beta)}$ and $R_{(A,n\beta)}$, respectively.
-Let *=$\mathcal{R}^{*}=\{R_{(A,p\beta)},R_{(A,a\beta)},R_{(A,m\beta)},R_{(A,s\beta)},\newline R_{(A,n\beta)}\}$, then $(U,\mathcal{R}^{*})$ is a knowledge base.
-Definition 5.49.
-Let $(U,F,A)_{\beta}$ be a hesitant fuzzy information system, $x_{i},x_{j}\in U$, $e\in A\sqsubset E$, and $\neg A=\{\neg e_{i}:e_{i}\in A\}$,
-(1) Attribute recognition functions:
-$\xi^{p\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{p}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
-$\xi^{a\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{a}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
-$\xi^{m\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{m}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
-$\xi^{s\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{s}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
-$\xi^{n\beta}(x_{i})=\widehat{A}$, where $\widehat{A}\sqsubset A\sqcup\neg A$, and $e\in\widehat{A}$ if $\beta(e)\subset_{n}F(e)(x_{i})$; otherwise, $\neg e\in\widehat{A}$.
-(2) Similar degree functions:
-$Sim^{p\beta}(x_{i},x_{j})=\frac{|\xi^{p\beta}(x_{i})\sqcap\xi^{p\beta}(x_{j})|}{|A|}$.
-$Sim^{a\beta}(x_{i},x_{j})=\frac{|\xi^{a\beta}(x_{i})\sqcap\xi^{a\beta}(x_{j})|}{|A|}$.
-$Sim^{m\beta}(x_{i},x_{j})=\frac{|\xi^{m\beta}(x_{i})\sqcap\xi^{m\beta}(x_{j})|}{|A|}$.
-$Sim^{s\beta}(x_{i},x_{j})=\frac{|\xi^{s\beta}(x_{i})\sqcap\xi^{s\beta}(x_{j})|}{|A|}$.
-$Sim^{n\beta}(x_{i},x_{j})=\frac{|\xi^{n\beta}(x_{i})\sqcap\xi^{n\beta}(x_{j})|}{|A|}$.
-As shown in Example 4.44, $\beta(e_{1})\not\subset_{p}F(e_{1})(x_{1})$; and $\beta(e_{i})\subset_{p}F(e_{i})(x_{1})$ for $e_{i}\in E-\{e_{1}\}$.
-It means that the object $x_{1}$ has no characteristic concerning $e_{1}$ based on $p\beta$-partition.
-Then, $\xi^{p\beta}(x_{1})=\{\neg e_{1},e_{2},e_{3},e_{4},e_{5},e_{6},e_{7}\}$.
-In the same manner, $\xi^{p\beta}(x_{5})=\{e_{1},\neg e_{2},\neg e_{3},e_{4},e_{5},e_{6},\neg e_{7}\}$.
-Furthermore, $Sim^{p\beta}(x_{1},x_{5})=\frac{|\xi^{p\beta}(x_{1})\sqcap\xi^{p\beta}(x_{5})|}{|E|}=\frac{|\{e_{4},e_{5},e_{6}\}|}{|E|}=\frac{3}{7}$.
-Theorem 5.50.
-For $X\sqsubset U$, the following statements hold in the knowledge base $(U,\mathcal{R}^{*})$.
-(1) $Sim^{p\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,p\beta)})$.
-(2) $Sim^{a\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,a\beta)})$.
-(3) $Sim^{m\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,m\beta)})$.
-(4) $Sim^{s\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,s\beta)})$.
-(5) $Sim^{n\beta}(x,y)=1$ for $x,y\in X\in(U,R_{(A,n\beta)})$.
-Proof (1) For $X\in(U,R_{(A,p\beta)})=Cov^{p\beta}(A)$ and $x\in X$, since $\{x,y\}\sqsubset X=\sqcap\{N_{e}^{p\beta}(x):e\in A\}$, then $y\in N_{e}^{p\beta}(x)$ for all $e\in A$.
-Therefore, we can obtain that $\xi^{p\beta}(x)=\xi^{p\beta}(y)$, then $Sim^{p\beta}(x,y)=1$.
-The proofs of (2)-(5) are similar to the proof of (1)$\blacksquare$
-5.2 Multi-strength intelligent classifier for health status diagnosis of complex systems
-This subsection presents a multi-strength intelligent classifier in knowledge bases to diagnose the health status of complex systems.
-In the following, the health state partitions of the complex system, the termination training conditions of the intelligent classifier, and the classification way of the intelligent classifier will be introduced.
-5.2.1 Partitions of health states
-In order to use the results of the previous section to diagnose the health status of the system, we divide the health level of the system into five levels according to Definition 2.7, shown in Table 2.
-For example, $x\in U_{n}$ means that it is necessary credible that $x$ is a healthy equipment; $y\in U_{f}$ means that it is not credible that $y$ is a healthy equipment, i.e., $y$ is a failed equipment.
-Table 2. Credible degrees that the equipment is healthy
-Kinds of Credible degrees of Health Sample
-partition systems’ health states grade set
-$n\beta$-partition necessary credible G1 $U_{n}$
-$s\beta$-partition strong credible G2 $U_{s}$
-$m\beta$-partition middle credible G3 $U_{m}$
-$a\beta$-partition middle credible G3 $U_{a}$
-$p\beta$-partition weak credible G4 $U_{p}$
-other cases not credible G5 $U_{f}$
-5.2.2 Conditions under which the multi-strength intelligent classifier terminates training
-If $Sim^{n\beta}(x_{i},x_{j})\geqslant PT$ ($Sim^{n\beta}(x_{i},x_{j})<NT$), then $x_{i}$ and $x_{j}$ are (not) considered as two samples in the same category based on $n\beta$-partition, where PT is a positive threshold and NT is a negative threshold.
-$\frac{1}{C_{|U^{*}_{n}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{C_{|U^{*}_{s}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{s}}Sim^{s\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{|U^{*}_{s}|\times|U^{*}_{n}|}\sum\limits_{x_{i}\in U^{*}_{s},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})<NT,\\ \frac{1}{C_{|U^{*}_{m}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{m}}Sim^{m\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{|U^{*}_{m}|\times|U^{*}_{s}|}\sum\limits_{x_{i}\in U^{*}_{m},x_{j}\in U^{*}_{s}}Sim^{s\beta}(x_{i},x_{j})<NT,\\ \frac{1}{C_{|U^{*}_{a}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{a}}Sim^{a\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{|U^{*}_{a}|\times|U^{*}_{s}|}\sum\limits_{x_{i}\in U^{*}_{a},x_{j}\in U^{*}_{s}}Sim^{s\beta}(x_{i},x_{j})<NT,\\ \frac{1}{C_{|U^{*}_{p}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{p}}Sim^{p\beta}(x_{i},x_{j})\geqslant PT,\\ \frac{1}{|U_{1}|\times|U^{*}_{f}|}\sum\limits_{x_{i}\in U_{1},x_{j}\in U^{*}_{f}}Sim^{p\beta}(x_{i},x_{j})<NT,\\ \frac{1}{|U_{2}|\times|U^{*}_{f}|}\sum\limits_{x_{i}\in U_{2},x_{j}\in U^{*}_{f}}Sim^{m\beta}(x_{i},x_{j})<NT.\\ \end{aligned}\right.$ (1)
-Based on the Proposition 2.9, $\subset_{n}\Rightarrow\subset_{s}\Rightarrow\subset_{a}\Rightarrow\subset_{p}$ and $\subset_{n}\Rightarrow\subset_{s}\Rightarrow\subset_{m}$, then we set $U_{1}=U^{*}_{n}\sqcup U^{*}_{s}\sqcup U^{*}_{a}\sqcup U^{*}_{p}$ and $U_{2}=U^{*}_{n}\sqcup U^{*}_{s}\sqcup U^{*}_{m}$ to distinguish $U^{*}_{f}$, where $U^{*}_{n}=U_{n}\sqcap U_{tain}$, $U^{*}_{s}=U_{s}\sqcap U_{tain}$, $U^{*}_{m}=U_{m}\sqcap U_{tain}$, $U^{*}_{a}=U_{a}\sqcap U_{tain}$, $U^{*}_{p}=U_{p}\sqcap U_{tain}$ and $U^{*}_{f}=U_{f}\sqcap U_{tain}$.
-The conditions for the intelligent classifier to terminate the training are shown in the inequality array (1) above, where $C_{N}^{2}=N\times(N-1)/2$.
-The last two inequalities of the inequality array (1) are used to distinguish the failure samples of the healthy samples, the samples in $U_{1}$ and $U^{*}_{f}$ share fewer of the same attributes based on $p\beta$-partition, and the samples in $U_{2}$ and $U^{*}_{f}$ share fewer of the same attributes based on $m\beta$-partition.
-That is the last two inequalities of the inequality array (1) restrict samples from having many features of both healthy and failed samples and prevent the same sample from being classified as both healthy and failed by the classifier.
-The inequality $\frac{1}{|U^{*}_{s}|\times|U^{*}_{n}|}\sum\limits_{x_{i}\in U^{*}_{s},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})<NT$ restricts samples $x_{i}\in U_{s}$ and $x_{j}\in U_{n}$ from having many features of both G1 and G2 health grades and prevents a sample $x_{i}\in U_{s}$ from being classified as $x_{i}\in U_{n}$, i.e., it prevents the classifier from classifying samples of low-level health states into a set of high-level health states.
-The other inequalities with a condition of “$<NT$” in array (1) serve a similar purpose.
-5.2.3 The classification way of the multi-strength intelligent classifier
-When the training of a multi-strength intelligent classifier is finished, it is used to classify other samples.
-Let $Cl(x)=\{Cl_{n}(x),Cl_{s}(x),Cl_{m}(x),Cl_{a}(x),Cl_{p}(x),\newline Cl_{f}(x)\}$, where
-{$Cl_{n}(x)=\frac{1}{|U^{*}_{n}|}\sum\limits_{x_{i}\in U^{*}_{n}}Sim^{n\beta}(x,x_{i}),\\ Cl_{s}(x)=\frac{1}{|U^{*}_{s}|}\sum\limits_{x_{i}\in U^{*}_{s}}Sim^{s\beta}(x,x_{i}),\\ Cl_{m}(x)=\frac{1}{|U^{*}_{m}|}\sum\limits_{x_{i}\in U^{*}_{m}}Sim^{m\beta}(x,x_{i}),\\ Cl_{a}(x)=\frac{1}{|U^{*}_{a}|}\sum\limits_{x_{i}\in U^{*}_{a}}Sim^{a\beta}(x,x_{i}),\\ Cl_{p}(x)=\frac{1}{|U^{*}_{p}|}\sum\limits_{x_{i}\in U^{*}_{p}}Sim^{p\beta}(x,x_{i}),\\ Cl_{f}(x)=\frac{1}{|U^{*}_{f}|}\sum\limits_{x_{i}\in U^{*}_{f}}Sim^{p\beta}(x,x_{i}).\\ \end{aligned}\right.$ (2)
-The values in $Cl(x)$ have three cases (i), (ii) and (iii) as follows.
-(i) One of the values in $Cl(x)$ is greater than or equal to $PT$, then $x$ is assigned to the class corresponding to the maximum value in $Cl(x)$.
-(ii) More than one value in $Cl(x)$ is greater than or equal to $PT$, then $x$ is assigned to more advanced health states.
-For example, $Cl_{n}(x)\geqslant PT$ and $Cl_{s}(x)\geqslant PT$, then $x$ is assigned to $U_{n}$.
-By Proposition 2.9, $\subset_{n}\Rightarrow\subset_{s}$, then case (ii) exists.
-On the other hand, the inequality $\frac{1}{|U^{*}_{s}|\times|U^{*}_{n}|}\sum\limits_{x_{i}\in U^{*}_{s},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})<NT$ prevents the sample $x\in U_{s}$ from being classified as $x\in U_{n}$.
-Hence, $x\in U_{s}$ cannot obtain $Cl_{n}(x)\geqslant PT$, but $x\in U_{n}$ can obtain $Cl_{s}(x)\geqslant PT$.
-If $Cl_{n}(x)\geqslant PT$ and $Cl_{s}(x)\geqslant PT$, then $x\in U_{n}$.
-(iii) No value in $Cl(x)$ is greater than or equal to $PT$, the classifier is invalidated for $x$.
-For small sample size and high dimension data, it is difficult for the classifier to learn all the features of the data, case (iii) may be present in a new monitoring sample.
-5.2.4 The algorithm of the multi-strength intelligent classifier
-The algorithm for training the multi-strength intelligent classifier is shown as Algorithm 2, where $AcT$ is the threshold of the correct classification rate of the testing samples.
-And, referring back the assumption of Section 5, $|F(e)(x)|=|F(e)(y)|$ for all $x,y\in U$.
-Algorithm 2: Multi-strength intelligent classifier.
-Input: $PT$, $NT$, $AcT$, sample sets $U_{n}$, $U_{s}$, $U_{a}$, $U_{m}$, $U_{p}$ and $U_{f}$, and hesitant fuzzy information system $(U,F,E)$, where $U=U_{n}\sqcup U_{s}\sqcup U_{a}\sqcup U_{m}\sqcup U_{p}\sqcup U_{f}$.
-Output: $U^{*}_{tain}$, $A^{*}$ and $\beta^{*}$.
-1. $\beta(e)\leftarrow\{0,0,\cdots,0\}$, where $|\beta(e)|=|F(e)(x)|$ for each $e\in E$.
-2. $AC\leftarrow 0$.
-3. While $AC<AcT$ do
-4. repeat
-5. Two-thirds of the samples in $U$ are randomly selected as training samples $U_{tain}$ and the other third as testing samples $U_{test}$.
-6. $Sim^{n\beta}(x_{i},x_{j})\leftarrow 0$ for all $x_{i},x_{j}\in U_{n}$.
-7. While One of the inequalities in the inequality array (1) does not hold do
-8. repeat
-9. The genetic algorithm is used to iterate an indicator $\beta$.
-10. Apply Algorithm 1 to calculate $\mathcal{I}^{p}$, $\mathcal{I}^{a}$, $\mathcal{I}^{m}$, $\mathcal{I}^{s}$ and $\mathcal{I}^{n}$.
-11.
-$A\leftarrow I_{p}\sqcup I_{a}\sqcup I_{m}\sqcup I_{s}\sqcup I_{n}$, where $I_{p},I_{a},I_{m},I_{s}$ and $I_{n}$ are the random elements in $\mathcal{I}^{p}$, $\mathcal{I}^{a}$, $\mathcal{I}^{m}$, $\mathcal{I}^{s}$ and $\mathcal{I}^{n}$, respectively.
-12.
-Calculate the factors in the array (1) ($\frac{1}{C_{|U^{*}_{n}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})$, etc), and judge whether all the inequalities in the inequality array (1) hold.
-13. until All the inequalities in the inequality array (1) hold.
-14. end while
-15. Calculate $Cl(x)$ for $x\in U_{test}$ and classify $x$.
-16. Calculate $AC$, where $AC=\frac{N_{0}}{|U_{tset}|}$ and $N_{0}$ is the number of testing samples that were correctly classified.
-17. until $AC\geqslant AcT$.
-18. end while
-19. $U^{*}_{tain}\leftarrow U_{tain}$, $A^{*}\leftarrow A$ and $\beta^{*}\leftarrow\beta$.
-When $U^{*}_{tain}$, $A^{*}$ and $\beta^{*}$ are obtained, a multi-strength intelligent classifier $(U^{*}_{tain},\mathcal{R}^{*},\{Sim\})$ is obtained, where *=$\mathcal{R}^{*}=\{R_{(A,p\beta)},R_{(A,a\beta)},R_{(A,m\beta)},R_{(A,s\beta)},\newline R_{(A,n\beta)}\}$ is illustrated below Definition 5.48, $\{Sim\}$ is a family of similar degree functions and $\{Sim\}=\{Sim^{p\beta}(\cdot),Sim^{a\beta}(\cdot),Sim^{m\beta}(\cdot),Sim^{s\beta}(\cdot),Sim^{n\beta}(\cdot)\}$ that is illustrated in the Definition 5.49.
-5.2.5 An example of the multi-strength intelligent classifier application
-Monitoring data of large wind turbines of the same model with different health states are used to do case analysis as follows.
-For the raw data of devices, let $U=\{x_{1},x_{2},\cdots,x_{|U|}\}$ be a sample set and $data(e)(x_{i})$ be the raw data of $x_{i}\in U$ with respect to the attribute $e$.
-Let $data(e)=\sqcup_{x_{i}\in U}data(e)(x_{i})$ and $h=\frac{d-min(data(e))}{max(data(e))-min(data(e))}$, where $d\in data(e)(x_{i})$ and $h\in F(e)(x_{i})$.
-Then, we can convert the raw data into the data conformed to the value range of the hesitant fuzzy membership degree.
-Example 5.51.
-$U=\{x_{1},x_{2},\cdots,x_{17}\}$ is a sample set. $E=\{e_{1},e_{2},\cdots,e_{26}\}$ is an attribute set. $PT$, $NT$ and $AcT$ are set to 85%, 50% and 95%, respectively.
-The samples are classified based on different health states $U_{n}=\{x_{1},x_{6},x_{13}\}$, $U_{s}=\{x_{2},x_{12},x_{14}\}$, $U_{m}=U_{a}=\{x_{3},x_{5},x_{15}\}$ (the samples $x_{3}$, $x_{5}$ and $x_{15}$ are in a same health grade G3), $U_{p}=\{x_{7},x_{8},x_{11},x_{16}\}$ and $U_{f}=\{x_{4},x_{9},x_{10},x_{17}\}$.
-About two-thirds of the samples from various health states are taken as training samples, and the remaining samples are used as testing samples.
-In the last iteration, $U^{*}_{tain}=\{x_{2},x_{3},x_{4},x_{6},x_{7},x_{9},\newline x_{11},x_{13},x_{14},x_{15},x_{16},x_{17}\}$, $U_{test}=\{x_{1},x_{5},x_{8},x_{10},x_{12}\}$ and $A^{*}=\{e_{1},e_{2},e_{6},e_{7},e_{8},e_{11},e_{14},e_{17},e_{18},e_{19},e_{21},e_{26}\}$, shown as the Table 3 in Appendix.
-In the following, we demonstrate some numerical calculations of Algorithm 2 with the results of the last iteration.
-In $(U^{*}_{tain},\mathcal{R}^{*},\{Sim\})$, $\xi^{n\beta}(x_{6})=(\neg e_{1},e_{2},e_{6},e_{7},e_{8},\newline \neg e_{11},\neg e_{14},\neg e_{17},e_{18},e_{19},\neg e_{21},\neg e_{26})$ and $\xi^{n\beta}(x_{13})=(\neg e_{1},e_{2},e_{6},e_{7},e_{8},\neg e_{11},\neg e_{14},e_{17},e_{18},e_{19},\neg e_{21},\neg e_{26})$.
-Then, $Sim^{n\beta}(x_{6},x_{13})=\frac{|\xi^{n\beta}(x_{6})\sqcap\xi^{n\beta}(x_{13})|}{|A^{*}|}=\frac{11}{12}$.
-Then, $\frac{1}{C_{|U^{*}_{n}|}^{2}}\sum\limits_{x_{i},x_{j}\in U^{*}_{n}}Sim^{n\beta}(x_{i},x_{j})=\frac{11}{12}>PN=85\%$.
-In the same manner, to calculate the other factors in the inequality array (1), we can verify that $U^{*}_{tain}$, $A^{*}$ and $\beta^{*}$ satisfy the other conditions of the inequality array (1).
-Then, we classify the samples in $U_{test}$ to calculate the classification accuracy.
-For $x_{1}\in U_{test}$, $Cl(x_{1})=\{Cl_{n}(x_{1}),Cl_{s}(x_{1}),Cl_{m}(x_{1}),\newline Cl_{a}(x_{1}),Cl_{p}(x_{1}),Cl_{f}(x_{1})\}=\{0.9583,0.9583,0.7083,\newline 0.7083,0.4722,0.0556\}$, $x_{1}$ is assigned to $U_{n}$ (health grade is G1) and classified correctly.
-In the same manner, we can obtain $x_{5}\in U_{m}=U_{a}$ (G3), $x_{8}\in U_{p}$ (G4), $x_{10}\in U_{f}$ (G5), and $x_{12}\in U_{s}$ (G2).
-All five testing samples are correctly classified to their classes, then $AC=\frac{5}{|U_{test}|}=100\%>95\%$, the intelligent classifier passes the test.
-And the intelligent classifier $(U^{*}_{tain},\mathcal{R}^{*},\{Sim\})$ is stored in the knowledge base $(U^{*}_{tain},\mathcal{R}^{*})$.
-When we obtain a new sample $x_{new}$, we can calculate $Cl(x_{new})$ and classify it by using the multi-strength intelligent classifier $(U^{*}_{tain},\mathcal{R}^{*},\{Sim\})$ like the way above.
-Regarding attribute $e$, the larger the value of the hesitant fuzzy membership, the higher the level of the system health status, that is, the system health status is positively correlated with attribute $e$.
-Regarding attribute $e$, the smaller the value of hesitant fuzzy membership is, the higher the level of system health status is, that is, the system health status is negatively correlated with attribute $e$.
-The advantage of the proposed multi-strength intelligent classifier: training classifier does not require analyzing the positive or negative correlation between attributes and the health state of the system before.
-If there are some attributes that are not positively or negatively correlated with the health state of the system, these attributes will affect the similarity functions of the classifier.
-The classifier can eliminate these attributes during the training processes to satisfy the constraints of the inequality array (1).
-This means that the classifier can reduce the dimensionality of high-dimensional data, removing attributes that are not relevant to the result.
-If the objects in $U$ are candidates for investment, the multi-strength intelligent classifier can evaluate the projects and select the optimal investment objects, and the classifier can help us make investment decisions.
-In summary, this multi-strength intelligent classifier is a candidate tool for dimensionality reduction of high-dimensional data, classification, diagnosis, evaluation, decision-making, and so on.
+It means that $sup((\bigcap\{H:H\in\mathscr{H}_{1}\})(x)-V)<sup_{2}^{-}$.
+On the other hand, $sup_{2}^{-}<inf((\bigcup\{H:H\in\mathscr{H}_{2}\})(x)-V)$, by Remark 2.11 (4), then ($(\bigcap\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$$\blacksquare$
 4 Conclusion
-The definition of inclusion relationship is one of the most foundational concepts of sets.
-Based on the hesitant fuzzy membership degree of discrete form, we propose several kinds of inclusion relationships for hesitant fuzzy sets.
-Following the definitions of several kinds of inclusion relationships, some foundational propositions of hesitant fuzzy sets and the families of hesitant fuzzy sets are investigated.
-Some rules hold in classical sets but not in hesitant fuzzy sets are also shown.
-In subsequent studies of hesitant fuzzy sets, researchers must avoid the intuitions of classical sets and cannot use these rules.
-Furthermore, we investigate some foundational theorems of hesitant fuzzy information systems with respect to parameter reductions and present an example and an algorithm in hesitant fuzzy information systems to illustrate the processes of parameter reductions.
-Finally, we propose a multi-strength intelligent classifier to make health state diagnoses for complex systems.
-In our future works, we will investigate more applications of the multi-strength intelligent classifiers, saturation reductions in hesitant fuzzy information systems, and foundational theories of knowledge bases with different strengths of knowledge, among other topics.
+The definition of the inclusion relationship is one of the most foundational concepts of sets.
+Based on the discrete form of hesitant fuzzy membership degrees, we propose several kinds of inclusion relationships for hesitant fuzzy sets.
+Furthermore, we present some propositions of hesitant fuzzy sets and some propositions of the families of hesitant fuzzy sets based on the proposed inclusion relationships.
+Some rules that hold in classical sets do not apply to hesitant fuzzy sets, as shown in Section 2.
+In subsequent studies of hesitant fuzzy sets, researchers must avoid the intuitive understanding of classical sets and cannot use these rules.
+In our future work, we will investigate more applications of multi-strength intelligent classifiers, saturation reductions in HFISs, and foundational theories of knowledge bases with varying strengths of knowledge, among other topics.
 Acknowledgments
 The authors thank the editors and the anonymous reviewers for their helpful comments and suggestions that have led to this improved version of the paper.
-The work was supported by the National Natural Science Foundations of China (No.
-71904043).
+This work was supported by the National Social Science Foundation Youth Project of China (No.
+24CTQ029).